### v1

# TFusionOcc: Student’s t-Distribution Based Object-Centric Multi-Sensor Fusion Framework for 3D Occupancy Prediction

Abstract
3D semantic occupancy prediction enables autonomous vehicles (AVs) to perceive fine-grained geometric and semantic structure of their surroundings from onboard sensors, which is essential for safe decision-making and navigation. Recent models for 3D semantic occupancy prediction have successfully addressed the challenge of describing real-world objects with varied shapes and classes. However, the intermediate representations used by existing methods for 3D semantic occupancy prediction rely heavily on 3D voxel volumes or a set of 3D Gaussians, hindering the model’s ability to efficiently and effectively capture fine-grained geometric details in the 3D driving environment. This paper introduces TFusionOcc, a novel object-centric multi-sensor fusion framework for predicting 3D semantic occupancy. By leveraging multi-stage multi-sensor fusion, Student’s t-distribution, and the T-Mixture model (TMM), together with more geometrically flexible primitives, such as the deformable superquadric (superquadric with inverse warp), the proposed method achieved state-of-the-art (SOTA) performance on the nuScenes benchmark. In addition, extensive experiments were conducted on the nuScenes-C dataset to demonstrate the robustness of the proposed method in different camera and lidar corruption scenarios. The code will be available at: https://github.com/DanielMing123/TFusionOcc
Index Terms:
Autonomous driving, 3D semantic occupancy prediction, Multi-sensor fusion, Environment perception
I Introduction
Accurate and efficient modeling of the scene around autonomous vehicles (AV) is essential for safe and collision-free navigation. As technology advances, the introduction of a 3D semantic occupancy representation has successfully addressed the limitations of traditional 3D object detection and 2D bird’s-eye-view (BEV) segmentation networks, particularly for detecting irregular objects and predicting occupied space in 3D. Thus, 3D semantic occupancy provides a suitable representation of the vehicle’s surrounding environment. This representation inherently ensures geometric consistency of the AV’s driving scene and can accurately describe occluded areas. In addition, a multi-sensor fusion technique for 3D semantic occupancy prediction further enhances an AVs’ ability to model the 3D world.
Recent advances in voxel-based multi-sensor fusion methods [18, 21, 35, 16, 33, 4, 26] have achieved remarkable progress by integrating complementary information from different sensors, such as surround-view cameras, surround-view radars, and lidars. Demonstrating superior robustness and overall performance in several challenging light and weather conditions, these methods heavily rely on volumetric grid calculations (Fig 1, voxel-based approach), which hinders real-time application and makes them unsuitable for edge device deployment. Meanwhile, recent advances in object-centric multi-sensor fusion methods [38, 23] aim to address the limitations of the voxel-based approach by utilizing a set of 3D Gaussian primitives to only model the occupied area in the driving scene (Fig 1, 3D Gaussian-based approach), reducing the computational overhead caused by empty regions by concentrating representational capacity where it matters most. However, their modeling performance is inferior to that of voxel-based approaches. This performance gap is mainly due to the limited geometric shape representation capability of primitives and their weaker robustness to outliers.
Figure 1: Pipeline of three approaches: Voxel-based approach (top-left), 3D-Gaussian-Primitive-based object-centric approach (top-right), and our approach (bottom-left).
To address the aforementioned limitations, we propose TFusionOcc (Fig 1, our approach), a novel Student’s t-distribution-based object-centric multi-sensor fusion framework. We adopt a multi-stage feature-level fusion strategy, including early-stage, middle-stage, and late-stage fusion, to better integrate complementary information across modalities and mitigate the drawbacks of each modality. For object-centric modeling, we adopt a Student’s t-distribution and the T-Mixture model (TMM) as a probability kernel, combined with several types of primitives, such as the general T-Primitive, superquadric, and deformable superquadric (superquadric with inverse-warp), to achieve greater geometric shape representation flexibility and robustness against outliers. The extensive experiments conducted on nuScenes [2] and nuScenes-C [31] datasets demonstrate that our proposed approach not only achieves state-of-the-art (SOTA) performance but also exhibits superior robustness against several sensor corruptions. The main contributions of this paper are summarized as follows:
  • We propose TFusionOcc, a novel object-centric multi-stage multi-sensor fusion framework that leverages different types of T primitives to efficiently and effectively model the 3D driving scene for 3D semantic occupancy prediction.
  • We propose a multi-gate cross-attention fusion module (MGCAFusion), which consists of a weighted summation and gated concatenation fusion modules, to combine the multi-modal features in an effective way.
  • We propose a set of Student’s t-distribution-based primitives, including the general T-Primitive, the T-Superquadric primitive, and the T-Superquadric with inverse warp primitive, to effectively model 3D driving scenes.
  • We thoroughly investigated the robustness of the proposed method with respect to the impact of different perception ranges and different lidar and camera corruption scenarios.
The remainder of this paper is structured as follows. Section II provides an overview of related research and identifies the key differences between this study and previous publications. Section III outlines the general framework of TFusionOcc and offers a detailed explanation of the implementation of each module. Section IV presents the results of our experiments. Finally, Section V provides the conclusion of our work.
II Related Work
II-A Voxel-Based Approach
In OccFusion [18], the authors extract a set of 3D voxel volumes from each different modality sensor, such as surround-view cameras, surround-view radars and lidar, followed by the feature fusion operation through the proposed dynamic feature fusion module to perform multi-sensor fusion-based 3D semantic occupancy prediction. The model exhibits strong robustness under challenging rain and night scenarios. In Co-Occ [21], the authors leverage the volume rendering technique, which originally derived from NeRF [15], to refine the intermediate 3D feature volumes. They also proposed a GSFusion module for efficient multimodal feature fusion, achieving SOTA performance. In OccCylindrical [16], the authors proposed a multi-sensor fusion-based framework that uses a cylindrical partition to perform voxelisation of a lidar-captured 3D point cloud and a pseudo 3D point cloud predicted by surround-view cameras, resulting in better preservation of 3D geometric information. The feature fusion and further refinement operations are performed under the cylindrical coordinate system. In FusionOcc [35], the authors proposed a cross-modality fusion module to generate a high-quality depth-aware visual feature, followed by a 2D to 3D view transformation, resulting in a depth-aware visual feature volumes under the Cartesian coordinate system. The lidar branch conducts voxelization and voxel encoding, resulting in a lidar feature volume also under the Cartesian coordinate system. Different modality feature volumes were fused via feature channel concatenation and further refined by a 3D Encoder. In DAOcc [33], the authors leverage multi-task learning to introduce an additional 3D supervision signal to guide intermediate features, leading to robust generalization and overall SOTA performance. Similarly, in Inverse++ [19], the authors leverage multi-task learning by incorporating an extra auxiliary 3D object detection branch. In addition, SDGOcc [4] incorporates a multi-task head to introduce a 2D semantic mask into the model training procedure, leveraging an additional 2D semantic mask to enhance the precision and densification of depth maps obtained from lidar point cloud projection, thereby boosting the model’s overall performance.
Despite achieving remarkable performance, these voxel-based methods incur redundant calculations for empty voxels, which are often dominant in 3D driving scenarios, resulting in inefficient computation. Our study presents a novel object-centric approach, aiming to address the aforementioned constraints by leveraging T-Primitives to model only occupied space in driving scenarios.
II-B Object-Centric Modelling Approach
In the GaussianFormer series [9, 7], the authors proposed a 3D Gaussian-based object-centric modeling approach to improve computational efficiency. They further proposed a Gaussian-mixture model (GMM) to improve the model’s robustness. The 3D-to-3D Gaussian splatting operator, initially proposed in [10] for 3D-to-2D splatting, is further proposed to boost the model’s efficiency. In QuadricFormer [40], the authors proposed leveraging a superquadric primitive combined with a Gaussian probability kernel to achieve object-centric modeling and benefit from the geometric shape representation’s flexibility, thereby improving model performance. In GaussianFormer3D [38], the authors extend the GaussianFormer work into a multi-sensor fusion scenario. By utilizing a set of 3D Gaussian primitives as an intermediate feature representation, the authors designed a pipeline that enables each 3D Gaussian primitive to aggregate features from different modalities and fuse them to perform 3D primitive-to-3D voxel volume Gaussian splatting. Benefiting from the object-centric modeling technique and multi-modal fusion, the model achieves significant performance and efficiency improvements. In GaussianFusionOcc [23], the authors proposed a GaussianFusionBlock module that performs efficient and high-quality feature aggregation and fusion across different modalities, allowing high-quality 3D Gaussian primitive modeling and leading to SOTA performance.
Though our proposed approach belongs to the object-centric modeling category, the key difference of our work compared to previous ones is that we explored Student’s t-distribution and T-mixture model to enhance the model robustness. Furthermore, we proposed a multi-stage feature fusion pipeline to mitigate the drawbacks of each modality, resulting in better fused features. Lastly, we enhanced the primitive’s geometric flexibility to provide a better intermediate representation, thereby achieving SOTA performance.
III TFusionOcc
This work leverages a set of Student-T-distribution-based primitives combined with a 3D to 3D splatting operation to generate a dense 3D semantic occupancy grid of the surrounding scene by integrating information from surround-view cameras and lidar. Thus, the problem can be formulated as follows:
$P^{T}_{refined}=F\left(Cam^{1},Cam^{2},...,Cam^{N},Lidar,P^{T}\right)$ (1)
$Occ=Splat\left(P^{T}_{refined}\right)$ (2)
where $F$ refers to a multi-sensor fusion framework, $P^{T}$ refers to a set of primitives based on T-distribution $P^{T}=\left\{p^{T}_{1},p^{T}_{2},...,p^{T}_{K}\right\}$, along with a set of query vectors $Q=\left\{q_{1},q_{2},...,q_{K}\right\}$, $P^{T}_{refined}$ refers to a set of T-distribution-based primitives refined through multi-sensor fusion framework, $N$ refers to the total number of surround-view images and $Splat\left(\cdot\right)$ refers to the splatting operation. In the final predicted 3D semantic occupancy $Occ\in R^{\left\{X\times Y\times Z\right\}}$, each grid is assigned a semantic property ranging from 0 to $C$, where $C$ refers to the total amount of semantic classes. In our case, a class value of 0 corresponds to an empty grid.
III-A Overview
Figure 2: Overall architecture of TFusionOcc. The pipeline comprises two different modality branches and a multi-stage feature fusion branch. The camera branch extracts multi-scale visual features and predicts a pseudo 3D point cloud from surround-view images. The pseudo 3D point cloud is further projected and cylindrically partitioned, resulting in camera-based, multi-scale, dense depth maps and a voxel volume defined under cylindrical coordinates. The lidar branch applies a cylindrical partition followed by a 3D encoder to extract the lidar feature. Meanwhile, the lidar point cloud is also projected to generate lidar-based, multi-scale, sparse depth maps. The feature fusion branch adopts a multi-stage fusion strategy to merge all outputs from the two-modality branch and leverages a proposed transformer to refine the T-primitives property through fused features.
The overall architecture is shown in Figure 2. For the camera branch, given $\left\{img^{1},img^{2},...,img^{N}\right\}$ surround-view images, a 2D encoder and a decoder are leveraged to extract $V^{Cam}=\left\{\left\{V_{n}^{l}\right\}_{n=1}^{N}\in\mathbb{R}^{C\times H_{l}\times W_{l}}\right\}_{l=1}^{L=3}$ multi-scale visual features. Meanwhile, a DepthNet is leveraged, resulting in a camera-based pseudo 3D point cloud $Pt^{cam}$ and a multi-scale camera-based dense depth maps $V_{depth}^{cam}=\left\{\left\{V_{depth\_n}^{l}\right\}_{n=1}^{N}\in\mathbb{R}^{D\times H_{l}\times W_{l}}\right\}_{l=1}^{L=3}$, where $D$ refers to the total amount of depth-bins. The pseudo 3D point cloud $Pt^{cam}$ is further divided into cylindrical voxels, resulting in a camera-based 3D voxel volume $Vol^{Cam}_{Cylin}$ defined in cylindrical coordinates. For the lidar branch, given a 3D point cloud $Pt^{lidar}$ from lidar, the cylindrical partition is first applied, resulting in a lidar-based 3D voxel volume $Vol^{Lidar}_{Cylin}$. Then a 3D encoder (e.g., Cylinder3D [39]) is used to extract lidar features $V^{Lidar}$ in cylindrical coordinates. Meanwhile, $Pt^{Lidar}$ is projected onto each surround-view camera frame, resulting in sparse multi-scale lidar-based depth maps $V_{depth}^{lidar}=\left\{\left\{V_{depth\_n}^{l}\right\}_{n=1}^{N}\in\mathbb{R}^{D\times H_{l}\times W_{l}}\right\}_{l=1}^{L=3}$. Then, a multi-stage feature fusion strategy is adopted. For the camera branch side, $V_{depth}^{cam}\oplus V_{depth}^{lidar}$ results in multi-scale fused dense depth maps $V_{depth}^{fuse}$. Then $V_{depth}^{fuse}$ is encoded by a multi-layer perceptron (MLP) and added to $V^{Cam}$, resulting in multi-scale depth-aware visual features $F_{depth}^{Cam}$. In addition, an outer product operation $F_{depth}^{Cam}\otimes V_{depth}^{fuse}$ is conducted, resulting in a multi-scale lifted depth-aware visual feature $Vol^{Cam}_{depth}$. For the lidar branch side, each occupied voxel center of $Vol^{Lidar}_{Cylin}$ serves as an anchor to project onto $V^{Cam}$, thereby aggregating semantic information. The aggregated semantic information is further fused with the lidar feature $Vol^{Lidar}_{Cylin}$ through an early-fusion module, resulting in a semantic-aware lidar feature $Vol^{Lidar}_{sem}$. The projection operation begins with a transition from a cylindrical coordinate system to a Cartesian coordinate system, followed by a transition to a new coordinate frame. The $Vol^{Cam}_{Cylin}$ and $Vol^{Lidar}_{Cylin}$ are fused through the skeleton merge module, resulting in a fused voxel volume $Vol^{Fused}_{Cylin}$ under cylindrical coordinates. The $Vol^{Fused}_{Cylin}$ is then used as primitive anchors to initialize T-Primitives $P^{T}$ under cylindrical coordinates. The $P^{T}$, $Vol^{Cam}_{depth}$ and $Vol^{Lidar}_{sem}$ are fed to the transformer module, which stacks the $N\times$ blocks to obtain a set of refined T-Primitives. Lastly, the refined set of T-Primitives performs a splatting operation, resulting in the final 3D semantic occupancy grid.
III-B Image 2D Encoder and Decoder for Surround-View Images
The purpose of the image encoder is to capture both the spatial and semantic features of the surround-view images. These features serve as the foundation for the subsequent feature fusion procedure, and lastly, the task of 3D semantic occupancy prediction. In our approach, we first use a 2D backbone network (e.g. ResNet50[5]) to extract visual features at multiple scales $V^{Cam}_{Encode}=\left\{\left\{V_{n}^{l}\right\}_{n=1}^{N}\in\mathbb{R}^{C_{l}\times H_{l}\times W_{l}}\right\}_{l=1}^{L=3}$. Subsequently, these features are fused using feature-pyramid networks (FPN), resulting in visual features $V^{Cam}=\left\{\left\{V_{n}^{l}\right\}_{n=1}^{N}\in\mathbb{R}^{C\times H_{l}\times W_{l}}\right\}_{l=1}^{L=3}$ that have the same feature channel dimmension and resolutions that are $\frac{1}{8}$, $\frac{1}{16}$, and $\frac{1}{32}$ of the input image resolution, respectively. The deeper visual feature, with a smaller resolution, contains more semantic information. In contrast, the relatively shallower visual feature, with larger resolutions, provides richer spatial details.
III-C Camera-based Pseudo 3D Point Cloud and Multi-Scale Cam Dense Depth Maps Generation
Figure 3: Inner Structure of DepthNet. The 1/8-scale visual features are first used to generate a 1/8-scale depth map. Then, bilinear interpolation is leveraged to generate 1/16- and 1/32-scale depth maps. Meanwhile, the image-based pseudo-point cloud is generated solely from a 1/8-scale depth map.
The pseudo 3D point cloud $Pt^{cam}$ and camera-based multi-scale dense depth maps $V^{cam}_{depth}$ are generated through the DepthNet module as shown in Figure 3. The DepthNet module takes $V^{Cam}_{Encode}$ as input and leverages a SECONDFPN [32] to generate a single-scale visual feature $V^{cam}_{sec}=\left\{V_{n}\right\}_{n=1}^{N}\in\mathbb{R}^{C\times H_{l}\times W_{l}}$, whose resolution is $\frac{1}{8}$ of the input image. The $V^{cam}_{sec}$ is further passed to an MLP to estimate the ray depth per-pixel, resulting in a single-scale depth map $V_{depth\_n}\in\mathbb{R}^{N\times D\times H_{l}\times W_{l}}$. Bilinear interpolation with a 0.5 downsampling factor is applied iteratively to $V_{depth\_n}$, resulting in two additional depth maps that have $\frac{1}{16}$ and $\frac{1}{32}$ resolution of the input image. Instead of performing depth estimation on each scale visual feature, our approach saves a certain amount of computation overhead while still maintaining a relatively high depth estimation precision. The predicted depth map and the two extra downsampled depth maps form $V_{depth}^{cam}=\left\{\left\{V_{depth\_n}^{l}\right\}_{n=1}^{N}\in\mathbb{R}^{D\times H_{l}\times W_{l}}\right\}_{l=1}^{L=3}$. Meanwhile, the pseudo 3D point cloud $Pt^{cam}$ is acquired only from $\frac{1}{8}$ scale depth map $V_{depth\_n}$. Each pixel ray has a set of depth bins separated by 0.5m intervals. We followed the approach proposed in [7] to use 3D semantic occupancy annotations to supervise depth estimation. For each pixel ray, we take the highest probability depth bins of $K$ to form $Pt^{cam}$.
III-D Skeleton Merge Module
Figure 4: Skeleton Merge Module. The upper lidar-branch serves as the main skeleton to provide a foundation structure for the 3D scene, and the bottom camera-branch serves as an augmentation based on the main skeleton to provide more detailed local structure to compensate for the fine-grained geometry of the main skeleton.
The primitive anchor initialization process is shown in Figure 4. Given the pseudo 3D point cloud derived from the camera $Pt^{cam}$ and the real 3D point cloud of the lidar $Pt^{lidar}$, we first apply the cylindrical partition to both point clouds, resulting in two volumes of voxels $Vol^{Cam}_{Cylin}$ and $Vol^{Lidar}_{Cylin}$ under cylindrical coordinates. We then predefine the total primitive anchors of $M+N$, where the anchors $M$ are sampled from $Vol^{Lidar}_{Cylin}$ and the anchors $N$ are sampled from $Vol^{Cam}_{Cylin}$, with a ratio of $M:N=3:1$. Following [37, 7], we adopt the farthest point sampling strategy to obtain $M$ Lidar anchors from $Vol^{Lidar}_{Cylin}$, which act as the main structural skeleton. Using these Lidar anchors as reference, any voxel in $Vol^{Cam}_{Cylin}$ that spatially overlaps with the LiDAR anchors is removed. In addition, we impose a radius constraint r=5r=5m to perform range filtering. Any voxel in $Vol^{Cam}_{Cylin}$ whose closest Lidar anchor is farther away than $r$ is also discarded. Then, $N$ Camera anchors are obtained using the farthest point sampling strategy on the remaining voxels in $Vol^{Cam}_{Cylin}$. Based on the aforementioned filtering and sampling strategy, we acquire a final set of $M+N$ anchors whose locations will be used to initialize a set of T-primitives.
The skeleton merge module takes advantage of the lidar strength in preserving the fine-grained 3D geometry of the scene and takes advantage of its $Vol^{Lidar}_{Cylin}$ to extract the main skeleton. Meanwhile, the module also leverages the camera’s strength in handling occlusion scenes and uses its $Vol^{Cam}_{Cylin}$ to extract the augmented skeleton surrounding the main skeleton. The fused skeleton preserves the finer-grained 3D geometry of the scene by combining the advantages of both modalities.
III-E Multi-scale-Fused Dense Depth-Maps Generation
Figure 5: Multi-Scale Fused Dense Depth Maps Generation. The upper part exhibits general multi-modality depth map fusion from a depth map perspective, and the bottom part demonstrates the detailed multi-modality depth map fusion from a per-pixel ray perspective.
Given multi-scale camera dense depth maps $V^{cam}_{depth}$ and multi-scale lidar sparse depth maps $V_{depth}^{lidar}$, we first go through per-pixel level value addition followed by a value clamp, resulting in the final multi-scale fused dense depth maps as shown in Figure 5 upper. The detail of the addition of the per-pixel ray value is shown in Figure 5 bottom. For the camera side, each pixel ray has $D$ number of depth bins, and the interval between each depth bin is 0.5m. The DepthNet predicts the probability that each depth bin being occupied with a value in the range of [0, 1]. On the lidar side, each pixel ray also has $D$ depth bins, but only one depth bin has been assigned a probability value of 1.0, while all other depth bins are assigned a probability value of 0. When adding the camera pixel ray to the lidar pixel ray, the resulting fused pixel ray may have some depth-bin probability values exceeding 1.0; therefore, these values need to be clamped within the range [0, 1].
This multi-modality depth map fusion, on the one hand, solves the sparsity problem of lidar-based depth maps, and on the other hand, mitigates the inaccurate depth estimation problem of camera-based depth maps, resulting in denser and highly precise fused depth maps.
III-F Early-Fusion Module
Figure 6: Early-Fusion Module. Each occupied voxel centre in $Vol^{Lidar}_{Cylin}$ serves as an anchor and is projected onto multi-scale visual features to aggregate semantic information, yielding a semantic-aware lidar feature.
The detailed inner structure of the Early-Fusion module is shown in Figure 6. Given the sparse volume of the voxel $Vol_{Cylin}^{Lidar}$ in cylindrical coordinates, we leverage each occupied voxel center as an anchor and convert each anchor position from cylindrical coordinates to Cartesian coordinates. Then, each anchor is projected onto multi-scale visual features $V^{Cam}$ to aggregate semantic information at three scales. The aggregated three-scale visual features are weighted, summed, and concatenated back to the corresponding voxel feature, then fed through an MLP layer to perform the feature fusion, resulting in the last semantic-aware lidar feature $Vol_{sem}^{Lidar}$.
III-G T-Primitive Representation
Following the GaussianFormer [9, 7, 40] series, we propose the probabilistic Student’s T-distribution superposition as an efficient and effective 3D scene representation. Specifically, we propose three types of primitives that are enhanced by the inner kernel of T-distribution or T-distribution-like. The first is the Student-T-distribution-based primitive, which has the property as follows:
$P_{i}^{T}(m_{i},s_{i},r_{i},\alpha_{i},c_{i})$ (3)
where $m_{i}$, $s_{i}$, $r_{i}$, $\alpha_{i}$ and $c_{i}$ refer to the mean, scale value for the $x,y,z$ axes, quaternion vector, opacity and semantic vector of the $i$-th primitive, respectively. The second is the Student-T-distribution-based superquadric primitive, which has the property as follows:
$P_{SQi}^{T}(m_{i},s_{i},r_{i},\alpha_{i},\epsilon_{1\_i},\epsilon_{2\_i},c_{i})$ (4)
where $\epsilon_{1}$ and $\epsilon_{2}$ are the shape exponents of the superquadric. The third is the Student-T-distribution-based superquadric with inverse-warp primitive that has the property as follows:
$P_{SQIWi}^{T}(m_{i},s_{i},r_{i},\alpha_{i},\epsilon_{1\_i},\epsilon_{2\_i},w_{i},c_{i})$ (5)
where $\omega_{i}$ is the learnable weighting scaler valued between -1 and 1.
Similarly to [7], we decompose the 3D modeling target into geometry and semantics predictions, and adopt the multiplication theorem of probability and the T-mixture model to address them probabilistically, respectively. What is more, compared to the Gaussian mixture model, the T-mixture model significantly improves robustness against outliers.
III-G1 Student-T-Distribution Based Geometry Prediction
To restrict T-Primitives to represent only occupied regions for geometry prediction, we interpret the T-distribution primitives $P^{T}=\left\{p^{T}_{i}\right\}_{i=1}^{K}$ as the probability of their surrounding space being occupied. In detail, we assign a probability value of 100% to the centers of the T-distribution, which decays with respect to the distance from the centers $m$:
$\alpha\left(x;p^{T}_{i}\right)=\left[1+\frac{1}{\nu}\left(x-m\right)^{T}\Sigma^{-1}\left(x-m\right)\right]^{-\frac{\nu+3}{2}}$ (6)
where $\nu$ refers to the degree of freedom, it is obtained by calculating how many voxels are intercepted by each T-primitive minus 1, and $\alpha\left(x;p^{T}_{i}\right)$ denotes the probability of the point x being occupied induced by the $i$-th T-primitive. Equation 6 assigns a high probability of occupancy when the point x is close to the center $m$ of the $i$-th T-primitive, which prevents any T-primitive from describing an empty area. To further derive the overall probability of occupancy, we assume that the probabilities of a point being occupied by different T-primitives are mutually independent, and thus we can aggregate them according to the multiplication theorem of probability:
$\alpha\left(x\right)=1-\prod_{i=1}^{K}\left(1-\alpha\left(x;p^{T}_{i}\right)\right)$ (7)
where $\alpha(x)$ represents the overall probability of occupancy at point x.
III-G2 T-Mixture Model Based Semantic Prediction
In addition to the Student-T-distribution-based Geometry modeling, we also need to perform the Student-T-distribution-based semantics prediction. Following [7], we interpret the set of T-primitives as a T-mixture model, where the semantics prediction could be formulated as calculating the expectation of semantics given the probabilistic T-mixture model.
To be specific, we take the original opacity properties $\alpha$ as the prior distribution of the T-primitives, which is normalized by $l1$. Then, the T-primitive probabilistic distribution is adopted as a conditional probability. In addition, we normalize each T-primitive’s original semantics properties $c_{i}$ with a softmax activation function, yielding $\tilde{c}_{i}$ to ensure the boundness of predicted semantics. Lastly, the expectation calculation of the T-mixture model can be formulated as:
$e\left(x;P^{T}\right)=\sum_{i=1}^{K}p\left(P^{T}_{i}|x\right)\tilde{c}_{i}=\frac{{\textstyle\sum_{i=1}^{K}p\left(x|P^{T}_{i}\right)a_{i}\tilde{c}_{i}}}{{\textstyle\sum_{j=1}^{K}p\left(x|P^{T}_{j}\right)a_{j}}}$ (8)
where $p\left(P^{T}_{i}|x\right)$, $p\left(x|P^{T}_{i}\right)$ and $\tilde{c}_{i}$ denote the posterior probability of point x belonging to the $i$-th T-primitive distribution, the conditional probability of point x given the $i$-th T-primitive distribution, and the softmax-normalized semantic properties, respectively.
Lastly, based on the geometry and semantic predictions, we combine the two prediction results together, resulting in the final 3D semantic occupancy prediction. The combination process is formulated as follows.
$Occ=\left[1-\alpha\left(x\right);\alpha\left(x\right)\cdot e\left(x;P^{T}\right)\right]$ (9)
where $1-\alpha\left(x\right)$ refers to the probability of the empty class and $\alpha\left(x\right)\cdot e\left(x;P^{T}\right)$ refers to the final semantic probabilities. The $\alpha\left(x\right)$ is used to weight the initial semantic prediction.
III-G3 Student-T-Distribution Based Primitive
The straightforward T-distribution primitive $p\left(x|P^{T}_{i}\right)$ can be formulated as:
$p\left(x|P^{T}_{i}\right)$ $=\frac{\Gamma\left(\frac{\nu+3}{2}\right)}{\Gamma\left(\frac{\nu}{2}\right)\cdot\left(\nu\pi\right)^{3/2}\cdot\left|\Sigma\right|^{1/2}}\cdot$
$\left[1+\frac{1}{\nu}\left(x-m\right)^{T}\Sigma^{-1}\left(x-m\right)\right]^{-\frac{\nu+3}{2}}$ (10)
where $\nu$ refers to the degree of freedom, the same as in Equation 6, it is obtained by calculating how many voxels are intercepted by each T-primitive minus 1. $\Gamma$ refers to the gamma function, and performs the $\Gamma\left(n\right)=\left(n-1\right)!$ operation. Compared to Gaussian-primitive, this T-primitive is highly tailed in its probability distribution, leading to better robustness against outliers.
III-G4 Student-T-Distribution Based SuperQuadric Primitive
Following the quadricformer [40], we use the T-probabilistic modeling mechanism to convert superquadrics into occupancy probabilities. We first transform the 3D point position $(X,Y,Z)$, which is occupied by the T-Superquadric primitive, into its local coordinate system as:
$\begin{bmatrix}X_{S}\\ Y_{S}\\ Z_{S}\\ \end{bmatrix}=R\times\begin{bmatrix}X-m_{x}\\ Y-m_{y}\\ Z-m_{z}\\ \end{bmatrix}$ (11)
where $(X_{S},Y_{S},Z_{S})$ refers to the local coordinate under the T-Superquadric primitive, and R is the rotation matrix of each T-Superquadric primitive. Then, the occupancy probability of the 3D point $(X,Y,Z)$ associated with each T-Superquadric primitive can be computed as:
$f\left(X_{S},Y_{S},Z_{S}\right)=\left(\left(\frac{X_{S}}{S_{x}}\right)^{\frac{2}{\epsilon_{2}}}+\left(\frac{Y_{S}}{S_{y}}\right)^{\frac{2}{\epsilon_{2}}}\right)^{\frac{\epsilon_{2}}{\epsilon_{1}}}+\left(\frac{Z_{S}}{S_{z}}\right)^{\frac{2}{\epsilon_{1}}}$ (12)
$p\left(x|P^{T}_{SQi}\right)=\left(1+\frac{1}{\nu}\cdot f\left(X_{S},Y_{S},Z_{S}\right)\right)^{-\frac{\nu+3}{2}}$ (13)
where $(S_{X},S_{Y},S_{Z})$ refers to the scale parameters along the X, Y, Z axes, $\epsilon_{1}$ and $\epsilon_{2}$ are the shape exponents of the superquadric.
III-G5 Student-T-Distribution Based SuperQuadric With Inverse-Warp Primitive
So far, thanks to the T-distribution kernel enhancement, the T-Superquadric primitive is robust to outliers. Nevertheless, the T-Superquadric primitive has a very limited shape variety. It can only exhibit convex, symmetric shapes and cannot capture non-convex, asymmetric shapes, which are common in the daily 3D driving scene. To solve this limitation, we make the T-Superquadric primitive deformable through inverse-warp enhancement. Specifically, we choose 24 basis field functions $\left\{B_{i}(u,v,w)\right\}_{i=1}^{24}$, where $u,v,w$ refers to three inputs, respectively, to deform the local coordinates $(X_{S},Y_{S},Z_{S})$ as follows:
$u=\frac{X_{S}}{S_{x}},v=\frac{Y_{S}}{S_{y}},w=\frac{Z_{S}}{S_{z}}$ (14)
$\begin{bmatrix}\tilde{X}_{S}\\ \tilde{Y}_{S}\\ \tilde{Z}_{S}\\ \end{bmatrix}=\begin{bmatrix}X_{S}\\ Y_{S}\\ Z_{S}\\ \end{bmatrix}-\sum_{i=1}^{24}\omega_{i}\cdot B_{i}\left(u,v,w\right)$ (15)
where $\omega_{i}$ is the learnable weighting scaler valued between -1 and 1. Then, the occupancy probability of the 3D point associated with each T-Superquadric-Inverse-Warp primitive can be calculated as follows:
$f\left(\tilde{X}_{S},\tilde{Y}_{S},\tilde{Z}_{S}\right)=\left(\left(\frac{\tilde{X}_{S}}{S_{x}}\right)^{\frac{2}{\epsilon_{2}}}+\left(\frac{\tilde{Y}_{S}}{S_{y}}\right)^{\frac{2}{\epsilon_{2}}}\right)^{\frac{\epsilon_{2}}{\epsilon_{1}}}+\left(\frac{\tilde{Z}_{S}}{S_{z}}\right)^{\frac{2}{\epsilon_{1}}}$ (16)
$p\left(x|P^{T}_{SQIWi}\right)=\left(1+\frac{1}{\nu}\cdot f\left(\tilde{X}_{S},\tilde{Y}_{S},\tilde{Z}_{S}\right)\right)^{-\frac{\nu+3}{2}}$ (17)
Each basis field function is listed in Table I. $B_{1}$ to $B_{9}$ provide constants, linear, basic shears, and taper deformation. $B_{10}$ to $B_{18}$ enable full shear, twist, and bending deformation. Moreover, $B_{19}$ to $B_{24}$ provide quadratics, radial bulge, and smooth corner deformation.
Basis Field $u$ $v$ $w$ $B_{1}$ 1.0 0.0 0.0 $B_{2}$ 0.0 1.0 0.0 $B_{3}$ 0.0 0.0 1.0 $B_{4}$ $u$ 0.0 0.0 $B_{5}$ 0.0 $v$ 0.0 $B_{6}$ 0.0 0.0 $w$ $B_{7}$ $v$ 0.0 0.0 $B_{8}$ $w$ 0.0 0.0 $B_{9}$ 0.0 $w$ 0.0 $B_{10}$ 0.0 $u$ 0.0 $B_{11}$ 0.0 0.0 $u$ $B_{12}$ 0.0 0.0 $v$ $B_{13}$ wv-w*v wuw*u 0.0 $B_{14}$ 0.0 uw-u*w uvu*v $B_{15}$ vwv*w 0.0 vu-v*u $B_{16}$ $w^{2}$ 0.0 0.0 $B_{17}$ 0.0 $w^{2}$ 0.0 $B_{18}$ 0.0 0.0 $u^{2}+v^{2}$ $B_{19}$ $u^{2}$ 0.0 0.0 $B_{20}$ 0.0 $v^{2}$ 0.0 $B_{21}$ 0.0 0.0 $w^{2}$ $B_{22}$ $(u^{2}+v^{2})*u$ $(u^{2}+v^{2})*v$ 0.0 $B_{23}$ uvu*v uvu*v 0.0 $B_{24}$ $u*v^{2}$ $u^{2}*v$ 0.0
TABLE I: The 24 basis field functions make the T-Superquadric deformable. Those functions enable full shears, twists, bends, quadratics, radial bulge and smooth corners.
III-H Transformer
The general inner structure of the proposed transformer is shown in Figure 2. It takes a set of T-primitives, multi-scale lifted depth-aware visual features, and a semantic-aware lidar feature as input, and proceeds through a fused depth-map-guided 3D deformable attention module, a multi-gate cross-attention fusion module, two MLP-based feed-forward modules, a 3D sparse convolution-based self-attention module, and a T-primitive properties refine module iteratively $N$ times, resulting in the last refined T-primitives. The following sections will introduce the key implementation details of each module.
III-H1 Fused Depth Map based 3D Deformable Attention
Figure 7: 3D Deformable Attention For the same visual feature at location $(u_{i},v_{j})$, several different depth values ranging between $[0,1.0]$ along the depth-axis are used to weight the same pixel feature, resulting in the depth encoded visual feature to mitigate the projection ambiguity problem stemming from the projection process.
Taking multi-scale lifted depth-aware visual features and a set of T-primitive as input, we use a 3D deformable attention operator, namely DFA3D [11], which performs feature aggregation on lifted visual features to mitigate the 3D to 2D projection ambiguity problem, to aggregate visual features for each T-primitive. As shown in Figure 2 camera branch and Figure 7, the multi-scale lifted depth-aware visual features $Vol^{Cam}_{depth}$ are obtained through outer product operation $F_{depth}^{Cam}\otimes V_{depth}^{fuse}$, this operation gives an accurate depth encoding for each visual feature in the feature map. For each T-primitive feature sampling $p^{T}_{i}$, we first use the mean location $m_{i}$ of each $p^{T}_{i}$ as the center, convert the center location from the cylindrical coordinate to the cartesian coordinate, and feed its associate query feature $q_{i}$ to an MLP layer, resulting in 13 offsets $\Delta m_{j}$ with respect to the center to form a set of 3D sampling points $m^{samp}=\left\{m^{samp}_{i}=m_{i}+\Delta m_{j}\mid j=1,...,13\right\}$. Then, we leverage extrinsic and intrinsic matrices to project each 3D sampling point $m^{samp}_{i}$ onto the lifted depth-aware visual features, resulting in the location of the sampling of the pixel frame $\bar{m}^{samp}_{i}=\left(u_{i},v_{i},d_{i}\right)$. Lastly, a 3D deformable attention operator DFA3D takes $\bar{m}^{samp}_{i}$ as a sampling location to perform cross-attention visual feature aggregation on $Vol^{Cam}_{depth}$, resulting in an updated query feature $q_{i}^{updated}$. The overall sampling process can be formulated as follows:
$q_{i}^{updated}=\sum_{i=1}^{13}DFA3D\left(q_{i},\bar{m}^{samp}_{i},Vol^{Cam}_{depth}\right)$ (18)
After the 3D cross-attention process, $q_{i}^{updated}$ already aggregates sufficient depth-aware visual features and is passed to the MGCAFusion module for feature fusion. Meanwhile, the set of 3D sampling points $m^{samp}$ for each T-primitive is also passed on to the MGCAFusion module for semantic-aware lidar feature aggregation.
III-H2 Multi-Gate Cross-Attention Fusion (MGCAFusion) Module
Figure 8: Multi-Gate Cross Attention Fusion Module. The left part exhibits the multi-modality weighted summation fusion part, and the right part exhibits the multi-modality gated concatenation fusion part. The outputs of the two major parts are further fused to produce the final fused feature.
Given a set of 3D sampling points $m^{samp}$ for each T-primitive, we apply bilinear interpolation on $Vol^{Lidar}_{sem}$ to aggregate a semantic-aware lidar feature for each T-primitive. The detailed implementation of the multi-gate cross attention fusion module (MGCAFusion) is demonstrated in Figure 8. It consists of two central parts, the weighted summation fusion part, as highlighted in the green rectangle on the left of Figure 8 and the gated concatenation fusion part, as highlighted in the pink rectangle on the right of Figure 8. Both parts take the same inputs, but perform different feature fusion strategies.
For the weighted summation fusion part, the semantic-aware lidar feature and the depth-aware visual feature are summed in elements first, then passed through an MLP layer followed by a softmax layer, resulting in two weight values, $W_{Lidar}$ and $W_{Vis}$. The $W_{Lidar}$ and $W_{Vis}$ are multiplied again with the corresponding semantic-aware lidar feature and depth-aware visual feature, respectively, resulting in the weighted lidar feature and the weighted visual feature. Lastly, the weighted lidar feature and the weighted visual feature are summed and passed through an MLP layer, yielding the summation fusion feature.
For the gated concatenation fusion part, the semantic-aware lidar feat and the depth-aware visual feat are passed through an MLP layer, followed by a sigmoid function, respectively, resulting in a visual gate vector $G_{Vis}$ and a lidar gate vector $G_{Lidar}$. $G_{Vis}$ is element-wise multiplied with the semantic-aware lidar feature, followed by a skip connection that results in the gated lidar feature. $G_{Lidar}$ is element-wise multiplied with the depth-aware visual feature, followed by a skip connection that results in the gated visual feature. The gated Lidar and visual features are concatenated along the feature channel dimension, followed by an MLP layer that reduces the fused feature channel dimension, yielding the concatenation fusion feature.
Lastly, the summation and concatenation fusion features are combined via feature-channel concatenation, followed by an MLP layer, yielding the final fused feature.
III-H3 3D Sparse Conv-based Self-Attention Module
Figure 9: Sparse-Convolution-Based Self-Attention Pipeline. The center of T-primitives is treated as a 3D point to form a primitive-based 3D point cloud. We apply a cylindrical partition to the obtained 3D point cloud, followed by a 3D sparse convolution to realise self-attention.
The self-attention module is demonstrated in Figure 9. Each T-primitive center $m_{i}$ is treated as a 3D point, and we can acquire a set of 3D points based on T-primitive. Then, this T-primitive-based sparse 3D point cloud undergoes a coordinate transformation from Cartesian to cylindrical coordinates, followed by voxelization, yielding a sparse voxel volume in cylindrical coordinates. Lastly, the 3D sparse convolution applies to this sparse voxel volume to perform the self-attention operation. Since our T-primitives are initialized in cylindrical coordinates, we perform the self-attention operation in the same coordinates to better preserve the scene’s fine-grained 3D geometry.
III-H4 T-Primitive Property Refinement Module
Since T-primitive queries have aggregated sufficient semantic and 3D geometry information in the previous 3D deformable attention, self-attention, and MGCAFusion modules, we use each query vector to update its associated T-primitive properties. Specifically, for three types of T-primitives, we feed each type of T-primitive query to an MLP layer to obtain the intermediate properties, and we combine the intermediate properties with the initial properties to achieve the refinement of the T-primitive property. The property refinement details for three types of T-primitives are formulated as:
$MLP(Q_{P_{i}})\Longrightarrow(\hat{m}_{i},\hat{s}_{i},\hat{r}_{i},\hat{c}_{i})$
$MLP(Q_{SQ_{i}})\Longrightarrow(\hat{m}_{i},\hat{s}_{i},\hat{r}_{i},\hat{\epsilon}_{1_{i}},\hat{\epsilon}_{2_{i}},\hat{c}_{i})$
$MLP(Q_{SQIW_{i}})\Longrightarrow(\hat{m}_{i},\hat{s}_{i},\hat{r}_{i},\hat{\epsilon}_{1_{i}},\hat{\epsilon}_{2_{i}},\hat{w}_{i},\hat{c}_{i})$
$P^{T}_{New_{i}}=(m+\hat{m}_{i},\hat{s}_{i},\hat{r}_{i},\hat{c}_{i})$
$P^{T}_{New_{SQ_{i}}}=(m+\hat{m}_{i},\hat{s}_{i},\hat{r}_{i},\hat{\epsilon}_{1_{i}},\hat{\epsilon}_{2_{i}},\hat{c}_{i})$
$P^{T}_{New_{SQIW_{i}}}=(m+\hat{m}_{i},\hat{s}_{i},\hat{r}_{i},\hat{\epsilon}_{1_{i}},\hat{\epsilon}_{2_{i}},\hat{w}_{i},\hat{c}_{i})$ (19)
IV Experimental Results
IV-A Implementation Details
The TFusionOcc model takes six surround-view images and 10 lidar sweeps per data sample and leverages ResNet50-DCN [5] as a 2D Backbone, initialized with weights from FCOS3D [28]. The transformer layer is stacked four times iteratively. AdamW optimizer is used, with an initial learning rate of 2e-4 and a weight decay of 0.01. The learning rate is decayed using a multistep scheduler. The predicted occupancy has a resolution of $200\times 200\times 16$ for full-scale evaluation on nuScenes-Surroundocc, nuScenes-Occ3D and nuScenes-C datasets. For data augmentation, we employ photometric distortions and a grid mask to the input images during training. Model training is conducted on two H100 GPUs with 94GB of memory.
IV-B Loss Function
To train TFusionOcc, we take advantage of the Lovász-Softmax loss [1] $L_{lovasz}$ and the binary cross-entropy loss $L_{BCE}$ to optimize the 3D semantic occupancy task. For depth supervision, a binary cross-entropy loss $L^{depth}_{BCE}$ is used to refine the depth distribution feature. The final loss is formulated as follows.
$L_{Total}=L_{lovasz}+\lambda\cdot L_{BCE}+L_{BCE}^{depth}$ (20)
where $\lambda$ balance the loss weight, in our case, we set $\lambda=10$. The final loss is applied only to the output of the last iterated transformer.
IV-C Dataset
The public nuScenes dataset [2], specifically designed for autonomous driving purposes, serves as the primary data source for our experiments. Furthermore, nuScenes-C [31], which not only augments the existing nuScenes dataset with additional simulated challenging scenarios, such as snow, fog, lowlight and brightness, etc, but also simulates several engineering failure cases, such as frame lost and camera crash, etc, is employed as a supplementary dataset to further evaluate the robustness of the proposed algorithm. Each corruption scenario in nuScenes-C has three severities, ranked easy, mid, and hard, respectively.
To perform the 3D semantic occupancy prediction task, we use dense labels obtained from SurroundOcc [30] and Occ3D [27], respectively. Since the test set lacks semantic labels, we train our model on the training set and evaluate its performance using the validation set. For annotations from SurroundOcc, we set the range of the X and Y axes to [-50, 50] meters and the Z axis to [-5, 3] meters under lidar coordinates. For annotations from Occ3D, we set the range of the X and Y axes to [-40, 40] meters and the Z axis to [-1, 5.4] meters under the ego coordinates. The input images have a resolution of $1600\times 900$ pixels, while the final output of the semantic occupancy prediction branch is represented with a resolution of $200\times 200\times 16$. The annotations from SurroundOcc contain a total of 17 semantic classes, while the annotations from Occ3D comprise 18 semantic classes. Label 0 refers to free voxels in both annotations.
Additionally, following the methodology proposed in [18], we conduct an in-depth analysis of our model performance in challenging scenarios, specifically in rainy and night conditions. This evaluation is carried out using the annotation file provided by [18].
IV-D Performance Evaluate Metrics
To assess the performance of various SOTA algorithms and compare them with our approach in the 3D semantic occupancy prediction task, we utilize the intersection over union (IoU) to evaluate each semantic class. Moreover, we employ the mean IoU over all semantic classes (mIoU) as a comprehensive evaluation metric:
$IoU=\frac{TP}{TP+FP+FN}$ (21)
and
$mIoU=\frac{1}{Cls}\sum_{i=1}^{Cls}\frac{TP_{i}}{TP_{i}+FP_{i}+FN_{i}}$ (22)
where $TP$, FPFP, and $FN$ represent the counts of true positives, false positives, and false negatives in our predictions, respectively, while $Cls$ denotes the total class number.
IV-E Model Performance Analysis
To demonstrate the performance of our proposed TFusionOcc, we evaluated it on the Surroundocc-nuScenes and Occ3D-nuScenes benchmarks. For each benchmark, we evaluate our model across three primitive settings, including T-P (general T-primitive), T-SQ (T-superquadric), and T-SQ-IW (deformable T-superquadric or T-superquadric with inverse-warp), with 12800 and 25600 total primitives, respectively. The results of the experiment on the surroundocc-nuScenes and Occ3D-nuScenes benchmarks are shown in Table II and Table III, respectively.
In both benchmarks, TFusionOcc (T-SQ-IW) settings achieve the best and second-best performance, demonstrating the importance of the primitives’ shape variety and robustness. Specifically, our proposed method excels at detecting background objects, including drivable surfaces, other flat areas, sidewalks, terrain, man-made structures, and vegetation. In addition, our model performs significantly well on relatively small foreground objects, such as barriers, bicycles, cars, motorcycles, pedestrians, and traffic cones.
Our method, which only adopts a single 3D supervision signal, achieved par-performance in comparison with algorithms that leverage multi-task learning to receive dual supervision signals, such as DAOcc [33] in the Surroundocc-nuScenes benchmark that incorporates a 3D object detection head and a 3D semantic occupancy prediction head during training, and SDGOcc [4] in the Occ3D-nuScenes benchmark, which incorporates a 2D semantic segmentation head and a 3D semantic occupancy prediction head during training.
Method 2D-Backbone Modality IoU mIoU $\bullet$ barrier $\bullet$ bicycle $\bullet$ bus $\bullet$ car $\bullet$ const. veh. $\bullet$ motorcycle $\bullet$ pedestrian $\bullet$ traffic cone $\bullet$ trailer $\bullet$ truck $\bullet$ drive. surf. $\bullet$ other flat $\bullet$ sidewalk $\bullet$ terrain $\bullet$ manmade $\bullet$ vegetation MonoScene [3] R101-DCN C 23.96 7.31 4.03 0.35 8.00 8.04 2.90 0.28 1.16 0.67 4.01 4.35 27.72 5.20 15.13 11.29 9.03 14.86 Atlas [20] - C 28.66 15.00 10.64 5.68 19.66 24.94 8.90 8.84 6.47 3.28 10.42 16.21 34.86 15.46 21.89 20.95 11.21 20.54 BEVFormer [13] R101-DCN C 30.50 16.75 14.22 6.58 23.46 28.28 8.66 10.77 6.64 4.05 11.20 17.78 37.28 18.00 22.88 22.17 13.80 22.21 TPVFormer [8] R101-DCN C 30.86 17.10 15.96 5.31 23.86 27.32 9.79 8.74 7.09 5.20 10.97 19.22 38.87 21.25 24.26 23.15 11.73 20.81 C-CONet [29] R101 C 26.10 18.40 18.60 10.00 26.40 27.40 8.60 15.70 13.30 9.70 10.90 20.20 33.00 20.70 21.40 21.80 14.70 21.30 InverseMatrixVT3D [17] R101-DCN C 30.03 18.88 18.39 12.46 26.30 29.11 11.00 15.74 14.78 11.38 13.31 21.61 36.30 19.97 21.26 20.43 11.49 18.47 OccFormer [36] R101-DCN C 31.39 19.03 18.65 10.41 23.92 30.29 10.31 14.19 13.59 10.13 12.49 20.77 38.78 19.79 24.19 22.21 13.48 21.35 FB-Occ [14] R101 C 31.50 19.60 20.60 11.30 26.90 29.80 10.40 13.60 13.70 11.40 11.50 20.60 38.20 21.50 24.60 22.70 14.80 21.60 RenderOcc [22] R101 C 29.20 19.00 19.70 11.20 28.10 28.20 9.80 14.70 11.80 11.90 13.10 20.10 33.20 21.30 22.60 22.30 15.30 20.90 GaussianFormer [9] R101-DCN C 29.83 19.10 19.52 11.26 26.11 29.78 10.47 13.83 12.58 8.67 12.74 21.57 39.63 23.28 24.46 22.99 9.59 19.12 Co-Occ [21] R101 C 30.00 20.30 22.50 11.20 28.60 29.50 9.90 15.80 13.50 8.70 13.60 22.20 34.90 23.10 24.20 24.10 18.00 24.80 GaussianFormer2-256 [7] R101-DCN C 31.74 20.82 21.39 13.44 28.49 30.82 10.92 15.84 13.55 10.53 14.04 22.92 40.61 24.36 26.08 24.27 13.83 21.98 SurroundOcc [30] R101-DCN C 31.49 20.30 20.59 11.68 28.06 30.86 10.70 15.14 14.09 12.06 14.38 22.26 37.29 23.70 24.49 22.77 14.89 21.86 Inverse++ [19] R101-DCN C 31.73 20.91 20.90 13.27 28.40 31.37 11.90 17.76 15.39 13.49 13.32 23.19 39.37 22.85 25.27 23.68 13.43 20.98 LMSCNet [25] - L 36.60 14.90 13.10 4.50 14.70 22.10 12.60 4.20 7.20 7.10 12.20 11.50 26.30 14.30 21.10 15.20 18.50 34.20 L-CONet [29] - L 39.40 17.70 19.20 4.00 15.10 26.90 6.20 3.80 6.80 6.00 14.10 13.10 39.70 19.10 24.00 23.90 25.10 35.70 Co-Occ [21] - L 42.20 22.90 22.00 6.90 25.70 32.40 14.50 13.50 21.00 10.50 18.00 22.50 36.60 21.80 24.60 25.70 31.20 39.90 Co-Occ [21] R101-DCN C+L 41.10 27.10 28.10 16.10 34.00 37.20 17.00 21.60 20.80 15.90 21.90 28.70 42.30 25.40 29.10 28.60 28.20 38.00 GaussianFormer3D [38] R101-DCN C+L 43.30 27.10 26.90 15.80 32.70 36.10 18.60 21.70 24.10 13.00 21.30 29.00 40.60 23.70 27.30 28.20 32.60 42.30 OccFusion [18] R101-DCN C+R 33.97 20.73 20.46 13.98 27.99 31.52 13.68 18.45 15.79 13.05 13.94 23.84 37.85 19.60 22.41 21.20 16.16 21.81 OccFusion [18] R101-DCN C+L 43.53 27.55 25.15 19.87 34.75 36.21 20.03 23.11 25.25 17.50 22.70 30.06 39.47 23.26 25.68 27.57 29.54 40.60 OccFusion [18] R101-DCN C+L+R 43.96 28.27 28.32 20.95 35.06 36.84 20.33 26.22 25.86 19.17 21.27 30.64 40.08 23.75 25.56 27.63 29.82 40.82 GaussianFusionOcc [23] R101-DCN C+L 45.16 30.21 30.22 18.70 35.91 39.57 22.67 27.36 30.10 18.59 24.45 31.25 43.06 25.76 29.12 29.33 34.65 42.70 GaussianFusionOcc [23] R101-DCN C+L+R 45.20 30.37 30.43 18.54 36.23 39.66 22.57 27.35 30.30 19.14 24.56 31.95 42.60 25.82 29.48 29.70 34.78 42.95 OccCylindrical [16] Res-50 C+L 44.94 28.67 26.17 22.12 31.47 36.84 17.95 27.77 29.85 23.90 20.64 28.27 43.00 23.14 27.99 27.81 30.81 40.95 DAOcc [33] Res-50 C+L 45.00 30.50 30.80 19.50 34.00 38.80 25.00 27.70 29.90 22.50 23.20 31.60 41.00 25.90 29.40 29.90 35.20 43.50 TFusionOcc (T-SQ-12800) Res-50 C+L 46.70 30.76 30.67 21.27 33.12 38.96 22.09 27.54 30.47 22.87 21.65 29.86 42.63 27.18 29.80 30.91 36.81 46.26 TFusionOcc (T-SQ-IW-12800) Res-50 C+L 46.92 31.21 31.65 20.13 34.75 39.70 23.08 27.00 30.89 23.33 22.02 30.40 43.11 27.78 30.25 31.47 37.15 46.60 TFusionOcc (T-P-25600) Res-50 C+L 45.49 30.04 30.15 19.58 34.20 38.64 22.10 26.60 29.85 22.24 22.44 29.65 41.08 25.19 28.43 29.71 35.77 45.01 TFusionOcc (T-SQ-25600) Res-50 C+L 46.14 30.88 30.87 21.26 33.68 39.25 22.54 28.34 31.45 23.61 22.80 30.16 41.76 26.12 29.23 30.26 36.77 46.06 TFusionOcc (T-SQ-IW-25600) Res-50 C+L 47.01 31.47 31.96 20.78 34.57 39.73 24.02 27.82 31.19 23.82 23.05 30.67 43.17 27.18 30.13 31.18 37.52 46.71
TABLE II: 3D semantic occupancy prediction results on SurroundOcc-nuScenes validation set. Our approach outperforms other existing methods with the same input modality. For readers’ reference, the bottom of the table presents results from three additional methods using different input modalities. All methods are trained with dense occupancy labels from SurroundOcc [30]. Notion of modality: Camera (C), Lidar (L), Radar (R). The best and second-best performance is indicated in bold red and bold blue, respectively.
Method Mask Backbone Modality mIoU $\bullet$ others $\bullet$ barrier $\bullet$ bicycle $\bullet$ bus $\bullet$ car $\bullet$ const. veh. $\bullet$ motorcycle $\bullet$ pedestrian $\bullet$ traffic cone $\bullet$ trailer $\bullet$ truck $\bullet$ drive. surf. $\bullet$ other flat $\bullet$ sidewalk $\bullet$ terrain $\bullet$ manmade $\bullet$ vegetation MonoScene [3] Res-101 C 6.06 1.75 7.23 4.26 4.93 9.38 5.67 3.98 3.01 5.90 4.45 7.17 14.91 6.32 7.92 7.43 1.01 7.65 BEVDet [6] Res-101 C 11.73 2.09 15.29 0.0 4.18 12.97 1.35 0.0 0.43 0.13 6.59 6.66 52.72 19.04 26.45 21.78 14.51 15.26 BEVFormer [13] Res-101 C 23.67 5.03 38.79 9.98 34.41 41.09 13.24 16.50 18.15 17.83 18.66 27.70 48.95 27.73 29.08 25.38 15.41 14.46 BEVStereo [12] Res-101 C 24.51 5.73 38.41 7.88 38.70 41.20 17.56 17.33 14.69 10.31 16.84 29.62 54.08 28.92 32.68 26.54 18.74 17.49 TPVFormer [8] Res-101 C 28.34 6.67 39.20 14.24 41.54 46.98 19.21 22.64 17.87 14.54 30.20 35.51 56.18 33.65 35.69 31.61 19.97 16.12 OccFormer [36] Res-101 C 21.93 5.94 30.29 12.32 34.40 39.17 14.44 16.45 17.22 9.27 13.90 26.36 50.99 30.96 34.66 22.73 6.76 6.97 RenderOcc [22] Swin-B C 26.11 4.84 31.72 10.72 27.67 26.45 13.87 18.20 17.67 17.84 21.19 23.25 63.20 36.42 46.21 44.26 19.58 20.72 CTF-Occ [27] Res-101 C 28.53 8.09 39.33 20.56 38.29 42.24 16.93 24.52 22.72 21.05 22.98 31.11 53.33 33.84 37.98 33.23 20.79 18.00 Inverse++ [19] Res-101 C 31.04 9.56 41.91 23.53 42.38 46.35 18.61 28.03 26.61 24.77 25.93 34.81 60.10 33.23 37.62 34.83 19.20 20.26 RadOcc [34] Swin-B C+L 49.38 10.90 58.20 25.00 57.90 62.90 34.00 33.50 50.10 32.10 48.90 52.10 82.90 42.70 55.30 58.30 68.60 66.00 OccFusion [18] R101-DCN C+L+R 46.67 12.40 50.30 31.50 57.60 58.80 34.00 41.00 47.20 29.70 42.00 48.00 78.40 35.70 47.30 52.70 63.50 63.30 OccFusion [18] R101-DCN C+L 48.74 12.40 51.80 33.00 54.60 57.70 34.00 43.00 48.40 35.50 41.20 48.60 83.00 44.70 57.10 60.00 62.50 61.30 GaussianFormer3D [38] R101-DCN C+L 46.40 9.80 50.00 31.30 54.00 59.40 28.10 36.20 46.20 26.70 40.20 49.70 79.10 37.30 49.00 55.00 69.10 67.60 RM2Occ [24] R101-DCN C+L 47.82 13.34 54.53 33.81 58.30 59.97 34.45 34.89 29.43 31.53 39.91 42.44 96.01 49.76 61.28 67.62 53.71 51.96 SDGOcc [4] Res-50 C+L 51.66 13.20 57.80 24.30 60.30 64.30 36.20 39.40 52.40 35.80 50.90 53.70 84.60 47.50 58.00 61.60 70.70 67.70 EFFOcc [26] Res-50 C+L 52.82 12.09 59.67 33.39 61.76 64.98 35.46 46.01 57.09 41.04 47.87 54.59 82.76 43.95 56.37 60.23 71.12 69.60 TFusionOcc (T-SQ-12800) Res-50 C+L 52.34 14.82 58.85 39.48 55.10 62.24 35.42 44.19 58.61 47.17 44.92 48.43 82.54 42.87 54.65 59.52 71.05 69.96 TFusionOcc (T-SQ-IW-12800) Res-50 C+L 52.97 14.90 60.59 40.93 56.48 63.68 35.25 45.96 58.37 48.62 44.54 49.84 83.42 43.85 55.09 59.67 69.86 69.51 TFusionOcc (T-P-25600) Res-50 C+L 51.51 14.75 58.10 37.73 51.21 62.03 35.55 44.82 55.18 43.70 44.33 49.24 82.71 42.59 54.83 59.27 70.21 69.41 TFusionOcc (T-SQ-25600) Res-50 C+L 51.70 14.89 59.19 37.64 53.96 62.40 33.88 44.74 58.41 47.17 41.08 47.23 82.46 43.25 54.80 59.04 70.34 68.39 TFusionOcc (T-SQ-IW-25600) Res-50 C+L 53.35 14.66 60.25 40.80 58.39 63.67 34.81 45.97 59.04 48.72 45.32 50.43 83.21 44.70 55.55 60.22 71.41 69.75
TABLE III: 3D semantic occupancy prediction results on Occ3D-nuScenes benchmark. All methods are trained with dense occupancy labels from Occ3D [27]. The camera visibility mask is not used during the training phase. Notion of modality: Camera (C), Lidar (L), Radar (R). The best and second-best performance is indicated in bold red and bold blue, respectively.
IV-F Challenging Scenarios Performance Analysis
To further examine the performance and robustness of the proposed algorithm, we follow [18] to evaluate TFusionOcc with different T-primitive representations under challenging rainy and night-time scenarios and the experiment results are shown in Table IV and Table V, respectively.
In the rainy scenario, raindrops on the camera lens cause focus blur and degrade the quality of semantic information provided by the camera branch. Furthermore, laser beam reflections during rainy conditions degraded the quality of the geometry information provided by the lidar branch. As a result, all algorithms in the benchmark exhibit varying degrees of performance degradation. However, our proposed method benefits from a multi-stage feature fusion design and the flexible T-Primitive representation, exhibits the least performance degradation in the benchmark, demonstrating high robustness against challenging rainy conditions.
In the nighttime scenario, due to the nature of the camera sensor, which is sensitive to the ambient lightning conditions, the dim lighting condition during the nighttime leads to semantic information loss in the camera branch. Meanwhile, the depth-estimation accuracy from the camera branch drops. As a result, all algorithms in the benchmark experience severe performance degradation around 1.56% to 11.10% under the IoU metric and 8.89% to 12.50% under the mIoU metric. Among all performance degradation algorithms, our method still achieves SOTA performance, demonstrating superior feature fusion and information filtering capabilities.
Method Backbone Modality IoU mIoU $\bullet$ barrier $\bullet$ bicycle $\bullet$ bus $\bullet$ car $\bullet$ const. veh. $\bullet$ motorcycle $\bullet$ pedestrian $\bullet$ traffic cone $\bullet$ trailer $\bullet$ truck $\bullet$ drive. surf. $\bullet$ other flat $\bullet$ sidewalk $\bullet$ terrain $\bullet$ manmade $\bullet$ vegetation InverseMatrixVT3D [17] R101-DCN C 29.72 (-0.31%) 18.99 (+0.11%) 18.55 14.29 22.28 30.02 10.19 15.20 10.03 9.71 13.28 20.98 37.18 23.47 27.74 17.46 10.36 23.13 GaussianFormer [9] R101-DCN C 27.37 (-2.46%) 16.96 (-2.14%) 18.16 9.58 21.09 26.83 8.04 10.13 7.80 5.84 12.66 18.24 35.53 18.51 27.79 19.23 11.04 20.85 Co-Occ [21] R101 C 28.90 (-1.10%) 19.70 (-0.60%) 22.10 17.60 26.30 30.80 10.90 9.90 8.20 9.70 11.40 19.30 39.00 22.20 32.60 23.00 11.50 21.30 GaussianFormer2-256 [7] R101-DCN C 31.14 (-0.60%) 20.36 (-0.46%) 19.84 13.52 26.89 31.65 10.82 15.16 9.04 8.41 13.72 21.84 40.51 24.57 32.21 20.65 12.64 24.33 SurroundOcc [30] R101-DCN C 30.57 (-0.92%) 19.85 (-0.45%) 21.40 12.75 25.49 31.31 11.39 12.65 8.94 9.48 14.51 21.52 35.34 25.32 29.89 18.37 14.44 24.78 Inverse++ [19] R101-DCN C 31.32 (-0.41%) 20.66 (-0.25%) 22.52 13.79 25.49 31.80 11.70 16.72 11.14 10.12 12.29 22.25 38.78 23.93 31.62 21.14 12.65 24.61 Co-Occ [21] R101 C+L 40.30 (-0.80%) 26.60 (-0.50%) 26.60 19.10 37.60 37.20 15.90 20.30 16.30 12.30 23.30 27.00 41.00 22.80 35.20 24.60 27.80 39.30 OccFusion [18] R101-DCN C+L 42.67 (-0.86%) 26.68 (-0.87%) 20.91 18.39 35.26 36.19 17.69 19.05 19.40 17.08 23.86 28.86 38.28 26.37 31.44 21.35 29.48 43.22 OccFusion [18] R101-DCN C+L+R 42.67 (-1.29%) 27.39 (-0.88%) 27.82 21.10 36.00 37.10 17.23 21.67 20.34 17.46 20.93 28.57 38.99 24.72 31.96 21.26 29.64 43.53 GaussianFusionOcc [23] R101-DCN C+L 44.28 (-0.88%) 29.19 (-1.02%) 28.10 19.84 36.28 38.90 18.11 21.13 26.14 17.95 25.79 29.92 41.72 27.35 34.99 22.85 33.84 44.10 GaussianFusionOcc [23] R101-DCN C+L+R 44.36 (-0.84%) 29.86 (-0.51%) 28.40 19.88 38.87 39.33 20.61 26.05 25.66 17.97 26.07 31.02 41.70 24.94 35.27 24.08 33.90 44.00 OccCylindrical [16] Res-50 C+L 44.08 (-0.86%) 28.07 (-0.60%) 24.50 22.66 33.47 36.79 16.96 23.63 23.79 23.49 22.70 28.14 41.67 22.32 33.55 21.44 30.58 43.45 DAOcc [33] Res-50 C+L 44.51 (-0.49%) 29.65 (-0.85%) 28.50 19.63 38.01 38.98 21.42 20.19 24.63 21.52 23.60 31.12 40.56 26.68 35.05 24.94 34.29 45.28 TFusionOcc (T-SQ-12800) Res-50 C+L 45.79 (-0.91%) 30.28 (-0.48%) 29.32 22.06 35.67 38.47 18.43 23.29 26.19 24.65 23.34 28.40 41.79 28.40 34.99 26.16 35.75 47.54 TFusionOcc (T-SQ-IW-12800) Res-50 C+L 45.81 (-1.11%) 30.74 (-0.47%) 30.54 21.85 37.99 39.11 20.43 21.56 26.44 24.82 23.39 28.74 41.99 28.68 35.70 26.86 36.08 47.63 TFusionOcc (T-P-25600) Res-50 C+L 44.77 (-0.72%) 29.22 (-0.82%) 27.86 21.11 36.24 38.33 17.79 18.33 25.25 23.53 25.50 28.59 40.27 24.95 33.54 24.82 34.90 46.57 TFusionOcc (T-SQ-25600) Res-50 C+L 45.40 (-0.74%) 30.38 (-0.50%) 29.18 22.52 36.87 38.61 19.41 22.19 26.98 24.84 24.78 28.70 41.21 27.39 34.74 25.57 35.69 47.44 TFusionOcc (T-SQ-IW-25600) Res-50 C+L 46.25 (-0.76%) 30.82 (-0.65%) 30.38 21.70 38.23 39.28 20.29 19.78 26.75 25.35 24.88 29.05 42.35 28.14 35.91 26.52 36.49 48.01
TABLE IV: 3D semantic occupancy prediction results on SurroundOcc-nuScenes validation rainy scenario subset. All methods are trained with dense occupancy labels from SurroundOcc [30]. Notion of modality: Camera (C), Lidar (L), Radar (R). The best and second-best performance is indicated in bold red and bold blue, respectively.
Method Backbone Modality IoU mIoU $\bullet$ barrier $\bullet$ bicycle $\bullet$ bus $\bullet$ car $\bullet$ const. veh. $\bullet$ motorcycle $\bullet$ pedestrian $\bullet$ traffic cone $\bullet$ trailer $\bullet$ truck $\bullet$ drive. surf. $\bullet$ other flat $\bullet$ sidewalk $\bullet$ terrain $\bullet$ manmade $\bullet$ vegetation InverseMatrixVT3D [17] R101-DCN C 22.48 (-7.55%) 9.99 (-8.89%) 10.40 12.03 0.00 29.94 0.00 9.92 4.88 0.91 0.00 17.79 29.10 2.37 10.80 9.40 8.68 13.57 GaussianFormer [9] R101-DCN C 20.30 (-9.53%) 9.07 (-10.03%) 6.11 8.70 0.00 25.75 0.00 10.44 2.85 0.55 0.00 17.26 30.65 2.95 12.53 9.94 6.65 10.71 Co-Occ [21] R101 C 18.90 (-11.10%) 9.40 (-10.90%) 4.50 9.30 0.00 29.50 0.00 8.40 3.50 0.00 0.00 15.10 29.40 0.60 12.40 11.50 10.70 15.50 GaussianFormer2-256 [7] R101-DCN C 21.19 (-10.55%) 10.14 (-10.68%) 5.25 9.29 0.00 29.33 0.00 13.65 5.80 0.90 0.00 20.22 31.80 1.94 14.83 10.48 5.96 12.72 SurroundOcc [30] R101-DCN C 24.38 (-7.11%) 10.80 (-9.50%) 10.55 14.60 0.00 31.05 0.00 8.26 5.37 0.58 0.00 18.75 30.72 2.74 12.39 11.53 10.52 15.77 Inverse++ [19] R101-DCN C 23.70 (-8.03%) 10.93 (-9.98%) 8.87 10.19 0.00 32.62 0.00 11.77 7.46 0.72 0.00 22.20 32.95 2.15 13.01 9.79 8.61 14.48 Co-Occ [21] R101 C+L 35.60 (-5.50%) 14.60 (-12.50%) 8.40 16.40 0.00 37.20 0.00 13.80 10.90 0.40 0.00 24.30 36.40 2.20 14.60 17.00 19.90 31.70 OccFusion [18] R101-DCN C+L 40.87 (-2.66%) 15.87 (-11.68%) 13.28 17.53 0.00 36.42 0.00 16.16 11.37 1.42 0.00 25.71 32.64 0.63 16.06 20.87 24.52 37.27 OccFusion [18] R101-DCN C+L+R 41.01 (-2.95%) 16.61 (-11.66%) 15.70 16.26 0.00 38.09 0.00 22.18 13.24 0.08 0.00 25.92 33.15 1.57 16.08 21.09 24.51 37.83 GaussianFusionOcc [23] R101-DCN C+L 42.78 (-2.38%) 18.66 (-11.55%) 16.09 12.27 0.00 39.82 0.00 27.66 13.68 0.07 0.00 38.25 40.10 2.07 19.64 19.82 29.55 39.61 GaussianFusionOcc [23] R101-DCN C+L+R 42.51 (-2.69%) 18.45 (-11.92%) 12.15 11.47 0.00 39.77 0.00 29.58 15.27 0.04 0.00 37.08 37.13 2.58 19.94 20.63 29.42 40.10 OccCylindrical [16] Res-50 C+L 43.38 (-1.56%) 17.79 (-10.88%) 16.19 10.04 0.00 37.84 0.00 25.63 13.28 0.23 0.00 32.93 39.59 3.89 17.89 21.61 26.99 38.47 DAOcc [33] Res-50 C+L 42.85 (-2.15%) 18.53 (-11.97%) 17.18 15.07 0.00 39.05 0.00 25.77 14.70 0.07 0.00 32.51 33.95 4.02 19.78 22.58 30.91 40.87 TFusionOcc (T-SQ-12800) Res-50 C+L 44.74 (-1.96%) 18.81 (-11.95%) 16.01 17.51 0.00 39.84 0.00 23.82 12.76 1.13 0.00 32.79 37.67 3.48 19.70 21.42 31.41 43.40 TFusionOcc (T-SQ-IW-12800) Res-50 C+L 44.82 (-2.10%) 19.07 (-12.14%) 16.81 15.53 0.00 40.36 0.00 21.27 13.33 3.79 0.00 34.89 38.04 3.92 19.86 21.89 31.67 43.72 TFusionOcc (T-P-25600) Res-50 C+L 43.25 (-2.24%) 17.83 (-12.21%) 14.12 10.10 0.00 38.45 0.00 21.99 13.10 1.16 0.00 32.05 34.60 4.61 18.91 23.14 30.76 42.35 TFusionOcc (T-SQ-25600) Res-50 C+L 43.86 (-2.28%) 18.95 (-11.93%) 17.10 16.09 0.00 40.17 0.00 23.21 15.13 2.97 0.00 33.83 36.18 3.75 20.12 20.48 31.30 42.94 TFusionOcc (T-SQ-IW-25600) Res-50 C+L 44.97 (-2.04%) 19.67 (-11.80%) 17.98 19.96 0.00 40.69 0.00 23.30 14.76 3.85 0.00 34.49 38.35 3.43 19.70 22.42 31.95 43.79
TABLE V: 3D semantic occupancy prediction results on SurroundOcc-nuScenes validation night scenario subset. All methods are trained with dense occupancy labels from SurroundOcc [30]. Notion of modality: Camera (C), Lidar (L), Radar (R). The best and second-best performance is indicated in bold red and bold blue, respectively.
IV-G Performance Analysis On nuScenes-C Dataset
To further assess the robustness of the proposed algorithm in extreme conditions, such as snow, fog, and sensor malfunctions, we use the nuScenes-C dataset and evaluate TFusionOcc at different levels of corruption and severity. The results of the experiment for several camera corruption settings are shown in Tables VI, VII, and X, and the results for several lidar corruption settings are shown in Tables VIII, IX, and XI. Note that since the nuScenes-C dataset only provides single-lidar-sweep corruption data, and our model leverages 10 lidar sweeps. Hence, when evaluating our model across different lidar corruption settings, we use only a single corrupted lidar sweep from the nuScenes-C dataset, and the model performance under the clean setting differs from that in Table II due to a much sparser lidar point cloud. In the camera corruption scenario, our model shows strong robustness against camera crashes, frame loss, brightness, and fog, as measured by the mIoU metric. Nevertheless, the motion blur and snow scenarios cause significant performance degradation in our model. With around 5.04% to 8.43% and 4.91% to 6.64% degradation under the IoU metric, 10.00% to 10.63% and 14.02% to 14.55% degradation under the mIoU metric for the motion blur and snow corruption case, respectively. In the lidar corruption scenario, our model is sensitive to lidar point cloud density and quality degradation, demonstrating its firm reliance on lidar point cloud quality.
Method Metric Clean Cam Crash Frame Lost Color Quant Motion Blur Bright Low Light Fog Snow OccCylindrical IoU 44.94 41.76 (-3.18%) 40.84 (-4.10%) 43.64 (-1.30%) 41.45 (-3.49%) 44.78 (-0.16%) 43.48 (-1.46%) 44.59 (-0.35%) 38.92 (-6.02%) DAOcc IoU 45.00 42.75 (-2.25%) 42.16 (-2.84%) 43.88 (-1.12%) 42.75 (-2.25%) 43.72 (-1.28%) 41.25 (-3.75%) 43.92 (-1.08%) 40.72 (-4.28%) TFusionOcc (T-P-25600) IoU 45.49 42.59 (-2.90%) 41.97 (-3.52%) 43.14 (-2.35%) 40.45 (-5.04%) 44.81 (-0.68%) 41.24 (-4.25%) 44.14 (-1.35%) 40.58 (-4.91%) TFusionOcc (T-SQ-25600) IoU 46.14 42.59 (-3.55%) 41.93 (-4.21%) 43.03 (-3.11%) 38.36 (-7.78%) 45.41 (-0.73%) 41.99 (-4.15%) 44.90 (-1.24%) 40.09 (-6.05%) TFusionOcc (T-SQ-IW-25600) IoU 47.01 42.61 (-4.40%) 41.77 (-5.24%) 43.97 (-3.04%) 38.58 (-8.43%) 46.31 (-0.70%) 42.19 (-4.82%) 45.95 (-1.06%) 40.37 (-6.64%)
TABLE VI: 3D semantic occupancy prediction IoU results on nuScenes-C Camera Corruption benchmark. All methods are evaluated with different corruption type input images provided by nuScenes-C [31] and ground-truth occupancy labels from SurroundOcc [30].
Method Metric Clean Cam Crash Frame Lost Color Quant Motion Blur Bright Low Light Fog Snow SurroundOcc mIoU 20.30 11.60 (-8.70%) 10.00 (-10.30%) 14.03 (-6.27%) 12.41 (-7.89%) 19.18 (-1.12%) 12.15 (-8.15%) 18.42 (-1.88%) 7.39 (-12.91%) Inverse++ mIoU 20.91 11.76 (-9.15%) 10.22 (-10.69%) 14.88 (-6.03%) 12.84 (-8.07%) 19.71 (-1.20%) 12.05 (-8.86%) 18.86 (-2.05%) 7.62 (-13.29%) OccCylindrical mIoU 28.67 23.38 (-5.29%) 21.87 (-6.80%) 25.38 (-3.29%) 22.57 (-6.10%) 27.88 (-0.79%) 22.17 (-6.50%) 27.38 (-1.29%) 18.29 (-10.38%) DAOcc mIoU 30.50 22.35 (-8.15%) 20.78 (-9.72%) 27.46 (-3.04%) 26.25 (-4.25%) 26.78 (-3.72%) 21.15 (-9.35%) 26.87 (-3.63%) 19.42 (-11.08%) TFusionOcc (T-P-25600) mIoU 30.04 22.34 (-7.70%) 20.78 (-9.26%) 24.42 (-5.62%) 20.04 (-10.00%) 29.03 (-1.01%) 18.62 (-11.42%) 27.43 (-2.61%) 15.62 (-14.42%) TFusionOcc (T-SQ-25600) mIoU 30.88 23.50 (-7.38%) 22.04 (-8.84%) 25.65 (-5.23%) 20.25 (-10.63%) 29.92 (-0.96%) 20.98 (-9.90%) 28.85 (-2.03%) 16.33 (-14.55%) TFusionOcc (T-SQ-IW-25600) mIoU 31.47 24.39 (-7.08%) 23.03 (-8.44%) 26.14 (-5.33%) 21.06 (-10.41%) 30.58 (-0.89%) 21.67 (-9.80%) 29.69 (-1.78%) 17.45 (-14.02%)
TABLE VII: 3D semantic occupancy prediction mIoU results on nuScenes-C Camera Corruption benchmark. All methods are evaluated with different corruption type input images provided by nuScenes-C [31] and ground-truth occupancy labels from SurroundOcc [30].
Method Metric Clean Beam Missing Cross Sensor Cross Talk Incomplete Echo Motion Blur Wet Ground Fog Snow OccCylindrical IoU 35.89 30.73 (-5.16%) 25.01 (-10.88%) 33.29 (-2.60%) 35.37 (-0.52%) 32.44 (-3.45%) 35.04 (-0.85%) 25.79 (-10.10%) 34.00 (-1.89%) DAOcc IoU 42.39 38.17 (-4.22%) 33.79 (-8.60%) 33.68 (-8.71%) 42.01 (-0.38%) 38.05 (-4.34%) 41.67 (-0.72%) 31.47 (-10.92%) 37.01 (-5.38%) TFusionOcc (T-P-25600) IoU 40.29 32.75 (-7.54%) 25.65 (-14.64%) 36.25 (-4.04%) 39.67 (-0.62%) 34.79 (-5.50%) 38.17 (-2.12%) 26.44 (-13.85%) 36.88 (-3.41%) TFusionOcc (T-SQ-25600) IoU 40.53 33.00 (-7.53%) 25.89 (-14.64%) 36.60 (-3.93%) 39.90 (-0.63%) 34.67 (-5.86%) 38.74 (-1.79%) 26.54 (-13.99%) 36.88 (-3.65%) TFusionOcc (T-SQ-IW-25600) IoU 41.53 34.09 (-7.44%) 26.82 (-14.71%) 37.50 (-4.03%) 40.94 (-0.59%) 35.54 (-5.99%) 39.84 (-1.69%) 27.10 (-14.43%) 37.70 (-3.83%)
TABLE VIII: 3D semantic occupancy prediction IoU results on nuScenes-C Lidar Corruption benchmark. All methods are evaluated with different corruption type input images provided by nuScenes-C [31] and ground-truth occupancy labels from SurroundOcc [30].
Method Metric Clean Beam Missing Cross Sensor Cross Talk Incomplete Echo Motion Blur Wet Ground Fog Snow OccCylindrical mIoU 24.06 19.83 (-4.23%) 16.20 (-7.86%) 22.60 (-1.46%) 20.93 (-3.13%) 20.85 (-3.21%) 23.31 (-0.75%) 17.32 (-6.74%) 22.38 (-1.68%) DAOcc mIoU 27.73 22.96 (-4.77%) 18.41 (-9.32%) 22.86 (-4.87%) 24.45 (-3.28%) 23.66 (-4.07%) 27.25 (-0.48%) 19.73 (-8.00%) 24.53 (-3.20%) TFusionOcc (T-P-25600) mIoU 26.13 20.49 (-5.64%) 15.38 (-10.75%) 23.76 (-2.37%) 22.69 (-3.44%) 21.39 (-4.74%) 25.05 (-1.08%) 17.44 (-8.69%) 23.15 (-2.98%) TFusionOcc (T-SQ-25600) mIoU 27.03 21.07 (-5.96%) 15.61 (-11.42%) 24.73 (-2.30%) 23.30 (-3.73%) 22.15 (-4.88%) 26.06 (-0.97%) 18.40 (-8.63%) 24.37 (-2.66%) TFusionOcc (T-SQ-IW-25600) mIoU 27.50 21.81 (-5.69%) 16.49 (-11.01%) 24.94 (-2.56%) 24.11 (-3.39%) 22.62 (-4.88%) 26.57 (-0.93%) 18.63 (-8.87%) 24.37 (-3.13%)
TABLE IX: 3D semantic occupancy prediction mIoU results on nuScenes-C Lidar Corruption benchmark. All methods are evaluated with different corruption type input lidar point cloud provided by nuScenes-C [31] and ground-truth occupancy labels from SurroundOcc [30].
Corruption Severity IoU mIoU $\bullet$ barrier $\bullet$ bicycle $\bullet$ bus $\bullet$ car $\bullet$ const. veh. $\bullet$ motorcycle $\bullet$ pedestrian $\bullet$ traffic cone $\bullet$ trailer $\bullet$ truck $\bullet$ drive. surf. $\bullet$ other flat $\bullet$ sidewalk $\bullet$ terrain $\bullet$ manmade $\bullet$ vegetation CameraCrash Easy 43.16 26.44 23.34 16.62 28.80 35.79 19.34 21.30 28.47 19.01 18.46 24.14 36.89 19.19 25.07 27.42 35.32 43.94 CameraCrash Mid 42.27 23.37 17.66 11.56 26.79 33.91 14.58 16.88 26.48 16.44 13.69 19.75 35.87 16.54 22.43 24.70 34.07 42.56 CameraCrash Hard 42.39 23.35 20.30 11.23 26.54 34.40 15.27 13.93 25.80 17.87 11.15 21.39 35.24 14.59 22.44 25.18 34.52 43.77 FrameLost Easy 44.33 27.42 25.40 16.12 30.44 36.89 19.15 21.10 28.59 20.72 18.97 26.33 38.65 21.48 26.20 28.11 35.81 44.84 FrameLost Mid 41.33 22.19 15.55 10.88 24.67 33.26 13.95 14.55 25.62 15.66 11.70 19.94 33.63 13.21 21.59 24.17 33.98 42.73 FrameLost Hard 39.66 19.48 9.91 9.09 22.02 31.26 11.74 11.63 23.98 12.70 8.77 15.30 30.99 9.11 18.80 21.75 33.03 41.52 ColorQuant Easy 46.55 30.61 31.18 19.09 33.46 39.63 22.86 26.15 30.57 23.20 22.87 30.29 42.26 25.40 28.71 30.42 37.19 46.44 ColorQuant Mid 44.19 26.96 25.94 14.03 30.21 38.26 19.43 18.13 29.03 22.00 18.56 26.63 38.54 19.68 24.39 26.26 35.72 44.54 ColorQuant Hard 41.18 20.85 13.87 8.24 22.61 34.82 12.93 8.22 26.70 18.66 9.86 20.29 33.70 9.37 19.14 33.79 42.30 20.85 MotionBlur Easy 41.43 26.89 28.45 15.39 31.42 37.33 20.59 19.98 29.74 23.25 19.18 25.99 36.16 17.03 22.63 25.86 34.73 42.48 MotionBlur Mid 37.26 18.80 15.08 8.51 21.43 33.12 11.83 13.43 22.58 17.93 6.36 13.80 29.91 3.61 12.99 20.31 31.56 38.34 MotionBlur Hard 37.06 17.48 12.18 7.54 19.65 32.08 9.80 12.42 20.60 16.37 4.50 12.63 29.47 2.63 11.26 19.46 31.17 37.93 Brightness Easy 46.90 31.35 31.83 20.44 34.29 39.73 23.55 27.70 31.19 23.89 23.07 30.52 43.06 27.37 29.94 30.92 37.45 46.57 Brightness Mid 46.45 30.67 31.12 19.86 32.69 39.37 22.29 26.38 31.08 23.73 22.70 29.40 42.35 27.07 28.96 30.20 37.26 46.31 Brightness Hard 45.59 29.71 30.29 19.08 31.22 38.80 20.78 25.06 30.92 23.76 21.97 28.14 40.90 25.49 27.29 28.93 36.92 45.88 LowLight Easy 43.86 25.38 19.52 14.11 30.88 37.84 16.55 20.19 28.43 17.38 13.94 25.64 37.39 15.11 23.22 25.82 35.57 44.48 LowLight Mid 42.12 21.69 11.25 11.54 27.65 36.35 11.88 18.08 26.75 12.07 5.65 22.09 34.43 8.43 19.90 23.02 34.45 43.44 LowLight Hard 40.60 17.93 3.52 6.49 23.37 34.05 7.87 15.49 23.81 5.33 1.57 17.02 32.01 3.56 16.69 20.34 33.18 42.59 Fog Easy 46.37 30.57 30.82 19.55 34.30 39.40 21.91 25.99 30.65 23.35 22.12 30.06 42.13 26.77 28.91 29.86 37.10 46.27 Fog Mid 46.03 29.89 30.01 18.93 33.72 39.29 20.89 25.12 30.42 22.91 21.21 29.32 41.37 25.36 27.89 28.93 36.87 46.07 Fog Hard 45.44 28.61 29.00 17.43 32.02 38.90 18.92 24.15 30.04 22.41 19.61 27.90 40.02 22.29 25.69 27.37 36.47 45.58 Snow Easy 41.26 20.46 12.34 8.18 24.01 34.48 14.53 13.38 26.04 20.12 6.13 19.37 32.52 6.58 12.73 21.68 33.02 42.31 Snow Mid 39.61 15.81 7.13 4.17 16.34 30.79 7.63 11.86 22.00 16.87 4.93 9.97 27.63 2.30 7.69 15.13 30.04 38.46 Snow Hard 40.23 16.08 5.29 4.19 16.00 31.17 8.68 11.28 22.39 16.59 4.58 10.08 29.80 2.64 9.17 14.27 31.40 39.78
TABLE X: The performance of TFusionOcc(T-SQ-IW-25600) on the nuScene-C [31] dataset under different camera corruption and severity settings.
Corruption Severity IoU mIoU $\bullet$ barrier $\bullet$ bicycle $\bullet$ bus $\bullet$ car $\bullet$ const. veh. $\bullet$ motorcycle $\bullet$ pedestrian $\bullet$ traffic cone $\bullet$ trailer $\bullet$ truck $\bullet$ drive. surf. $\bullet$ other flat $\bullet$ sidewalk $\bullet$ terrain $\bullet$ manmade $\bullet$ vegetation BeamMissing Light 38.03 24.83 25.39 12.35 30.38 33.20 17.80 20.84 23.97 16.98 19.35 26.19 37.83 21.87 23.66 24.77 26.92 35.79 BeamMissing Moderate 34.28 21.94 22.45 9.84 28.26 29.26 15.39 17.79 20.83 12.97 17.61 23.96 35.44 19.69 20.99 22.06 23.07 31.41 BeamMissing Heavy 29.97 18.67 18.85 6.93 25.51 24.95 13.02 13.23 15.80 11.40 15.58 21.18 32.63 17.35 18.16 18.45 19.03 26.71 CrossSensor Light 31.71 20.08 19.55 7.95 27.85 27.18 14.63 13.75 17.86 11.99 17.34 23.06 34.50 19.05 19.57 19.78 20.08 27.19 CrossSensor Moderate 26.52 16.07 15.50 5.00 23.42 21.71 12.15 7.42 13.75 9.30 14.33 19.55 31.52 15.85 16.60 16.43 14.97 19.66 CrossSensor Heavy 22.23 13.32 11.50 3.77 21.31 16.56 6.96 9.04 10.43 7.01 13.48 15.76 28.99 14.64 14.00 13.66 11.28 14.64 CrossTalk Light 39.77 26.33 28.50 14.08 31.14 35.44 18.51 21.00 25.73 17.29 19.97 27.30 39.52 24.07 26.08 27.46 29.48 35.63 CrossTalk Moderate 37.47 24.93 27.82 12.62 30.10 34.11 16.63 19.01 24.07 15.91 19.14 26.14 38.85 23.78 25.07 26.82 28.03 30.82 CrossTalk Heavy 35.26 23.57 26.75 11.55 29.19 32.80 15.39 16.61 22.60 14.60 18.47 25.09 38.09 23.34 23.41 25.86 26.44 26.93 IncompleteEcho Light 41.18 25.74 28.97 12.08 28.39 30.45 15.70 19.21 27.68 19.35 17.63 24.15 39.80 23.88 26.37 27.75 30.51 39.94 IncompleteEcho Moderate 41.01 24.66 28.96 10.54 25.72 26.81 13.15 17.27 27.67 19.26 15.40 21.58 39.74 23.90 26.36 27.73 30.52 39.93 IncompleteEcho Heavy 40.62 21.93 28.92 8.66 17.25 16.89 7.41 13.52 27.62 19.08 8.97 14.64 39.57 23.92 26.34 27.71 30.53 39.91 MotionBlur Light 38.10 24.89 25.61 12.99 30.89 33.74 19.39 20.91 24.15 14.99 19.40 26.77 36.30 21.25 23.10 24.59 26.09 38.00 MotionBlur Moderate 35.44 22.56 22.40 11.01 29.50 31.09 18.24 18.10 20.87 11.53 18.13 25.43 33.88 18.98 20.68 22.26 22.74 36.10 MotionBlur Heavy 33.07 20.42 19.24 9.15 28.01 28.37 17.26 15.57 17.81 8.80 16.49 23.64 32.07 17.44 18.77 20.20 19.85 34.03 WetGround Light 40.53 26.99 28.86 15.18 31.66 36.33 19.80 23.18 27.60 19.65 20.73 28.03 37.07 21.62 23.82 27.84 30.50 39.99 WetGround Moderate 39.84 26.57 28.86 15.12 31.65 36.23 19.77 22.70 27.53 19.86 20.71 27.91 35.13 19.21 21.98 28.01 30.50 39.98 WetGround Heavy 39.14 26.14 28.80 14.79 31.57 36.14 19.74 22.69 27.47 19.87 20.68 27.85 33.08 17.10 20.01 27.99 30.49 39.98 Fog Light 36.86 24.38 27.36 14.09 28.96 33.78 15.06 20.04 23.85 18.70 17.81 23.49 38.44 22.32 24.71 23.95 25.84 31.69 Fog Moderate 28.39 20.05 22.93 12.52 26.29 29.36 12.23 17.41 19.34 16.90 11.32 16.80 34.83 20.18 21.50 19.90 16.92 22.37 Fog Heavy 16.04 11.46 13.35 6.10 15.92 19.10 7.93 8.75 9.78 10.17 3.34 7.81 26.48 14.66 12.86 11.56 6.17 9.35 Snow Light 39.38 24.78 28.29 14.52 29.78 34.99 14.71 19.91 23.43 17.00 18.80 25.34 39.33 23.29 25.23 26.10 27.60 28.23 Snow Moderate 37.52 24.42 28.34 14.26 29.77 34.75 14.91 19.96 22.36 15.96 18.51 24.98 38.93 23.35 25.34 26.44 27.15 25.73 Snow Heavy 36.20 23.91 28.11 13.47 29.74 34.20 15.34 18.98 21.07 13.80 18.17 25.16 38.42 23.08 25.28 26.47 25.68 25.65
TABLE XI: The performance of TFusionOcc(T-SQ-IW-25600) on the nuScene-C [31] dataset under different lidar corruption and severity settings.
IV-H Performance Analysis On Varying Distance
IV-H1 Performance Variation Trend At Different Distances.
Following [18], we take the ego vehicle as the origin and $R$ as the radius. By adjusting the length of $R$, we evaluate our proposed method against other SOTA algorithms at varying perception distance radius [20m, 25m, 30m, 35m, 40m, 45m, 50m], including challenging rainy and night scenarios. The trend of performance variation is shown in Figure 10.
The trend of performance variation across the whole validation set, as shown in Figures 10(a) and 10(d), aligns with the performance benchmark in Table II. In the rainy scenario, the performance variation trend of our proposed method is similar to the results of the whole validation set experiment, except that the performance gap between different T-Primitive settings further narrows down to 0.11% to 2.90% under the mIoU metric and 0.44% to 1.74% under the IoU metric due to the semantic and geometric information loss from both sensors. In the nighttime scenario, we discover that our T-SQ-25600 setting outperforms the second-best T-SQ-IW-12800 setting within a 35-meter perception range under the mIoU metric, as shown in Figure 10(c). This phenomenon may stem from the fact that more primitives can better capture close-range details, leading to better robustness against challenging dim-lighting conditions.
Figure 10: Performance variation trend for 3D semantic occupancy prediction task. (a) mIoU performance variation trend on the whole SurroundOcc-nuScenes validation set, (b) IoU performance variation on the whole SurroundOcc-nuScenes validation set, (c) mIoU performance variation trend on the SurroundOcc-nuScenes validation rainy scenario subset, (d) IoU performance variation on the SurroundOcc-nuScenes validation rainy scenario subset, (e) mIoU performance variation on the SurroundOcc-nuScenes validation night scenario subset, and (f) IoU performance variation on the SurroundOcc-nuScenes validation night scenario subset. Better viewed when zoomed in.
IV-H2 Performance In Different Sectors Ranges
[…]
Figure 11: Performance at different sector ranges for the 3D semantic occupancy prediction task. (a) mIoU performance at different sector ranges on the whole SurroundOcc-nuScenes validation set, (b) IoU performance at different sector ranges on the whole SurroundOcc-nuScenes validation set, (c) mIoU performance at different sector ranges on the SurroundOcc-nuScenes validation rainy scenario subset, (d) IoU performance at different sector ranges on the SurroundOcc-nuScenes validation rainy scenario subset, (e) mIoU performance at different sector ranges on the SurroundOcc-nuScenes validation night scenario subset, and (f) IoU performance at different sector ranges on the SurroundOcc-nuScenes validation night scenario subset. Better viewed when zoomed in.
IV-I Qualitative Study
IV-I1 Qualitative Results On SurroundOcc-nuScenes Dataset
To qualitatively study the difference in each setting of the proposed TFusionOcc, we demonstrate the visualization result of each setting in daytime, rainy, and night-time scenarios in Figure 12, with the main discrepancy area highlighted. In the daytime scenarios as shown at the top of Figure 12, in the orange square highlighted area, except for the T-P-25600 setting, all other settings successfully detect those two pedestrians. What is more, T-SQ-IW Primitive preserves better the geometric contour shape of background buildings, such as sharper building edges and more accurate building boundaries, than T-P primitive and T-SQ primitive, demonstrating the importance of the geometric shape flexibility of intermediate primitives. In rainy scenarios, as shown in the middle of Figure 12, in the dark red square highlighted area. The better geometric shape representation capability of the primitives could more accurately capture the geometric structure of the object, leading to better detection performance; T-SQ-IW-25600 achieved the best performance due to its best geometric shape representation capability, and T-P-25600 the worst, stemming from the ellipsoid geometric shape prior. Similarly, in the nighttime scenarios, as shown in the bottom of Figure 12, in the square highlighted area. Even with severe occlusions and a vehicle headlight glare pollution problem, we observed that greater geometric shape flexibility of primitives leads to better detection performance, which aligns with rainy and daytime phenomena.
Figure 12: Qualitative results under different settings for daytime, rainy, and nighttime scenarios displayed in the upper, middle, and bottom sections, respectively. Better viewed when zoomed in.
Furthermore, to demonstrate the effectiveness and robustness of our method, we compare TFusionOcc in three different T-Primitive settings, such as T-P-25600, T-SQ-25600 and T-SQ-IW-25600, against other SOTA algorithms in challenging daytime, rainy and nighttime scenarios, as shown in Figure 13, with the primary difference area highlighted in each scenario. In the daytime scenario, as shown at the top of Figure 13, in the orange square highlighted area, benefit from multi-task learning, which provides an additional supervision signal, DAOcc successfully detects distant pedestrians and achieves similar performance as TFusionOcc. However, in the dark-red-square-highlighted area, DAOcc fails to detect the distant vehicle, demonstrating the superior performance of our proposed method. In the rainy scenario, as shown in the middle of Figure 13, in the purple square highlighted area, our proposed method not only successfully detects all dynamic objects, but also preserves their general contours well. In the highlighted dark-blue-square area, our approach achieves par performance with other SOTA algorithms for detecting challenging pedestrians. In the nighttime scenario, as shown at the bottom of Figure 13, our proposed method successfully detects very distant dynamic objects in the highlighted dark-red-square area, demonstrating its superior long-range dynamic object perception capability.
Figure 13: Qualitative results compare against other SOTA algorithms for daytime, rainy, and nighttime scenarios displayed in the upper, middle, and bottom sections, respectively. Better viewed when zoomed in.
IV-I2 Qualitative Results On nuScenes-C Dataset
Figure 14: Qualitative results compare against other SOTA algorithms under fog and snow scenarios, displayed in the top and bottom sections, respectively. Better viewed when zoomed in.
Since the nuScenes dataset does not contain fog and snow scenarios, which are common driving conditions in the real world, we adopted the nuScenes-C dataset to evaluate our model across three settings within the same scene under fog and snow conditions, with the severity of each condition set to hard, and compare the visualization results with other SOTA algorithms. The visualization results are shown in Figure 14, with the main area of the discrepancy highlighted in dark blue and dark red squares.
Under foggy conditions, the surround-view images suffer significant visibility degradation due to strong atmospheric scattering, leading to reduced illumination, low contrast, and severely blurred object boundaries, especially in distant regions of the scene. Compared to baseline SOTA algorithms, our method, especially the T-SQIW-25600 setting, successfully detects all dynamic pedestrian objects in the scene (as shown by the dark blue squares) and also accurately captures the background objects’ contour as well (as shown by the dark red squares), demonstrating the robustness of the proposed algorithm under foggy conditions. Nevertheless, in the snow condition, the snowfall captured in the surround-view image appears to introduce salt-and-pepper noise to the image, and snowflakes occlude some objects’ contours, even the main body parts. In this case, our proposed method is less tolerant of salt-and-pepper noise than the baseline SOTA algorithms.
IV-J Model Efficiency
Method Modality Latency (ms) ($\downarrow$) Memory (GB) ($\downarrow$) Params SurroundOcc* [30] C 472 5.98 180.51M InverseMatrixVT3D* [17] C 447 4.41 67.18M OccFusion(C+R)* [18] C+R 588 5.56 92.71M OccFusion(C+L)* [18] C+L 591 5.56 92.71M OccFusion(C+L+R)* [18] C+L+R 601 5.78 114.97M OccCylindrical* [16] C+L 398 10.63 111.62M TFusionOcc (T-SQ-12800) C+L 278.46 3.68 95.62M TFusionOcc (T-SQ-IW-12800) C+L 281.47 3.88 95.96M TFusionOcc (T-P-25600) C+L 385.49 4.73 97.51M TFusionOcc (T-SQ-25600) C+L 388.05 4.13 97.59M TFusionOcc (T-SQ-IW-25600) C+L 391.06 4.04 98.24M
TABLE XII: Efficiency comparison of multi-modal 3D semantic occupancy prediction on the nuScenes validation set. Algorithm results marked with * were taken from the paper that introduced the model and were measured on a different GPU.
We evaluated the real-time performance of our model in various settings on an RTX 4090 GPU and presented the results in Table XII. The experiment results show the trade-off between the total number of primitives and the model inference speed. More primitives require more rendering operations, leading to higher latency. Surprisingly, all five settings have similar total model parameters. Based on the main performance benchmark shown in Table II, the T-SQ-IW-12800 setting delivers the best balance between detection and real-time performance.
IV-K Ablation Study
IV-K1 Ablation Study On Skeleton Merge Module
Main-Skele Augment Overlap Remove Range Filter IoU ($\uparrow$) mIoU ($\uparrow$) 46.24 (Baseline) 30.79 (Baseline) 46.20 (-0.04%) 30.70 (-0.09%) 46.07 (-0.17%) 30.76 (-0.03%) 46.08 (-0.16%) 30.69 (-0.1%) 41.96 (-4.28%) 29.92 (-0.87%) 29.35 (-16.89%) 22.31 (-8.48%)
TABLE XIII: Ablation study on the Skeleton Merge Module. Main-Skele: Lidar-side main skeleton branch. Augment: Camera-side augment skeleton branch. $\uparrow$:the higher, the better.
Range Filter $r$ IoU ($\uparrow$) mIoU ($\uparrow$) 5 m 46.24 (Baseline) 30.79 (Baseline) 4 m 46.21 (-0.03%) 30.75 (-0.04%) 3 m 46.09 (-0.15%) 30.76 (-0.03%) 2 m 45.72 (-0.52%) 30.69 (-0.10%) 1 m 44.64 (-1.60%) 30.32 (-0.47%) 0.5 m 43.60 (-2.64%) 30.04 (-0.75%)
TABLE XIV: Ablation study on range filter hyperparameter $r$. $\uparrow$:the higher, the better.
We conducted an ablation study for each component of the skeleton merge module and demonstrated the results of the experiment in Table XIII. The experiment results prove the significant impact of the main skeleton structure, which supports the fundamental 29.92% mIoU and 41.96% IoU performance, and the auxiliary impact of the augment structure, which further boosts 0.87% mIoU and 4.28% IoU performance. The main skeleton, based on lidar, provides detailed 3D geometric information of the surrounding scene, and the augmented skeleton, based on the camera, provides additional information to mitigate occlusion impact in the scene. Furthermore, we studied the impact of the range filter hyperparameter $r$ on the overall performance of the model and present the results in Table XIV. The results show that a smaller filter range forces the augmented skeleton to get too close to the main skeleton, preventing it from overcoming the lidar sensor’s occlusion limitation.
IV-K2 Ablation Study On Depth-Fusion-Based 3D Deformable Attention Module
Vis-Depth Fusion Depth-Fusion Vis-Depth Lidar-Depth IoU ($\uparrow$) mIoU ($\uparrow$) 46.24 (Baseline) 30.79 (Baseline) 46.06 (-0.18%) 30.57 (-0.22%) 39.80 (-6.44%) 25.28 (-5.51%) 44.01 (-2.23%) 28.98 (-1.81%) 43.99 (-2.25%) 28.93 (-1.86%) 40.33 (-5.91%) 24.71 (-6.08%)
TABLE XV: Ablation study on Depth-Fusion-Based 3D Deformable Attention Module. $\uparrow$:the higher, the better. Vis-Depth Fusion: add the fused depth feature back to the visual feature, resulting in a depth-aware visual feature. Depth-Fusion: fuse the visual depth feature with the lidar depth feature. Vis-Depth: visual depth feature. Lidar-Depth: lidar depth feature. $\uparrow$:the higher, the better.
We conducted an ablation study with respect to each component of the depth-fusion-based 3D deformable attention module and presented the experiment results in Table XV. In practice, Vis-Depth provides dense but less accurate depth maps, while Lidar-Depth provides sparse but more accurate depth maps. The experiment shows that dense depth maps contribute significantly to the performance, even though they are less accurate. Relying solely on sparse but accurate depth maps from lidar results in a significant performance degradation of 5.5%-6.0%. Furthermore, adding fused depth maps back to the original visual features results in depth-aware visual features contributing the second-largest performance improvement, around 2.2%. The experiment results demonstrate the necessity of multi-stage multi-modality feature fusion to boost model performance.
IV-K3 Ablation Study On MGCAFusion Module
Skip-Connect WSFusion GCFusion IoU ($\uparrow$) mIoU ($\uparrow$) 46.24 (Baseline) 30.79 (Baseline) 38.60 (-7.64%) 18.18 (-12.61%) 39.39 (-6.85%) 18.57 (-12.22%) 24.20 (-22.04%) 19.91 (-10.88%)
TABLE XVI: Ablation study on the MGCAFusion Module. Skip-Connet: skip-connection structure at the last feature fusion stage. WSFusion: weighted summation fusion part. GCFusion: gated concatenation fusion part. $\uparrow$:the higher, the better.
We conducted an ablation study for each major component of the MGCAFusion module and present the experiment results in Table XVI. The experiment results show that the weighted summation and gated concatenation fusion components contribute almost equally to the model’s final performance, with the former slightly more important. Missing each part will cause 12.22% to 12.61% mIoU and 6.85% to 7.64% IoU performance degradation. The skip-connection structure plays an important role during the feature fusion.
V Conclusion and Future work
This paper presents TFusionOcc, a novel Student’s t-distribution-based object-centric multi-stage multi-sensor fusion framework. Our framework in the sensor fusion part adopts a multi-stage feature-level fusion strategy, including an early-stage, a middle-stage, and a late-stage fusion, to better integrate complementary information across modalities and mitigate the drawbacks of each modality. We also adopt an object-centric modeling strategy by leveraging the Student’s t-distribution and the T-Mixture model as a probability kernel, combined with several typical primitives, such as the general T-Primitive and the superquadric, to achieve greater robustness against outliers. We further propose a primitive called the deformable superquadric primitive (superquadric with inverse-warp), which offers greater geometric shape representation flexibility, thereby boosting the model’s 3D geometric shape-capturing capability in the driving scene. The extensive experiments conducted on the nuScenes [2] and nuScenes-C [31] datasets, including challenging cases of rainy, night, fog, snow and sensor-malfunction, demonstrate that our proposed approach not only achieves SOTA performance but also exhibits superior robustness. In our future research endeavors, we will investigate how to utilize a much smaller set of primitives and fewer rendering operations to achieve similar performance.

### v2

@@
-TFusionOcc: Student’s t-Distribution Based Object-Centric Multi-Sensor Fusion Framework for 3D Occupancy Prediction
+TFusionOcc: T-Primitive Based Object-Centric Multi-Sensor Fusion Framework for 3D Occupancy Prediction
 Abstract
-3D semantic occupancy prediction enables autonomous vehicles (AVs) to perceive fine-grained geometric and semantic structure of their surroundings from onboard sensors, which is essential for safe decision-making and navigation.
-Recent models for 3D semantic occupancy prediction have successfully addressed the challenge of describing real-world objects with varied shapes and classes.
-However, the intermediate representations used by existing methods for 3D semantic occupancy prediction rely heavily on 3D voxel volumes or a set of 3D Gaussians, hindering the model’s ability to efficiently and effectively capture fine-grained geometric details in the 3D driving environment.
-This paper introduces TFusionOcc, a novel object-centric multi-sensor fusion framework for predicting 3D semantic occupancy.
-By leveraging multi-stage multi-sensor fusion, Student’s t-distribution, and the T-Mixture model (TMM), together with more geometrically flexible primitives, such as the deformable superquadric (superquadric with inverse warp), the proposed method achieved state-of-the-art (SOTA) performance on the nuScenes benchmark.
-In addition, extensive experiments were conducted on the nuScenes-C dataset to demonstrate the robustness of the proposed method in different camera and lidar corruption scenarios.
+The prediction of 3D semantic occupancy enables autonomous vehicles (AVs) to perceive the fine-grained geometric and semantic scene structure for safe navigation and decision-making.
+Existing methods mainly rely on either voxel-based representations, which incur redundant computation over empty regions, or on object-centric Gaussian primitives, which are limited in modeling complex, non-convex, and asymmetric structures.
+In this paper, we present TFusionOcc, a T-primitive-based object-centric multi-sensor fusion framework for 3D semantic occupancy prediction.
+Specifically, we introduce a family of Student’s t-distribution-based T-primitives, including the plain T-primitive, T-Superquadric and deformable T-Superquadric with inverse warping, where the deformable T-Superquadric serves as the key geometry-enhancing primitive.
+We further develop a unified probabilistic formulation based on the Student’s t-distribution and the T-mixture model (TMM) to jointly model occupancy and semantics, and design a tightly coupled multi-stage fusion architecture to effectively integrate camera and LiDAR cues.
+Extensive experiments on nuScenes show state-of-the-art performance, while additional evaluations on nuScenes-C demonstrate strong robustness under most corruption scenarios.
 The code will be available at: https://github.com/DanielMing123/TFusionOcc
@@
 I Introduction
-Accurate and efficient modeling of the scene around autonomous vehicles (AV) is essential for safe and collision-free navigation.
-As technology advances, the introduction of a 3D semantic occupancy representation has successfully addressed the limitations of traditional 3D object detection and 2D bird’s-eye-view (BEV) segmentation networks, particularly for detecting irregular objects and predicting occupied space in 3D.
-Thus, 3D semantic occupancy provides a suitable representation of the vehicle’s surrounding environment.
-This representation inherently ensures geometric consistency of the AV’s driving scene and can accurately describe occluded areas.
-In addition, a multi-sensor fusion technique for 3D semantic occupancy prediction further enhances an AVs’ ability to model the 3D world.
-Recent advances in voxel-based multi-sensor fusion methods [18, 21, 35, 16, 33, 4, 26] have achieved remarkable progress by integrating complementary information from different sensors, such as surround-view cameras, surround-view radars, and lidars.
-Demonstrating superior robustness and overall performance in several challenging light and weather conditions, these methods heavily rely on volumetric grid calculations (Fig 1, voxel-based approach), which hinders real-time application and makes them unsuitable for edge device deployment.
-Meanwhile, recent advances in object-centric multi-sensor fusion methods [38, 23] aim to address the limitations of the voxel-based approach by utilizing a set of 3D Gaussian primitives to only model the occupied area in the driving scene (Fig 1, 3D Gaussian-based approach), reducing the computational overhead caused by empty regions by concentrating representational capacity where it matters most.
-However, their modeling performance is inferior to that of voxel-based approaches.
-This performance gap is mainly due to the limited geometric shape representation capability of primitives and their weaker robustness to outliers.
-Figure 1: Pipeline of three approaches: Voxel-based approach (top-left), 3D-Gaussian-Primitive-based object-centric approach (top-right), and our approach (bottom-left).
-To address the aforementioned limitations, we propose TFusionOcc (Fig 1, our approach), a novel Student’s t-distribution-based object-centric multi-sensor fusion framework.
-We adopt a multi-stage feature-level fusion strategy, including early-stage, middle-stage, and late-stage fusion, to better integrate complementary information across modalities and mitigate the drawbacks of each modality.
-For object-centric modeling, we adopt a Student’s t-distribution and the T-Mixture model (TMM) as a probability kernel, combined with several types of primitives, such as the general T-Primitive, superquadric, and deformable superquadric (superquadric with inverse-warp), to achieve greater geometric shape representation flexibility and robustness against outliers.
-The extensive experiments conducted on nuScenes [2] and nuScenes-C [31] datasets demonstrate that our proposed approach not only achieves state-of-the-art (SOTA) performance but also exhibits superior robustness against several sensor corruptions.
+Accurate and efficient scene representation is essential for autonomous driving, since downstream planning and decision-making rely on a reliable understanding of both geometry and semantics.
+Compared with conventional 3D object detection and BEV segmentation, 3D semantic occupancy prediction provides a more complete and denser description of the surrounding environment, especially for irregular structures, free space, and partially occluded regions.
+Multi-sensor fusion further improves this representation by combining the complementary strengths of cameras and LiDAR.
+Figure 1: Pipeline of three approaches: Voxel-based approach (top), 3D-Gaussian-Primitive-based object-centric approach (middle), and our approach (bottom).
+Recent progress in multi-sensor 3D semantic occupancy prediction has been largely driven by voxel-based fusion methods [18, 21, 37, 16, 35, 4, 26], which integrate complementary signals from cameras and LiDAR.
+Although these methods have achieved strong performance, they generally rely on dense volumetric processing (Fig.
+1, voxel-based approach), leading to substantial redundant computation over empty regions that dominate outdoor driving scenes.
+This inefficiency has motivated the development of object-centric representations, where a compact set of learnable primitives is used to model occupied regions more selectively.
+Among object-centric approaches[39, 23], Gaussian-based primitives have recently shown promise for improving efficiency and compact representation (Fig 1, 3D Gaussian-based approach).
+However, their geometric modeling capability remains limited when faced with non-convex, asymmetric, and structurally irregular scene elements that are common in real driving environments.
+In addition, Gaussian-like formulations are less effective in handling outliers and long-tailed spatial distributions, which may reduce robustness under noisy or corrupted observations.
+As a result, existing object-centric occupancy frameworks still face a performance gap when compared with strong voxel-based methods, especially when an accurate geometric representation is required.
+To address these limitations, we propose TFusionOcc (Fig 1, our approach), a T-primitive-based object-centric multi-sensor fusion framework for 3D semantic occupancy prediction.
+Specifically, we introduce a family of Student’s t-distribution-based T-primitives, including the plain T-primitive, T-Superquadric and a deformable T-Superquadric with inverse warping, where the latter serves as the key geometry-enhancing primitive.
+We further develop a unified probabilistic formulation based on Student’s t-distribution and the T-mixture model (TMM) to jointly model occupancy and semantics, and design a compact multi-stage fusion pipeline to effectively integrate camera and LiDAR cues.
+In this way, TFusionOcc improves geometric representation while preserving the efficiency advantage of object-centric modeling.
+Extensive experiments on nuScenes [2] show that TFusionOcc achieves state-of-the-art (SOTA) overall performance and consistently benefits from stronger primitive geometry.
+In particular, our results indicate that the flexibility of primitive shapes is more decisive than simply increasing the primitive count, highlighting the importance of geometry-aware primitive design.
+Additional evaluations on nuScenes-C [32] in different corruption settings further demonstrate strong robustness in many challenging cases.
 The main contributions of this paper are summarized as follows:
-  • We propose TFusionOcc, a novel object-centric multi-stage multi-sensor fusion framework that leverages different types of T primitives to efficiently and effectively model the 3D driving scene for 3D semantic occupancy prediction.
-  • We propose a multi-gate cross-attention fusion module (MGCAFusion), which consists of a weighted summation and gated concatenation fusion modules, to combine the multi-modal features in an effective way.
-  • We propose a set of Student’s t-distribution-based primitives, including the general T-Primitive, the T-Superquadric primitive, and the T-Superquadric with inverse warp primitive, to effectively model 3D driving scenes.
-  • We thoroughly investigated the robustness of the proposed method with respect to the impact of different perception ranges and different lidar and camera corruption scenarios.
+  • We propose a family of Student’s t-distribution-based T-primitives for object-centric 3D occupancy prediction, including the plain T-primitive, T-Superquadric, and a deformable T-Superquadric with inverse warping.
+In particular, the deformable T-Superquadric significantly improves geometric modeling flexibility, enabling better representation of non-convex and asymmetric structures.
+  • We develop a unified probabilistic modeling framework based on Student’s t-distribution and the T-mixture model, through which both occupancy and semantics are jointly modeled.
+This design improves robustness to outliers and complex real-world noise while alleviating the limitations of Gaussian-based primitives in modeling long-tailed spatial distributions.
+  • We design a tightly coupled multi-sensor fusion pipeline to support effective T-primitive learning.
+Specifically, it includes fused-depth-guided visual feature lifting, transformer-based T-primitive-centric refinement, and the GCWSFusion module for adaptive cross-modal interaction, allowing T-primitives to better exploit cross-modal cues for 3D semantic occupancy prediction.
 The remainder of this paper is structured as follows.
@@
 The model exhibits strong robustness under challenging rain and night scenarios.
-In Co-Occ [21], the authors leverage the volume rendering technique, which originally derived from NeRF [15], to refine the intermediate 3D feature volumes.
+In Co-Occ [21], the authors leverage the volume rendering technique, which was originally derived from NeRF [15], to refine the intermediate 3D feature volumes.
 They also proposed a GSFusion module for efficient multimodal feature fusion, achieving SOTA performance.
+In FusionOcc [37], the authors proposed a cross-modality fusion module to generate a high-quality depth-aware visual feature, followed by a 2D to 3D view transformation, resulting in volumes of depth-aware visual features under the Cartesian coordinate system.
+The LiDAR branch conducts voxelization and voxel encoding, resulting in a LiDAR feature volume also under the Cartesian coordinate system.
+Different modality feature volumes were fused via feature channel concatenation and further refined by a 3D Encoder.
 In OccCylindrical [16], the authors proposed a multi-sensor fusion-based framework that uses a cylindrical partition to perform voxelisation of a LiDAR-captured 3D point cloud and a pseudo 3D point cloud predicted by surround-view cameras, resulting in better preservation of 3D geometric information.
 The feature fusion and further refinement operations are performed under the cylindrical coordinate system.
-In FusionOcc [35], the authors proposed a cross-modality fusion module to generate a high-quality depth-aware visual feature, followed by a 2D to 3D view transformation, resulting in a depth-aware visual feature volumes under the Cartesian coordinate system.
-The lidar branch conducts voxelization and voxel encoding, resulting in a lidar feature volume also under the Cartesian coordinate system.
-Different modality feature volumes were fused via feature channel concatenation and further refined by a 3D Encoder.
+In AdaptiveOcc [34], the author explored using an efficient octree data structure to preserve the fine-grained geometry of the 3D scene, achieving a good balance between computational complexity and prediction performance for the voxel-grid-based approach.
 In DAOcc [35], the authors leverage multi-task learning to introduce an additional 3D supervision signal to guide intermediate features, leading to robust generalization and overall SOTA performance.
-Similarly, in Inverse++ [19], the authors leverage multi-task learning by incorporating an extra auxiliary 3D object detection branch.
+Similarly, in Doracamom [40], the authors developed a joint 3D detection and occupancy prediction pipeline under a camera-radar fusion paradigm.
+By leveraging the strengths of multi-task learning and further incorporating temporal fusion, the proposed method achieved superior performance in both 3D object detection and 3D semantic occupancy prediction.
 In addition, SDGOcc [4] incorporates a multi-task head to introduce a 2D semantic mask into the model training procedure, leveraging an additional 2D semantic mask to enhance the precision and densification of depth maps obtained from LiDAR point cloud projection, thereby boosting the model’s overall performance.
@@
 Our study presents a novel object-centric approach, aiming to address the aforementioned constraints by leveraging T-Primitives to model only occupied space in driving scenarios.
-II-B Object-Centric Modelling Approach
-In the GaussianFormer series [9, 7], the authors proposed a 3D Gaussian-based object-centric modeling approach to improve computational efficiency.
-They further proposed a Gaussian-mixture model (GMM) to improve the model’s robustness.
-The 3D-to-3D Gaussian splatting operator, initially proposed in [10] for 3D-to-2D splatting, is further proposed to boost the model’s efficiency.
-In QuadricFormer [40], the authors proposed leveraging a superquadric primitive combined with a Gaussian probability kernel to achieve object-centric modeling and benefit from the geometric shape representation’s flexibility, thereby improving model performance.
-In GaussianFormer3D [38], the authors extend the GaussianFormer work into a multi-sensor fusion scenario.
-By utilizing a set of 3D Gaussian primitives as an intermediate feature representation, the authors designed a pipeline that enables each 3D Gaussian primitive to aggregate features from different modalities and fuse them to perform 3D primitive-to-3D voxel volume Gaussian splatting.
+II-B Object-Centric Modeling Approach
+In the GaussianFormer series [9, 7], the authors proposed an object-centric 3D Gaussian model together with a Gaussian mixture model (GMM) to improve computational efficiency and robustness.
+In GraphGSOcc [27], the authors combine the object-centric 3D Gaussian representation with a graph transformer to realize dynamic and static object decoupling modeling.
+Allow the model to optimize foreground dynamic objects and background static objects separately.
+The approach further incorporates a temporal fusion mechanism to achieve superior performance.
+In GaussianFormer3D [39], the authors extend the GaussianFormer work to a multi-sensor fusion scenario.
+Using a set of 3D Gaussian primitives as an intermediate feature representation, the authors designed a pipeline that enables each 3D Gaussian primitive to aggregate features from different modalities and fuse them to perform 3D primitive to-3D voxel volume Gaussian splatting.
 Benefiting from the object-centric modeling technique and multi-modal fusion, the model achieves significant performance and efficiency improvements.
-In GaussianFusionOcc [23], the authors proposed a GaussianFusionBlock module that performs efficient and high-quality feature aggregation and fusion across different modalities, allowing high-quality 3D Gaussian primitive modeling and leading to SOTA performance.
-Though our proposed approach belongs to the object-centric modeling category, the key difference of our work compared to previous ones is that we explored Student’s t-distribution and T-mixture model to enhance the model robustness.
-Furthermore, we proposed a multi-stage feature fusion pipeline to mitigate the drawbacks of each modality, resulting in better fused features.
-Lastly, we enhanced the primitive’s geometric flexibility to provide a better intermediate representation, thereby achieving SOTA performance.
+In GaussianFusionOcc [23], the authors proposed a GaussianFusionBlock module that performs efficient and high-quality feature aggregation and fusion across different modalities, allowing high-quality 3D Gaussian primitive modeling and leading to outstanding performance.
+Unlike previous object-centric methods, our work introduces a family of Student’s t-distribution-based T-primitives for 3D semantic occupancy prediction.
+In particular, we propose a deformable T-Superquadric with inverse warping, which substantially strengthens primitive-level geometric flexibility.
+We further formulate occupancy and semantics under a unified TMM framework and design a multi-stage camera–LiDAR fusion pipeline to support effective T-primitive refinement.
+In this way, our method aims to preserve the efficiency advantage of object-centric modeling while improving geometric representation capability and robustness.
 III TFusionOcc
@@
 where $F$ refers to a multi-sensor fusion framework, $P^{T}$ refers to a set of primitives based on T-distribution $P^{T}=\left\{p^{T}_{1},p^{T}_{2},...,p^{T}_{K}\right\}$, along with a set of query vectors $Q=\left\{q_{1},q_{2},...,q_{K}\right\}$, $P^{T}_{refined}$ refers to a set of T-distribution-based primitives refined through multi-sensor fusion framework, $N$ refers to the total number of surround-view images and $Splat\left(\cdot\right)$ refers to the splatting operation.
-In the final predicted 3D semantic occupancy $Occ\in R^{\left\{X\times Y\times Z\right\}}$, each grid is assigned a semantic property ranging from 0 to $C$, where $C$ refers to the total amount of semantic classes.
+In the final predicted 3D semantic occupancy $Occ\in R^{\left\{X\times Y\times Z\right\}}$, each grid is assigned a semantic property ranging from 0 to $C$, where $C$ refers to the total number of semantic classes.
 In our case, a class value of 0 corresponds to an empty grid.
@@
 Figure 2: Overall architecture of TFusionOcc.
-The pipeline comprises two different modality branches and a multi-stage feature fusion branch.
+The pipeline comprises a camera branch, a LiDAR branch, and a multi-stage fusion branch.
 The camera branch extracts multi-scale visual features and predicts a pseudo 3D point cloud from surround-view images.
@@
 The feature fusion branch adopts a multi-stage fusion strategy to merge all outputs from the two-modality branch and leverages a proposed transformer to refine the T-primitives property through fused features.
-The overall architecture is shown in Figure 2.
+The overall architecture is shown in Fig.
+2.
 For the camera branch, given $\left\{img^{1},img^{2},...,img^{N}\right\}$ surround-view images, a 2D encoder and a decoder are leveraged to extract $V^{Cam}=\left\{\left\{V_{n}^{l}\right\}_{n=1}^{N}\in\mathbb{R}^{C\times H_{l}\times W_{l}}\right\}_{l=1}^{L=3}$ multi-scale visual features.
@@
 The aggregated semantic information is further fused with the LiDAR feature $Vol^{Lidar}_{Cylin}$ through an early-fusion module, resulting in a semantic-aware LiDAR feature $Vol^{Lidar}_{sem}$.
-The projection operation begins with a transition from a cylindrical coordinate system to a Cartesian coordinate system, followed by a transition to a new coordinate frame.
 The $Vol^{Cam}_{Cylin}$ and $Vol^{Lidar}_{Cylin}$ are fused through the skeleton merge module, resulting in a fused voxel volume $Vol^{Fused}_{Cylin}$ under cylindrical coordinates.
@@
 Lastly, the refined set of T-Primitives performs a splatting operation, resulting in the final 3D semantic occupancy grid.
-III-B Image 2D Encoder and Decoder for Surround-View Images
-The purpose of the image encoder is to capture both the spatial and semantic features of the surround-view images.
-These features serve as the foundation for the subsequent feature fusion procedure, and lastly, the task of 3D semantic occupancy prediction.
-In our approach, we first use a 2D backbone network (e.g.
-ResNet50[5]) to extract visual features at multiple scales $V^{Cam}_{Encode}=\left\{\left\{V_{n}^{l}\right\}_{n=1}^{N}\in\mathbb{R}^{C_{l}\times H_{l}\times W_{l}}\right\}_{l=1}^{L=3}$.
-Subsequently, these features are fused using feature-pyramid networks (FPN), resulting in visual features $V^{Cam}=\left\{\left\{V_{n}^{l}\right\}_{n=1}^{N}\in\mathbb{R}^{C\times H_{l}\times W_{l}}\right\}_{l=1}^{L=3}$ that have the same feature channel dimmension and resolutions that are $\frac{1}{8}$, $\frac{1}{16}$, and $\frac{1}{32}$ of the input image resolution, respectively.
-The deeper visual feature, with a smaller resolution, contains more semantic information.
-In contrast, the relatively shallower visual feature, with larger resolutions, provides richer spatial details.
-III-C Camera-based Pseudo 3D Point Cloud and Multi-Scale Cam Dense Depth Maps Generation
+III-B Camera-based Pseudo 3D Point Cloud and Multi-Scale Cam Dense Depth Maps Generation
 Figure 3: Inner Structure of DepthNet.
@@
 Meanwhile, the image-based pseudo-point cloud is generated solely from a 1/8-scale depth map.
-The pseudo 3D point cloud $Pt^{cam}$ and camera-based multi-scale dense depth maps $V^{cam}_{depth}$ are generated through the DepthNet module as shown in Figure 3.
+The pseudo 3D point cloud $Pt^{cam}$ and camera-based multi-scale dense depth maps $V^{cam}_{depth}$ are generated through the DepthNet module as shown in Fig.
+3.
 The DepthNet module takes $V^{Cam}_{Encode}$ as input and leverages a SECONDFPN [33] to generate a single-scale visual feature $V^{cam}_{sec}=\left\{V_{n}\right\}_{n=1}^{N}\in\mathbb{R}^{C\times H_{l}\times W_{l}}$, whose resolution is $\frac{1}{8}$ of the input image.
-The $V^{cam}_{sec}$ is further passed to an MLP to estimate the ray depth per-pixel, resulting in a single-scale depth map $V_{depth\_n}\in\mathbb{R}^{N\times D\times H_{l}\times W_{l}}$.
-Bilinear interpolation with a 0.5 downsampling factor is applied iteratively to $V_{depth\_n}$, resulting in two additional depth maps that have $\frac{1}{16}$ and $\frac{1}{32}$ resolution of the input image.
-Instead of performing depth estimation on each scale visual feature, our approach saves a certain amount of computation overhead while still maintaining a relatively high depth estimation precision.
+The $V^{cam}_{sec}$ is further passed to an MLP to perform the multi-hot per-ray depth estimation, resulting in a depth map $V_{depth\_1}\in\mathbb{R}^{N\times D\times H_{l}\times W_{l}}$ with each pixel ray having a set of depth bins separated by 0.5m intervals, and each depth bin on the pixel ray has a probability range in $[0,1]$.
+Following [7], we adopt the 3D semantic occupancy label to supervise $V_{depth\_1}$.
+Then, bilinear interpolation with a 0.5 downsampling factor is applied iteratively to $V_{depth\_1}$, resulting in two additional depth maps that have $\frac{1}{16}$ and $\frac{1}{32}$ resolution of the input image.
 The predicted depth map and the two extra downsampled depth maps form $V_{depth}^{cam}=\left\{\left\{V_{depth\_n}^{l}\right\}_{n=1}^{N}\in\mathbb{R}^{D\times H_{l}\times W_{l}}\right\}_{l=1}^{L=3}$.
-Meanwhile, the pseudo 3D point cloud $Pt^{cam}$ is acquired only from $\frac{1}{8}$ scale depth map $V_{depth\_n}$.
-Each pixel ray has a set of depth bins separated by 0.5m intervals.
-We followed the approach proposed in [7] to use 3D semantic occupancy annotations to supervise depth estimation.
-For each pixel ray, we take the highest probability depth bins of $K$ to form $Pt^{cam}$.
+Meanwhile, the pseudo 3D point cloud $Pt^{cam}$ is acquired only from $\frac{1}{8}$ scale depth map $V_{depth\_1}$.
+III-C Multi-scale-Fused Dense Depth-Maps Generation
+Figure 4: Multi-Scale Fused Dense Depth Maps Generation. This demonstrates the detailed multi-source depth map fusion from a per-pixel ray perspective.
+Given the multi-scale dense depth maps from cameras, $V^{cam}_{depth}$, and the multi-scale sparse depth maps from LiDAR, $V^{lidar}_{depth}$, we perform element-wise addition at the per-pixel level, followed by a value clamping operation.
+This process produces the final multi-scale fused dense depth maps, as illustrated in Fig.
+4.
+All depth maps from different sensors share a consistent representation, with the same number of depth bins $D$ and a uniform interval of 0.5 m between adjacent bins along each pixel ray.
+However, their value distributions differ: on the camera side, $V^{cam}_{depth}$ is represented as multi-hot per-pixel ray depth maps, where each depth bin takes a value within the range $[0,1]$.
+In contrast, $V^{lidar}_{depth}$ is represented as single-hot per-pixel ray depth maps, where only one depth bin is assigned a value of 1.0.
+When the camera and LiDAR depth maps are combined via per-pixel addition, the resulting fused pixel rays may contain depth-bin values exceeding 1.0.
+To ensure valid probability values, all depth-bin responses are subsequently clamped to the range $[0,1]$.
+This fusion strategy combines the complementary strengths of the two modalities.
+The dense camera depth maps provide rich structural coverage but are less accurate, whereas the LiDAR depth maps are sparse but geometrically precise.
+Their combination yields denser, more reliable fused depth maps, which are subsequently used to enhance visual features and facilitate cross-modal fusion.
 III-D Skeleton Merge Module
@@
 The upper LiDAR-branch serves as the main skeleton to provide a foundation structure for the 3D scene, and the bottom camera-branch serves as an augmentation based on the main skeleton to provide more detailed local structure to compensate for the fine-grained geometry of the main skeleton.
-The primitive anchor initialization process is shown in Figure 4.
+The primitive anchor initialization process is shown in Fig.
+5.
 Given the pseudo 3D point cloud derived from the camera $Pt^{cam}$ and the real 3D point cloud of the LiDAR $Pt^{lidar}$, we first apply the cylindrical partition to both point clouds, resulting in two volumes of voxels $Vol^{Cam}_{Cylin}$ and $Vol^{Lidar}_{Cylin}$ under cylindrical coordinates.
 We then predefine the total primitive anchors of $M+N$, where the anchors $M$ are sampled from $Vol^{Lidar}_{Cylin}$ and the anchors $N$ are sampled from $Vol^{Cam}_{Cylin}$, with a ratio of $M:N=3:1$.
-Following [37, 7], we adopt the farthest point sampling strategy to obtain $M$ Lidar anchors from $Vol^{Lidar}_{Cylin}$, which act as the main structural skeleton.
+We adopt the farthest point sampling strategy to obtain $M$ LiDAR anchors from $Vol^{Lidar}_{Cylin}$, which act as the main structural skeleton.
 Using these LiDAR anchors as reference, any voxel in $Vol^{Cam}_{Cylin}$ that spatially overlaps with the LiDAR anchors is removed.
@@
 Any voxel in $Vol^{Cam}_{Cylin}$ whose closest LiDAR anchor is farther away than $r$ is also discarded.
-Then, $N$ Camera anchors are obtained using the farthest point sampling strategy on the remaining voxels in $Vol^{Cam}_{Cylin}$.
+Then, $N$ Camera anchors are obtained using the same sampling strategy on the remaining voxels in $Vol^{Cam}_{Cylin}$.
 Based on the aforementioned filtering and sampling strategy, we acquire a final set of $M+N$ anchors whose locations will be used to initialize a set of T-primitives.
-The skeleton merge module takes advantage of the lidar strength in preserving the fine-grained 3D geometry of the scene and takes advantage of its $Vol^{Lidar}_{Cylin}$ to extract the main skeleton.
-Meanwhile, the module also leverages the camera’s strength in handling occlusion scenes and uses its $Vol^{Cam}_{Cylin}$ to extract the augmented skeleton surrounding the main skeleton.
-The fused skeleton preserves the finer-grained 3D geometry of the scene by combining the advantages of both modalities.
-III-E Multi-scale-Fused Dense Depth-Maps Generation
-Figure 5: Multi-Scale Fused Dense Depth Maps Generation.
-The upper part exhibits general multi-modality depth map fusion from a depth map perspective, and the bottom part demonstrates the detailed multi-modality depth map fusion from a per-pixel ray perspective.
-Given multi-scale camera dense depth maps $V^{cam}_{depth}$ and multi-scale lidar sparse depth maps $V_{depth}^{lidar}$, we first go through per-pixel level value addition followed by a value clamp, resulting in the final multi-scale fused dense depth maps as shown in Figure 5 upper.
-The detail of the addition of the per-pixel ray value is shown in Figure 5 bottom.
-For the camera side, each pixel ray has $D$ number of depth bins, and the interval between each depth bin is 0.5m.
-The DepthNet predicts the probability that each depth bin being occupied with a value in the range of [0, 1].
-On the lidar side, each pixel ray also has $D$ depth bins, but only one depth bin has been assigned a probability value of 1.0, while all other depth bins are assigned a probability value of 0.
-When adding the camera pixel ray to the lidar pixel ray, the resulting fused pixel ray may have some depth-bin probability values exceeding 1.0; therefore, these values need to be clamped within the range [0, 1].
-This multi-modality depth map fusion, on the one hand, solves the sparsity problem of lidar-based depth maps, and on the other hand, mitigates the inaccurate depth estimation problem of camera-based depth maps, resulting in denser and highly precise fused depth maps.
-III-F Early-Fusion Module
+This design exploits the complementary strengths of the two modalities: LiDAR provides the main geometric structure with higher spatial accuracy, while the camera branch supplements local details and occluded regions.
+The resulting fused skeleton, therefore, offers a better initialization for subsequent T-primitive modeling.
+III-E Early-Fusion Module
 Figure 6: Early-Fusion Module.
-Each occupied voxel centre in $Vol^{Lidar}_{Cylin}$ serves as an anchor and is projected onto multi-scale visual features to aggregate semantic information, yielding a semantic-aware lidar feature.
-The detailed inner structure of the Early-Fusion module is shown in Figure 6.
+Each occupied voxel center in $Vol^{Lidar}_{Cylin}$ serves as an anchor and is projected onto multi-scale visual features to aggregate semantic information, yielding a semantic-aware LiDAR feature.
+The detailed inner structure of the Early-Fusion module is shown in Fig.
+6.
 Given the sparse volume of the voxel $Vol_{Cylin}^{Lidar}$ in cylindrical coordinates, we leverage each occupied voxel center as an anchor and convert each anchor position from cylindrical coordinates to Cartesian coordinates.
 Then, each anchor is projected onto multi-scale visual features $V^{Cam}$ to aggregate semantic information at three scales.
-The aggregated three-scale visual features are weighted, summed, and concatenated back to the corresponding voxel feature, then fed through an MLP layer to perform the feature fusion, resulting in the last semantic-aware lidar feature $Vol_{sem}^{Lidar}$.
-III-G T-Primitive Representation
-Following the GaussianFormer [9, 7, 40] series, we propose the probabilistic Student’s T-distribution superposition as an efficient and effective 3D scene representation.
-Specifically, we propose three types of primitives that are enhanced by the inner kernel of T-distribution or T-distribution-like.
-The first is the Student-T-distribution-based primitive, which has the property as follows:
+The aggregated three-scale visual features are weighted, summed, and concatenated back to the corresponding voxel-associated LiDAR feature, then fed through an MLP layer to perform the feature fusion, resulting in the last semantic-aware LiDAR feature $Vol_{sem}^{Lidar}$.
+This module injects image semantics into sparse LiDAR features at an early stage, allowing the LiDAR branch to retain its geometric precision while gaining richer semantic context from the camera branch.
+III-F T-Primitive Representation
+We propose a probabilistic superposition of Student’s t-distribution-based primitives as an efficient and effective 3D scene representation.
+Specifically, we propose three types of primitives that are enhanced by the inner kernel of the T-distribution or T-distribution-like.
+The first is the Student-T-distribution-based plain primitive, which has the attributes as follows:
 $P_{i}^{T}(m_{i},s_{i},r_{i},\alpha_{i},c_{i})$ (3)
 where $m_{i}$, $s_{i}$, $r_{i}$, $\alpha_{i}$ and $c_{i}$ refer to the mean, scale value for the $x,y,z$ axes, quaternion vector, opacity and semantic vector of the $i$-th primitive, respectively.
-The second is the Student-T-distribution-based superquadric primitive, which has the property as follows:
-$P_{SQi}^{T}(m_{i},s_{i},r_{i},\alpha_{i},\epsilon_{1\_i},\epsilon_{2\_i},c_{i})$ (4)
+The second is the Student-T-distribution-based superquadric primitive, which has the attributes as follows:
+$P_{SQi}^{T}(m_{i},s_{i},r_{i},\alpha_{i},\epsilon_{1_{i}},\epsilon_{2_{i}},c_{i})$ (4)
 where $\epsilon_{1}$ and $\epsilon_{2}$ are the shape exponents of the superquadric.
-The third is the Student-T-distribution-based superquadric with inverse-warp primitive that has the property as follows:
-$P_{SQIWi}^{T}(m_{i},s_{i},r_{i},\alpha_{i},\epsilon_{1\_i},\epsilon_{2\_i},w_{i},c_{i})$ (5)
-where $\omega_{i}$ is the learnable weighting scaler valued between -1 and 1.
-Similarly to [7], we decompose the 3D modeling target into geometry and semantics predictions, and adopt the multiplication theorem of probability and the T-mixture model to address them probabilistically, respectively.
-What is more, compared to the Gaussian mixture model, the T-mixture model significantly improves robustness against outliers.
-III-G1 Student-T-Distribution Based Geometry Prediction
-To restrict T-Primitives to represent only occupied regions for geometry prediction, we interpret the T-distribution primitives $P^{T}=\left\{p^{T}_{i}\right\}_{i=1}^{K}$ as the probability of their surrounding space being occupied.
-In detail, we assign a probability value of 100% to the centers of the T-distribution, which decays with respect to the distance from the centers $m$:
-$\alpha\left(x;p^{T}_{i}\right)=\left[1+\frac{1}{\nu}\left(x-m\right)^{T}\Sigma^{-1}\left(x-m\right)\right]^{-\frac{\nu+3}{2}}$ (6)
-where $\nu$ refers to the degree of freedom, it is obtained by calculating how many voxels are intercepted by each T-primitive minus 1, and $\alpha\left(x;p^{T}_{i}\right)$ denotes the probability of the point x being occupied induced by the $i$-th T-primitive.
-Equation 6 assigns a high probability of occupancy when the point x is close to the center $m$ of the $i$-th T-primitive, which prevents any T-primitive from describing an empty area.
-To further derive the overall probability of occupancy, we assume that the probabilities of a point being occupied by different T-primitives are mutually independent, and thus we can aggregate them according to the multiplication theorem of probability:
-$\alpha\left(x\right)=1-\prod_{i=1}^{K}\left(1-\alpha\left(x;p^{T}_{i}\right)\right)$ (7)
-where $\alpha(x)$ represents the overall probability of occupancy at point x.
-III-G2 T-Mixture Model Based Semantic Prediction
-In addition to the Student-T-distribution-based Geometry modeling, we also need to perform the Student-T-distribution-based semantics prediction.
-Following [7], we interpret the set of T-primitives as a T-mixture model, where the semantics prediction could be formulated as calculating the expectation of semantics given the probabilistic T-mixture model.
-To be specific, we take the original opacity properties $\alpha$ as the prior distribution of the T-primitives, which is normalized by $l1$.
-Then, the T-primitive probabilistic distribution is adopted as a conditional probability.
-In addition, we normalize each T-primitive’s original semantics properties $c_{i}$ with a softmax activation function, yielding $\tilde{c}_{i}$ to ensure the boundness of predicted semantics.
-Lastly, the expectation calculation of the T-mixture model can be formulated as:
-$e\left(x;P^{T}\right)=\sum_{i=1}^{K}p\left(P^{T}_{i}|x\right)\tilde{c}_{i}=\frac{{\textstyle\sum_{i=1}^{K}p\left(x|P^{T}_{i}\right)a_{i}\tilde{c}_{i}}}{{\textstyle\sum_{j=1}^{K}p\left(x|P^{T}_{j}\right)a_{j}}}$ (8)
-where $p\left(P^{T}_{i}|x\right)$, $p\left(x|P^{T}_{i}\right)$ and $\tilde{c}_{i}$ denote the posterior probability of point x belonging to the $i$-th T-primitive distribution, the conditional probability of point x given the $i$-th T-primitive distribution, and the softmax-normalized semantic properties, respectively.
-Lastly, based on the geometry and semantic predictions, we combine the two prediction results together, resulting in the final 3D semantic occupancy prediction.
-The combination process is formulated as follows.
-$Occ=\left[1-\alpha\left(x\right);\alpha\left(x\right)\cdot e\left(x;P^{T}\right)\right]$ (9)
-where $1-\alpha\left(x\right)$ refers to the probability of the empty class and $\alpha\left(x\right)\cdot e\left(x;P^{T}\right)$ refers to the final semantic probabilities.
-The $\alpha\left(x\right)$ is used to weight the initial semantic prediction.
-III-G3 Student-T-Distribution Based Primitive
-The straightforward T-distribution primitive $p\left(x|P^{T}_{i}\right)$ can be formulated as:
-$p\left(x|P^{T}_{i}\right)$ $=\frac{\Gamma\left(\frac{\nu+3}{2}\right)}{\Gamma\left(\frac{\nu}{2}\right)\cdot\left(\nu\pi\right)^{3/2}\cdot\left|\Sigma\right|^{1/2}}\cdot$
-$\left[1+\frac{1}{\nu}\left(x-m\right)^{T}\Sigma^{-1}\left(x-m\right)\right]^{-\frac{\nu+3}{2}}$ (10)
-where $\nu$ refers to the degree of freedom, the same as in Equation 6, it is obtained by calculating how many voxels are intercepted by each T-primitive minus 1.
-$\Gamma$ refers to the gamma function, and performs the $\Gamma\left(n\right)=\left(n-1\right)!$ operation.
-Compared to Gaussian-primitive, this T-primitive is highly tailed in its probability distribution, leading to better robustness against outliers.
-III-G4 Student-T-Distribution Based SuperQuadric Primitive
-Following the quadricformer [40], we use the T-probabilistic modeling mechanism to convert superquadrics into occupancy probabilities.
-We first transform the 3D point position $(X,Y,Z)$, which is occupied by the T-Superquadric primitive, into its local coordinate system as:
-$\begin{bmatrix}X_{S}\\ Y_{S}\\ Z_{S}\\ \end{bmatrix}=R\times\begin{bmatrix}X-m_{x}\\ Y-m_{y}\\ Z-m_{z}\\ \end{bmatrix}$ (11)
-where $(X_{S},Y_{S},Z_{S})$ refers to the local coordinate under the T-Superquadric primitive, and R is the rotation matrix of each T-Superquadric primitive.
-Then, the occupancy probability of the 3D point $(X,Y,Z)$ associated with each T-Superquadric primitive can be computed as:
-$f\left(X_{S},Y_{S},Z_{S}\right)=\left(\left(\frac{X_{S}}{S_{x}}\right)^{\frac{2}{\epsilon_{2}}}+\left(\frac{Y_{S}}{S_{y}}\right)^{\frac{2}{\epsilon_{2}}}\right)^{\frac{\epsilon_{2}}{\epsilon_{1}}}+\left(\frac{Z_{S}}{S_{z}}\right)^{\frac{2}{\epsilon_{1}}}$ (12)
-$p\left(x|P^{T}_{SQi}\right)=\left(1+\frac{1}{\nu}\cdot f\left(X_{S},Y_{S},Z_{S}\right)\right)^{-\frac{\nu+3}{2}}$ (13)
-where $(S_{X},S_{Y},S_{Z})$ refers to the scale parameters along the X, Y, Z axes, $\epsilon_{1}$ and $\epsilon_{2}$ are the shape exponents of the superquadric.
-III-G5 Student-T-Distribution Based SuperQuadric With Inverse-Warp Primitive
+The third is the Student-T-distribution-based superquadric with inverse-warp primitive that has the attributes as follows:
+$P_{SQIWi}^{T}(m_{i},s_{i},r_{i},\alpha_{i},\epsilon_{1_{i}},\epsilon_{2_{i}},\left\{\omega_{i,k}\right\}_{k=1}^{24},c_{i})$ (5)
+where $\left\{\omega_{i,k}\right\}_{k=1}^{24}$ is a set of learnable weighting scalers used to determine the magnitude of applied deformation, and each $\omega_{i,k}$ has a value range between -1 and 1.
+Inspired by [7], we perform 3D geometry prediction and semantic prediction separately and adopt the multiplication theorem of probability and the T-mixture model to address them probabilistically.
+III-F1 Student-T-Distribution and Multiplication Theorem Based Geometry Prediction
+In practice, we interpret the T-distribution primitives $P^{T}=\left\{P^{T}_{i}\right\}_{i=1}^{K}$ as the probability of their surrounding space being occupied.
+In detail, we assign a probability value of 100% to the centers of the $P^{T}_{i}$, which decays with respect to the distance from the centers $m_{i}$.
+Assuming conditional independence of occupancy among different T-primitives, the final occupancy probability at x can be computed as:
+$p_{o}\left(x\right)=1-\prod_{i=1}^{K}\left(1-p\left(x|P^{T}_{i}\right)\right)$ (6)
+where $p\left(x|P^{T}_{i}\right)$ denotes the probability of point x being occupied induced by the $i$-th T-primitive.
+Depending on the primitive type, it is instantiated by (8), (12), or (16).
+$p_{o}\left(x\right)$ refers to the overall probability of occupancy at point x.
+III-F2 T-Mixture Model Based Semantic Prediction
+A set of T-primitives can be considered as a T-mixture model, and the final semantic predictions can be calculated via a weighted aggregation of semantic probabilities from all contributing T-primitives in the scene as:
+$p_{sem}\left(x\right)=\frac{{\textstyle\sum_{i=1}^{K}p\left(x|P^{T}_{i}\right)a_{i}\tilde{c}_{i}}}{{\textstyle\sum_{j=1}^{K}p\left(x|P^{T}_{j}\right)a_{j}}}$ (7)
+where $\tilde{c}_{i}$ represents the softmax-normalized semantic property of $i$-th T-primitive, and $a_{i}$ refers to $l1$ normalized opacity property of $i$-th T-primitive.
+III-F3 Student-T-Distribution Based Plain Primitive
+The plain T-distribution primitive $p\left(x|P^{T}_{i}\right)$ with the lowest geometric flexibility can be formulated as:
+$p\left(x|P^{T}_{i}\right)$ $=\frac{\Gamma\left(\frac{\nu_{i}+3}{2}\right)}{\Gamma\left(\frac{\nu_{i}}{2}\right)\cdot\left(\nu_{i}\pi\right)^{3/2}\cdot\left|\Sigma\right|^{1/2}}\cdot$
+$\left[1+\frac{1}{\nu_{i}}\left(x-m\right)^{T}\Sigma^{-1}\left(x-m\right)\right]^{-\frac{\nu_{i}+3}{2}}$ (8)
+where $\nu_{i}$ denotes the degrees of freedom of the $i$-th T-primitive.
+In our setting, we treat the number of voxels intersected by a T-primitive in the final 3D voxel grid as an approximation of its observable local support.
+Therefore, we adaptively define
+$\nu_{i}=N_{i}-1,$ (9)
+where $N_{i}$ is the number of intersected voxels associated with the $i$-th T-primitive.
+In this way, primitives supported by fewer voxels retain heavier tails for better robustness, while primitives supported by more voxels yield more concentrated distributions.
+This adaptive design allows each T-primitive to adjust its probabilistic shape according to its local geometric support in the scene.
+$\Gamma\left(\cdot\right)$ denotes the gamma function, which satisfies $\Gamma\left(n\right)=\left(n-1\right)!$ for positive integers $n$.
+Compared to the plain Gaussian-primitive, this T-primitive is highly tailed in its probability distribution, leading to better robustness against outliers.
+III-F4 Student-T-Distribution Based SuperQuadric Primitive
+The T-probabilistic modeling mechanism is adopted to convert superquadrics into occupancy probabilities.
+We first transform the 3D point position $(x,y,z)$, which is occupied by the $i$-th T-Superquadric primitive, into its local coordinate system as:
+$\begin{bmatrix}X_{S_{i}}\\ Y_{S_{i}}\\ Z_{S_{i}}\\ \end{bmatrix}=R_{i}\times\begin{bmatrix}x-m_{x_{i}}\\ y-m_{y_{i}}\\ z-m_{z_{i}}\\ \end{bmatrix}$ (10)
+where $(X_{S_{i}},Y_{S_{i}},Z_{S_{i}})$ refers to the local coordinate under the $i$-th T-Superquadric primitive, and $R_{i}$ is the rotation matrix of the $i$-th T-Superquadric primitive.
+Then, the occupancy probability of the 3D point $(x,y,z)$ associated with the $i$-th T-Superquadric primitive can be computed as:
+$f\left(X_{S_{i}},Y_{S_{i}},Z_{S_{i}}\right)=\left(\left(\frac{X_{S_{i}}}{S_{x_{i}}}\right)^{\frac{2}{\epsilon_{2_{i}}}}+\left(\frac{Y_{S_{i}}}{S_{y_{i}}}\right)^{\frac{2}{\epsilon_{2_{i}}}}\right)^{\frac{\epsilon_{2_{i}}}{\epsilon_{1_{i}}}}+\left(\frac{Z_{S_{i}}}{S_{z_{i}}}\right)^{\frac{2}{\epsilon_{1_{i}}}}$ (11)
+$p\left(x|P^{T}_{SQi}\right)=\left(1+\frac{1}{\nu_{i}}\cdot f\left(X_{S_{i}},Y_{S_{i}},Z_{S_{i}}\right)\right)^{-\frac{\nu_{i}+3}{2}}$ (12)
+where $(S_{x_{i}},S_{y_{i}},S_{z_{i}})$ refers to the scale parameters of $i$-th T-Superquadric primitive along the $X,Y,Z$ axes and $\nu_{i}$ refers to the degree of freedom, and has the same definition as in (8).
+III-F5 Student-T-Distribution Based SuperQuadric With Inverse-Warp Primitive
 So far, thanks to the T-distribution kernel enhancement, the T-Superquadric primitive is robust to outliers.
-Nevertheless, the T-Superquadric primitive has a very limited shape variety.
-It can only exhibit convex, symmetric shapes and cannot capture non-convex, asymmetric shapes, which are common in the daily 3D driving scene.
+Nevertheless, due to limited shape variety, it cannot capture non-convex, asymmetric shapes, which are very common in the daily 3D driving scene.
 To solve this limitation, we make the T-Superquadric primitive deformable through inverse-warp enhancement.
-Specifically, we choose 24 basis field functions $\left\{B_{i}(u,v,w)\right\}_{i=1}^{24}$, where $u,v,w$ refers to three inputs, respectively, to deform the local coordinates $(X_{S},Y_{S},Z_{S})$ as follows:
-$u=\frac{X_{S}}{S_{x}},v=\frac{Y_{S}}{S_{y}},w=\frac{Z_{S}}{S_{z}}$ (14)
-$\begin{bmatrix}\tilde{X}_{S}\\ \tilde{Y}_{S}\\ \tilde{Z}_{S}\\ \end{bmatrix}=\begin{bmatrix}X_{S}\\ Y_{S}\\ Z_{S}\\ \end{bmatrix}-\sum_{i=1}^{24}\omega_{i}\cdot B_{i}\left(u,v,w\right)$ (15)
-where $\omega_{i}$ is the learnable weighting scaler valued between -1 and 1.
-Then, the occupancy probability of the 3D point associated with each T-Superquadric-Inverse-Warp primitive can be calculated as follows:
-$f\left(\tilde{X}_{S},\tilde{Y}_{S},\tilde{Z}_{S}\right)=\left(\left(\frac{\tilde{X}_{S}}{S_{x}}\right)^{\frac{2}{\epsilon_{2}}}+\left(\frac{\tilde{Y}_{S}}{S_{y}}\right)^{\frac{2}{\epsilon_{2}}}\right)^{\frac{\epsilon_{2}}{\epsilon_{1}}}+\left(\frac{\tilde{Z}_{S}}{S_{z}}\right)^{\frac{2}{\epsilon_{1}}}$ (16)
-$p\left(x|P^{T}_{SQIWi}\right)=\left(1+\frac{1}{\nu}\cdot f\left(\tilde{X}_{S},\tilde{Y}_{S},\tilde{Z}_{S}\right)\right)^{-\frac{\nu+3}{2}}$ (17)
+Specifically, we choose 24 basis field functions $\left\{B_{k}(u,v,w)\right\}_{k=1}^{24}$ to deform the local coordinates $(X_{S_{i}},Y_{S_{i}},Z_{S_{i}})$ as follows:
+$u=\frac{X_{S_{i}}}{S_{x_{i}}},v=\frac{Y_{S_{i}}}{S_{y_{i}}},w=\frac{Z_{S_{i}}}{S_{z_{i}}}$ (13)
+$\begin{bmatrix}\tilde{X}_{S_{i}}\\ \tilde{Y}_{S_{i}}\\ \tilde{Z}_{S_{i}}\\ \end{bmatrix}=\begin{bmatrix}X_{S_{i}}\\ Y_{S_{i}}\\ Z_{S_{i}}\\ \end{bmatrix}-\sum_{k=1}^{24}\omega_{i,k}\cdot B_{k}\left(u,v,w\right)$ (14)
+Then, the occupancy probability of the 3D point associated with $i$-th T-Superquadric-Inverse-Warp primitive can be calculated as follows:
+$f\left(\tilde{X}_{S_{i}},\tilde{Y}_{S_{i}},\tilde{Z}_{S_{i}}\right)=\left(\left(\frac{\tilde{X}_{S_{i}}}{S_{x_{i}}}\right)^{\frac{2}{\epsilon_{2_{i}}}}+\left(\frac{\tilde{Y}_{S_{i}}}{S_{y_{i}}}\right)^{\frac{2}{\epsilon_{2_{i}}}}\right)^{\frac{\epsilon_{2_{i}}}{\epsilon_{1_{i}}}}+\left(\frac{\tilde{Z}_{S_{i}}}{S_{z_{i}}}\right)^{\frac{2}{\epsilon_{1_{i}}}}$ (15)
+$p\left(x|P^{T}_{SQIWi}\right)=\left(1+\frac{1}{\nu_{i}}\cdot f\left(\tilde{X}_{S_{i}},\tilde{Y}_{S_{i}},\tilde{Z}_{S_{i}}\right)\right)^{-\frac{\nu_{i}+3}{2}}$ (16)
+Basis Field $u$ $v$ $w$ $B_{1}$ 1.0 0.0 0.0 $B_{2}$ 0.0 1.0 0.0 $B_{3}$ 0.0 0.0 1.0 $B_{4}$ $u$ 0.0 0.0 $B_{5}$ 0.0 $v$ 0.0 $B_{6}$ 0.0 0.0 $w$ $B_{7}$ $v$ 0.0 0.0 $B_{8}$ $w$ 0.0 0.0 $B_{9}$ 0.0 $w$ 0.0 $B_{10}$ 0.0 $u$ 0.0 $B_{11}$ 0.0 0.0 $u$ $B_{12}$ 0.0 0.0 $v$ $B_{13}$ wv-w*v wuw*u 0.0 $B_{14}$ 0.0 uw-u*w uvu*v $B_{15}$ vwv*w 0.0 vu-v*u $B_{16}$ $w^{2}$ 0.0 0.0 $B_{17}$ 0.0 $w^{2}$ 0.0 $B_{18}$ 0.0 0.0 $u^{2}+v^{2}$ $B_{19}$ $u^{2}$ 0.0 0.0 $B_{20}$ 0.0 $v^{2}$ 0.0 $B_{21}$ 0.0 0.0 $w^{2}$ $B_{22}$ $(u^{2}+v^{2})*u$ $(u^{2}+v^{2})*v$ 0.0 $B_{23}$ uvu*v uvu*v 0.0 $B_{24}$ $u*v^{2}$ $u^{2}*v$ 0.0
+TABLE I: The 24 basis field functions make the T-Superquadric deformable. Those functions enable full shears, twists, bends, quadratics, radial bulge, and smooth corners.
 Each basis field function is listed in Table I.
@@
 Moreover, $B_{19}$ to $B_{24}$ provide quadratics, radial bulge, and smooth corner deformation.
-Basis Field $u$ $v$ $w$ $B_{1}$ 1.0 0.0 0.0 $B_{2}$ 0.0 1.0 0.0 $B_{3}$ 0.0 0.0 1.0 $B_{4}$ $u$ 0.0 0.0 $B_{5}$ 0.0 $v$ 0.0 $B_{6}$ 0.0 0.0 $w$ $B_{7}$ $v$ 0.0 0.0 $B_{8}$ $w$ 0.0 0.0 $B_{9}$ 0.0 $w$ 0.0 $B_{10}$ 0.0 $u$ 0.0 $B_{11}$ 0.0 0.0 $u$ $B_{12}$ 0.0 0.0 $v$ $B_{13}$ wv-w*v wuw*u 0.0 $B_{14}$ 0.0 uw-u*w uvu*v $B_{15}$ vwv*w 0.0 vu-v*u $B_{16}$ $w^{2}$ 0.0 0.0 $B_{17}$ 0.0 $w^{2}$ 0.0 $B_{18}$ 0.0 0.0 $u^{2}+v^{2}$ $B_{19}$ $u^{2}$ 0.0 0.0 $B_{20}$ 0.0 $v^{2}$ 0.0 $B_{21}$ 0.0 0.0 $w^{2}$ $B_{22}$ $(u^{2}+v^{2})*u$ $(u^{2}+v^{2})*v$ 0.0 $B_{23}$ uvu*v uvu*v 0.0 $B_{24}$ $u*v^{2}$ $u^{2}*v$ 0.0
-TABLE I: The 24 basis field functions make the T-Superquadric deformable. Those functions enable full shears, twists, bends, quadratics, radial bulge and smooth corners.
-III-H Transformer
-The general inner structure of the proposed transformer is shown in Figure 2.
-It takes a set of T-primitives, multi-scale lifted depth-aware visual features, and a semantic-aware lidar feature as input, and proceeds through a fused depth-map-guided 3D deformable attention module, a multi-gate cross-attention fusion module, two MLP-based feed-forward modules, a 3D sparse convolution-based self-attention module, and a T-primitive properties refine module iteratively $N$ times, resulting in the last refined T-primitives.
-The following sections will introduce the key implementation details of each module.
-III-H1 Fused Depth Map based 3D Deformable Attention
-Figure 7: 3D Deformable Attention For the same visual feature at location $(u_{i},v_{j})$, several different depth values ranging between $[0,1.0]$ along the depth-axis are used to weight the same pixel feature, resulting in the depth encoded visual feature to mitigate the projection ambiguity problem stemming from the projection process.
-Taking multi-scale lifted depth-aware visual features and a set of T-primitive as input, we use a 3D deformable attention operator, namely DFA3D [11], which performs feature aggregation on lifted visual features to mitigate the 3D to 2D projection ambiguity problem, to aggregate visual features for each T-primitive.
-As shown in Figure 2 camera branch and Figure 7, the multi-scale lifted depth-aware visual features $Vol^{Cam}_{depth}$ are obtained through outer product operation $F_{depth}^{Cam}\otimes V_{depth}^{fuse}$, this operation gives an accurate depth encoding for each visual feature in the feature map.
+III-G Transformer
+III-G1 Fused Depth Map Guided 3D Deformable Attention
+Figure 7: Fused Depth Map Guided 3D Deformable Attention For the same visual feature at location $(u_{i},v_{j})$, several different depth values ranging between $[0.0,1.0]$ along the depth-axis are used to weight the same pixel feature, resulting in the depth encoded visual feature to mitigate the projection ambiguity problem stemming from the projection process.
+Taking multi-scale lifted depth-aware visual features and a set of T-primitive as input, we use a 3D deformable attention operator, namely 3DFA [11], which performs feature aggregation on lifted visual features to mitigate the 3D to 2D projection ambiguity problem, to aggregate visual features for each T-primitive.
+As shown in Fig.
+2 camera branch and Fig.
+7, the multi-scale lifted depth-aware visual features $Vol^{Cam}_{depth}$ are obtained through outer product operation $F_{depth}^{Cam}\otimes V_{depth}^{fuse}$, this operation gives an accurate depth encoding for each visual feature in the feature map.
 For each T-primitive feature sampling $p^{T}_{i}$, we first use the mean location $m_{i}$ of each $p^{T}_{i}$ as the center, convert the center location from the cylindrical coordinate to the cartesian coordinate, and feed its associate query feature $q_{i}$ to an MLP layer, resulting in 13 offsets $\Delta m_{j}$ with respect to the center to form a set of 3D sampling points $m^{samp}=\left\{m^{samp}_{i}=m_{i}+\Delta m_{j}\mid j=1,...,13\right\}$.
 Then, we leverage extrinsic and intrinsic matrices to project each 3D sampling point $m^{samp}_{i}$ onto the lifted depth-aware visual features, resulting in the location of the sampling of the pixel frame $\bar{m}^{samp}_{i}=\left(u_{i},v_{i},d_{i}\right)$.
-Lastly, a 3D deformable attention operator DFA3D takes $\bar{m}^{samp}_{i}$ as a sampling location to perform cross-attention visual feature aggregation on $Vol^{Cam}_{depth}$, resulting in an updated query feature $q_{i}^{updated}$.
-The overall sampling process can be formulated as follows:
-$q_{i}^{updated}=\sum_{i=1}^{13}DFA3D\left(q_{i},\bar{m}^{samp}_{i},Vol^{Cam}_{depth}\right)$ (18)
-After the 3D cross-attention process, $q_{i}^{updated}$ already aggregates sufficient depth-aware visual features and is passed to the MGCAFusion module for feature fusion.
-Meanwhile, the set of 3D sampling points $m^{samp}$ for each T-primitive is also passed on to the MGCAFusion module for semantic-aware lidar feature aggregation.
-III-H2 Multi-Gate Cross-Attention Fusion (MGCAFusion) Module
-Figure 8: Multi-Gate Cross Attention Fusion Module.
+Lastly, a 3D deformable attention operator 3DFA takes $\bar{m}^{samp}_{i}$ as a sampling location to perform cross-attention visual feature aggregation on $Vol^{Cam}_{depth}$, resulting in an updated query feature $q_{i}^{updated}$.
+The overall sampling process can be formulated as follows.
+$q_{i}^{updated}=\sum_{i=1}^{13}3DFA\left(q_{i},\bar{m}^{samp}_{i},Vol^{Cam}_{depth}\right)$ (17)
+After the 3D cross-attention process, $q_{i}^{updated}$ already aggregates sufficient depth-aware visual features and is passed to the GCWSFusion module for feature fusion.
+Meanwhile, the set of 3D sampling points $m^{samp}$ for each T-primitive is also passed on to the GCWSFusion module for semantic-aware LiDAR feature aggregation.
+III-G2 Gate-Concatenation and Weight-Summation Fusion (GCWSFusion) Module
+Figure 8: Gate-Concatenation and Weight-Summation Fusion Module.
 The left part exhibits the multi-modality weighted summation fusion part, and the right part exhibits the multi-modality gated concatenation fusion part.
-The outputs of the two major parts are further fused to produce the final fused feature.
+The outputs of the two major parts are further fused to produce the final fused feature for each T-Primitive.
 Given a set of 3D sampling points $m^{samp}$ for each T-primitive, we apply bilinear interpolation on $Vol^{Lidar}_{sem}$ to aggregate a semantic-aware LiDAR feature for each T-primitive.
-The detailed implementation of the multi-gate cross attention fusion module (MGCAFusion) is demonstrated in Figure 8.
-It consists of two central parts, the weighted summation fusion part, as highlighted in the green rectangle on the left of Figure 8 and the gated concatenation fusion part, as highlighted in the pink rectangle on the right of Figure 8.
+Then, both the semantic-aware LiDAR feature and the depth-aware visual feature of each T-primitive are fed into the GCWSFusion module for late-stage feature fusion.
+The detailed implementation of GCWSFusion module is demonstrated in Fig.
+8.
+It consists of two central parts, the weighted summation fusion part, as highlighted in the green rectangle on the left of Fig.
+8, and the gated concatenation fusion part, as highlighted in the pink rectangle on the right of Fig.
+8.
 Both parts take the same inputs, but perform different feature fusion strategies.
-For the weighted summation fusion part, the semantic-aware lidar feature and the depth-aware visual feature are summed in elements first, then passed through an MLP layer followed by a softmax layer, resulting in two weight values, $W_{Lidar}$ and $W_{Vis}$.
+For the weighted summation fusion part, the semantic-aware LiDAR feature and the depth-aware visual feature are summed in elements first, then passed through several linear layers followed by a softmax layer, resulting in two weight values, $W_{Lidar}$ and $W_{Vis}$.
 The $W_{Lidar}$ and $W_{Vis}$ are multiplied again with the corresponding semantic-aware LiDAR feature and depth-aware visual feature, respectively, resulting in the weighted LiDAR feature and the weighted visual feature.
-Lastly, the weighted lidar feature and the weighted visual feature are summed and passed through an MLP layer, yielding the summation fusion feature.
-For the gated concatenation fusion part, the semantic-aware lidar feat and the depth-aware visual feat are passed through an MLP layer, followed by a sigmoid function, respectively, resulting in a visual gate vector $G_{Vis}$ and a lidar gate vector $G_{Lidar}$.
+Lastly, the weighted LiDAR feature and the weighted visual feature are summed and passed through an extra set of linear layers, yielding the summation fusion feature.
+This design adaptively weights the dominance information across two sensors.
+For the gated concatenation fusion part, the semantic-aware LiDAR feat and the depth-aware visual feat are passed through several linear layers, followed by a sigmoid function, respectively, resulting in a visual gate vector $G_{Vis}$ and a LiDAR gate vector $G_{Lidar}$.
 $G_{Vis}$ is element-wise multiplied with the semantic-aware LiDAR feature, followed by a skip connection that results in the gated LiDAR feature.
 $G_{Lidar}$ is element-wise multiplied with the depth-aware visual feature, followed by a skip connection that results in the gated visual feature.
-The gated Lidar and visual features are concatenated along the feature channel dimension, followed by an MLP layer that reduces the fused feature channel dimension, yielding the concatenation fusion feature.
-Lastly, the summation and concatenation fusion features are combined via feature-channel concatenation, followed by an MLP layer, yielding the final fused feature.
-III-H3 3D Sparse Conv-based Self-Attention Module
-Figure 9: Sparse-Convolution-Based Self-Attention Pipeline.
-The center of T-primitives is treated as a 3D point to form a primitive-based 3D point cloud.
-We apply a cylindrical partition to the obtained 3D point cloud, followed by a 3D sparse convolution to realise self-attention.
-The self-attention module is demonstrated in Figure 9.
-Each T-primitive center $m_{i}$ is treated as a 3D point, and we can acquire a set of 3D points based on T-primitive.
+The gated LiDAR and visual features are concatenated along the feature channel dimension, followed by another set of linear layers that reduces the fused feature channel dimension, yielding the concatenation fusion feature.
+This design can dynamically amplify or suppress feature information within the two different modalities, thereby achieving denoising and information enhancement during the feature fusion process.
+Lastly, the summation and concatenation fusion features are combined via feature-channel concatenation, followed by extra linear layers, yielding the final fused feature.
+III-G3 3D Sparse Conv-based Self-Attention Module
+Each T-primitive center $m_{i}$ is treated as a 3D point, and we can acquire a set of 3D points based on a set of T-primitives.
 Then, this T-primitive-based sparse 3D point cloud undergoes a coordinate transformation from Cartesian to Cylindrical coordinates, followed by voxelization, yielding a sparse voxel volume in cylindrical coordinates.
@@
 Since our T-primitives are initialized in cylindrical coordinates, we perform the self-attention operation in the same coordinates to better preserve the scene’s fine-grained 3D geometry.
-III-H4 T-Primitive Property Refinement Module
-Since T-primitive queries have aggregated sufficient semantic and 3D geometry information in the previous 3D deformable attention, self-attention, and MGCAFusion modules, we use each query vector to update its associated T-primitive properties.
-Specifically, for three types of T-primitives, we feed each type of T-primitive query to an MLP layer to obtain the intermediate properties, and we combine the intermediate properties with the initial properties to achieve the refinement of the T-primitive property.
-The property refinement details for three types of T-primitives are formulated as:
-$MLP(Q_{P_{i}})\Longrightarrow(\hat{m}_{i},\hat{s}_{i},\hat{r}_{i},\hat{c}_{i})$
-$MLP(Q_{SQ_{i}})\Longrightarrow(\hat{m}_{i},\hat{s}_{i},\hat{r}_{i},\hat{\epsilon}_{1_{i}},\hat{\epsilon}_{2_{i}},\hat{c}_{i})$
-$MLP(Q_{SQIW_{i}})\Longrightarrow(\hat{m}_{i},\hat{s}_{i},\hat{r}_{i},\hat{\epsilon}_{1_{i}},\hat{\epsilon}_{2_{i}},\hat{w}_{i},\hat{c}_{i})$
-$P^{T}_{New_{i}}=(m+\hat{m}_{i},\hat{s}_{i},\hat{r}_{i},\hat{c}_{i})$
-$P^{T}_{New_{SQ_{i}}}=(m+\hat{m}_{i},\hat{s}_{i},\hat{r}_{i},\hat{\epsilon}_{1_{i}},\hat{\epsilon}_{2_{i}},\hat{c}_{i})$
-$P^{T}_{New_{SQIW_{i}}}=(m+\hat{m}_{i},\hat{s}_{i},\hat{r}_{i},\hat{\epsilon}_{1_{i}},\hat{\epsilon}_{2_{i}},\hat{w}_{i},\hat{c}_{i})$ (19)
+III-G4 T-Primitive Property Refinement
+As presented in (18), the physical attributes of the three proposed T-Primitives are refined by processing each associated query vector through an MLP to produce updated attributes, which are subsequently added to the original attributes for refinement.
+$(\Delta{m}_{i},\hat{s}_{i},\hat{r}_{i},\hat{\alpha}_{i},\hat{c}_{i})=MLP(Q_{P_{i}})$
+$(\Delta{m}_{i},\hat{s}_{i},\hat{r}_{i},\hat{\alpha}_{i},\hat{\epsilon}_{1_{i}},\hat{\epsilon}_{2_{i}},\hat{c}_{i})=MLP(Q_{SQ_{i}})$
+$(\Delta{m}_{i},\hat{s}_{i},\hat{r}_{i},\hat{\alpha}_{i},\hat{\epsilon}_{1_{i}},\hat{\epsilon}_{2_{i}},\left\{\hat{\omega_{i,k}}\right\}_{k=1}^{24},\hat{c}_{i})=MLP(Q_{SQIW_{i}})$
+$\hat{P^{T}_{i}}=(m_{i}+\Delta{m}_{i},\hat{s}_{i},\hat{r}_{i},\hat{\alpha}_{i},\hat{c}_{i})$
+$\hat{P^{T}_{SQ_{i}}}=(m_{i}+\Delta{m}_{i},\hat{s}_{i},\hat{r}_{i},\hat{\alpha}_{i},\hat{\epsilon}_{1_{i}},\hat{\epsilon}_{2_{i}},\hat{c}_{i})$
+$\hat{P^{T}_{SQIW_{i}}}=(m_{i}+\Delta{m}_{i},\hat{s}_{i},\hat{r}_{i},\hat{\alpha}_{i},\hat{\epsilon}_{1_{i}},\hat{\epsilon}_{2_{i}},\left\{\hat{\omega_{i,k}}\right\}_{k=1}^{24},\hat{c}_{i})$ (18)
 IV Experimental Results
@@
 The learning rate is decayed using a multistep scheduler.
-The predicted occupancy has a resolution of $200\times 200\times 16$ for full-scale evaluation on nuScenes-Surroundocc, nuScenes-Occ3D and nuScenes-C datasets.
+The predicted occupancy has a resolution of $200\times 200\times 16$ for full-scale evaluation on SurroundOcc-nuScenes, Occ3D-nuScenes and nuScenes-C datasets.
 For data augmentation, we employ photometric distortions and a grid mask to the input images during training.
@@
 For depth supervision, a binary cross-entropy loss $L^{depth}_{BCE}$ is used to refine the depth distribution feature.
-The final loss is formulated as follows.
+The final loss is formulated as:
 $L_{Total}=L_{lovasz}+\lambda\cdot L_{BCE}+L_{BCE}^{depth}$ (19)
@@
 Since the test set lacks semantic labels, we train our model on the training set and evaluate its performance using the validation set.
-For annotations from SurroundOcc, we set the range of the X and Y axes to [-50, 50] meters and the Z axis to [-5, 3] meters under lidar coordinates.
+For SurroundOcc annotations, we set the range of the X and Y axes to [-50, 50] meters and the Z axis to [-5, 3] meters under LiDAR coordinates.
 For annotations from Occ3D, we set the range of the X and Y axes to [-40, 40] meters and the Z axis to [-1, 5.4] meters under the ego coordinates.
@@
 Label 0 refers to free voxels in both annotations.
-Additionally, following the methodology proposed in [18], we conduct an in-depth analysis of our model performance in challenging scenarios, specifically in rainy and night conditions.
-This evaluation is carried out using the annotation file provided by [18].
-IV-D Performance Evaluate Metrics
-To assess the performance of various SOTA algorithms and compare them with our approach in the 3D semantic occupancy prediction task, we utilize the intersection over union (IoU) to evaluate each semantic class.
-Moreover, we employ the mean IoU over all semantic classes (mIoU) as a comprehensive evaluation metric:
-$IoU=\frac{TP}{TP+FP+FN}$ (21)
-and
-$mIoU=\frac{1}{Cls}\sum_{i=1}^{Cls}\frac{TP_{i}}{TP_{i}+FP_{i}+FN_{i}}$ (22)
-where $TP$, FPFP, and $FN$ represent the counts of true positives, false positives, and false negatives in our predictions, respectively, while $Cls$ denotes the total class number.
-IV-E Model Performance Analysis
+IV-D Model Performance Analysis
 To demonstrate the performance of our proposed TFusionOcc, we evaluated it on the SurroundOcc-nuScenes and Occ3D-nuScenes benchmarks.
-For each benchmark, we evaluate our model across three primitive settings, including T-P (general T-primitive), T-SQ (T-superquadric), and T-SQ-IW (deformable T-superquadric or T-superquadric with inverse-warp), with 12800 and 25600 total primitives, respectively.
+For each benchmark, we evaluate our model across three primitive settings, including T-P (plain T-primitive), T-SQ (T-superquadric), and T-SQ-IW (deformable T-superquadric or T-superquadric with inverse-warp), with 12800 and 25600 total primitives, respectively.
 The results of the experiment on the SurroundOcc-nuScenes and Occ3D-nuScenes benchmarks are shown in Table II and Table III, respectively.
-In both benchmarks, TFusionOcc (T-SQ-IW) settings achieve the best and second-best performance, demonstrating the importance of the primitives’ shape variety and robustness.
-Specifically, our proposed method excels at detecting background objects, including drivable surfaces, other flat areas, sidewalks, terrain, man-made structures, and vegetation.
-In addition, our model performs significantly well on relatively small foreground objects, such as barriers, bicycles, cars, motorcycles, pedestrians, and traffic cones.
-Our method, which only adopts a single 3D supervision signal, achieved par-performance in comparison with algorithms that leverage multi-task learning to receive dual supervision signals, such as DAOcc [33] in the Surroundocc-nuScenes benchmark that incorporates a 3D object detection head and a 3D semantic occupancy prediction head during training, and SDGOcc [4] in the Occ3D-nuScenes benchmark, which incorporates a 2D semantic segmentation head and a 3D semantic occupancy prediction head during training.
 Method 2D-Backbone Modality IoU mIoU $\bullet$ barrier $\bullet$ bicycle $\bullet$ bus $\bullet$ car $\bullet$ const.
@@
 surf.
-$\bullet$ other flat $\bullet$ sidewalk $\bullet$ terrain $\bullet$ manmade $\bullet$ vegetation MonoScene [3] R101-DCN C 23.96 7.31 4.03 0.35 8.00 8.04 2.90 0.28 1.16 0.67 4.01 4.35 27.72 5.20 15.13 11.29 9.03 14.86 Atlas [20] - C 28.66 15.00 10.64 5.68 19.66 24.94 8.90 8.84 6.47 3.28 10.42 16.21 34.86 15.46 21.89 20.95 11.21 20.54 BEVFormer [13] R101-DCN C 30.50 16.75 14.22 6.58 23.46 28.28 8.66 10.77 6.64 4.05 11.20 17.78 37.28 18.00 22.88 22.17 13.80 22.21 TPVFormer [8] R101-DCN C 30.86 17.10 15.96 5.31 23.86 27.32 9.79 8.74 7.09 5.20 10.97 19.22 38.87 21.25 24.26 23.15 11.73 20.81 C-CONet [29] R101 C 26.10 18.40 18.60 10.00 26.40 27.40 8.60 15.70 13.30 9.70 10.90 20.20 33.00 20.70 21.40 21.80 14.70 21.30 InverseMatrixVT3D [17] R101-DCN C 30.03 18.88 18.39 12.46 26.30 29.11 11.00 15.74 14.78 11.38 13.31 21.61 36.30 19.97 21.26 20.43 11.49 18.47 OccFormer [36] R101-DCN C 31.39 19.03 18.65 10.41 23.92 30.29 10.31 14.19 13.59 10.13 12.49 20.77 38.78 19.79 24.19 22.21 13.48 21.35 FB-Occ [14] R101 C 31.50 19.60 20.60 11.30 26.90 29.80 10.40 13.60 13.70 11.40 11.50 20.60 38.20 21.50 24.60 22.70 14.80 21.60 RenderOcc [22] R101 C 29.20 19.00 19.70 11.20 28.10 28.20 9.80 14.70 11.80 11.90 13.10 20.10 33.20 21.30 22.60 22.30 15.30 20.90 GaussianFormer [9] R101-DCN C 29.83 19.10 19.52 11.26 26.11 29.78 10.47 13.83 12.58 8.67 12.74 21.57 39.63 23.28 24.46 22.99 9.59 19.12 Co-Occ [21] R101 C 30.00 20.30 22.50 11.20 28.60 29.50 9.90 15.80 13.50 8.70 13.60 22.20 34.90 23.10 24.20 24.10 18.00 24.80 GaussianFormer2-256 [7] R101-DCN C 31.74 20.82 21.39 13.44 28.49 30.82 10.92 15.84 13.55 10.53 14.04 22.92 40.61 24.36 26.08 24.27 13.83 21.98 SurroundOcc [30] R101-DCN C 31.49 20.30 20.59 11.68 28.06 30.86 10.70 15.14 14.09 12.06 14.38 22.26 37.29 23.70 24.49 22.77 14.89 21.86 Inverse++ [19] R101-DCN C 31.73 20.91 20.90 13.27 28.40 31.37 11.90 17.76 15.39 13.49 13.32 23.19 39.37 22.85 25.27 23.68 13.43 20.98 LMSCNet [25] - L 36.60 14.90 13.10 4.50 14.70 22.10 12.60 4.20 7.20 7.10 12.20 11.50 26.30 14.30 21.10 15.20 18.50 34.20 L-CONet [29] - L 39.40 17.70 19.20 4.00 15.10 26.90 6.20 3.80 6.80 6.00 14.10 13.10 39.70 19.10 24.00 23.90 25.10 35.70 Co-Occ [21] - L 42.20 22.90 22.00 6.90 25.70 32.40 14.50 13.50 21.00 10.50 18.00 22.50 36.60 21.80 24.60 25.70 31.20 39.90 Co-Occ [21] R101-DCN C+L 41.10 27.10 28.10 16.10 34.00 37.20 17.00 21.60 20.80 15.90 21.90 28.70 42.30 25.40 29.10 28.60 28.20 38.00 GaussianFormer3D [38] R101-DCN C+L 43.30 27.10 26.90 15.80 32.70 36.10 18.60 21.70 24.10 13.00 21.30 29.00 40.60 23.70 27.30 28.20 32.60 42.30 OccFusion [18] R101-DCN C+R 33.97 20.73 20.46 13.98 27.99 31.52 13.68 18.45 15.79 13.05 13.94 23.84 37.85 19.60 22.41 21.20 16.16 21.81 OccFusion [18] R101-DCN C+L 43.53 27.55 25.15 19.87 34.75 36.21 20.03 23.11 25.25 17.50 22.70 30.06 39.47 23.26 25.68 27.57 29.54 40.60 OccFusion [18] R101-DCN C+L+R 43.96 28.27 28.32 20.95 35.06 36.84 20.33 26.22 25.86 19.17 21.27 30.64 40.08 23.75 25.56 27.63 29.82 40.82 GaussianFusionOcc [23] R101-DCN C+L 45.16 30.21 30.22 18.70 35.91 39.57 22.67 27.36 30.10 18.59 24.45 31.25 43.06 25.76 29.12 29.33 34.65 42.70 GaussianFusionOcc [23] R101-DCN C+L+R 45.20 30.37 30.43 18.54 36.23 39.66 22.57 27.35 30.30 19.14 24.56 31.95 42.60 25.82 29.48 29.70 34.78 42.95 OccCylindrical [16] Res-50 C+L 44.94 28.67 26.17 22.12 31.47 36.84 17.95 27.77 29.85 23.90 20.64 28.27 43.00 23.14 27.99 27.81 30.81 40.95 DAOcc [33] Res-50 C+L 45.00 30.50 30.80 19.50 34.00 38.80 25.00 27.70 29.90 22.50 23.20 31.60 41.00 25.90 29.40 29.90 35.20 43.50 TFusionOcc (T-SQ-12800) Res-50 C+L 46.70 30.76 30.67 21.27 33.12 38.96 22.09 27.54 30.47 22.87 21.65 29.86 42.63 27.18 29.80 30.91 36.81 46.26 TFusionOcc (T-SQ-IW-12800) Res-50 C+L 46.92 31.21 31.65 20.13 34.75 39.70 23.08 27.00 30.89 23.33 22.02 30.40 43.11 27.78 30.25 31.47 37.15 46.60 TFusionOcc (T-P-25600) Res-50 C+L 45.49 30.04 30.15 19.58 34.20 38.64 22.10 26.60 29.85 22.24 22.44 29.65 41.08 25.19 28.43 29.71 35.77 45.01 TFusionOcc (T-SQ-25600) Res-50 C+L 46.14 30.88 30.87 21.26 33.68 39.25 22.54 28.34 31.45 23.61 22.80 30.16 41.76 26.12 29.23 30.26 36.77 46.06 TFusionOcc (T-SQ-IW-25600) Res-50 C+L 47.01 31.47 31.96 20.78 34.57 39.73 24.02 27.82 31.19 23.82 23.05 30.67 43.17 27.18 30.13 31.18 37.52 46.71
+$\bullet$ other flat $\bullet$ sidewalk $\bullet$ terrain $\bullet$ manmade $\bullet$ vegetation MonoScene [3] R101-DCN C 23.96 7.31 4.03 0.35 8.00 8.04 2.90 0.28 1.16 0.67 4.01 4.35 27.72 5.20 15.13 11.29 9.03 14.86 Atlas [20] - C 28.66 15.00 10.64 5.68 19.66 24.94 8.90 8.84 6.47 3.28 10.42 16.21 34.86 15.46 21.89 20.95 11.21 20.54 BEVFormer [13] R101-DCN C 30.50 16.75 14.22 6.58 23.46 28.28 8.66 10.77 6.64 4.05 11.20 17.78 37.28 18.00 22.88 22.17 13.80 22.21 TPVFormer [8] R101-DCN C 30.86 17.10 15.96 5.31 23.86 27.32 9.79 8.74 7.09 5.20 10.97 19.22 38.87 21.25 24.26 23.15 11.73 20.81 C-CONet [30] R101 C 26.10 18.40 18.60 10.00 26.40 27.40 8.60 15.70 13.30 9.70 10.90 20.20 33.00 20.70 21.40 21.80 14.70 21.30 InverseMatrixVT3D [17] R101-DCN C 30.03 18.88 18.39 12.46 26.30 29.11 11.00 15.74 14.78 11.38 13.31 21.61 36.30 19.97 21.26 20.43 11.49 18.47 OccFormer [38] R101-DCN C 31.39 19.03 18.65 10.41 23.92 30.29 10.31 14.19 13.59 10.13 12.49 20.77 38.78 19.79 24.19 22.21 13.48 21.35 FB-Occ [14] R101 C 31.50 19.60 20.60 11.30 26.90 29.80 10.40 13.60 13.70 11.40 11.50 20.60 38.20 21.50 24.60 22.70 14.80 21.60 RenderOcc [22] R101 C 29.20 19.00 19.70 11.20 28.10 28.20 9.80 14.70 11.80 11.90 13.10 20.10 33.20 21.30 22.60 22.30 15.30 20.90 GaussianFormer [9] R101-DCN C 29.83 19.10 19.52 11.26 26.11 29.78 10.47 13.83 12.58 8.67 12.74 21.57 39.63 23.28 24.46 22.99 9.59 19.12 Co-Occ [21] R101 C 30.00 20.30 22.50 11.20 28.60 29.50 9.90 15.80 13.50 8.70 13.60 22.20 34.90 23.10 24.20 24.10 18.00 24.80 GaussianFormer2-256 [7] R101-DCN C 31.74 20.82 21.39 13.44 28.49 30.82 10.92 15.84 13.55 10.53 14.04 22.92 40.61 24.36 26.08 24.27 13.83 21.98 QuadricFormer [42] R101-DCN C 31.22 20.12 19.58 13.11 27.27 29.64 11.25 16.26 12.65 9.15 12.51 21.24 40.20 24.34 25.69 24.24 12.95 21.86 SurroundOcc [31] R101-DCN C 31.49 20.30 20.59 11.68 28.06 30.86 10.70 15.14 14.09 12.06 14.38 22.26 37.29 23.70 24.49 22.77 14.89 21.86 Inverse++ [19] R101-DCN C 31.73 20.91 20.90 13.27 28.40 31.37 11.90 17.76 15.39 13.49 13.32 23.19 39.37 22.85 25.27 23.68 13.43 20.98 LMSCNet [25] - L 36.60 14.90 13.10 4.50 14.70 22.10 12.60 4.20 7.20 7.10 12.20 11.50 26.30 14.30 21.10 15.20 18.50 34.20 L-CONet [30] - L 39.40 17.70 19.20 4.00 15.10 26.90 6.20 3.80 6.80 6.00 14.10 13.10 39.70 19.10 24.00 23.90 25.10 35.70 Co-Occ [21] - L 42.20 22.90 22.00 6.90 25.70 32.40 14.50 13.50 21.00 10.50 18.00 22.50 36.60 21.80 24.60 25.70 31.20 39.90 Co-Occ [21] R101-DCN C+L 41.10 27.10 28.10 16.10 34.00 37.20 17.00 21.60 20.80 15.90 21.90 28.70 42.30 25.40 29.10 28.60 28.20 38.00 GaussianFormer3D [39] R101-DCN C+L 43.30 27.10 26.90 15.80 32.70 36.10 18.60 21.70 24.10 13.00 21.30 29.00 40.60 23.70 27.30 28.20 32.60 42.30 OccFusion [18] R101-DCN C+R 33.97 20.73 20.46 13.98 27.99 31.52 13.68 18.45 15.79 13.05 13.94 23.84 37.85 19.60 22.41 21.20 16.16 21.81 OccFusion [18] R101-DCN C+L 43.53 27.55 25.15 19.87 34.75 36.21 20.03 23.11 25.25 17.50 22.70 30.06 39.47 23.26 25.68 27.57 29.54 40.60 OccFusion [18] R101-DCN C+L+R 43.96 28.27 28.32 20.95 35.06 36.84 20.33 26.22 25.86 19.17 21.27 30.64 40.08 23.75 25.56 27.63 29.82 40.82 GaussianFusionOcc [23] R101-DCN C+L 45.16 30.21 30.22 18.70 35.91 39.57 22.67 27.36 30.10 18.59 24.45 31.25 43.06 25.76 29.12 29.33 34.65 42.70 GaussianFusionOcc [23] R101-DCN C+L+R 45.20 30.37 30.43 18.54 36.23 39.66 22.57 27.35 30.30 19.14 24.56 31.95 42.60 25.82 29.48 29.70 34.78 42.95 OccCylindrical [16] Res-50 C+L 44.94 28.67 26.17 22.12 31.47 36.84 17.95 27.77 29.85 23.90 20.64 28.27 43.00 23.14 27.99 27.81 30.81 40.95 DAOcc [35] Res-50 C+L 45.00 30.50 30.80 19.50 34.00 38.80 25.00 27.70 29.90 22.50 23.20 31.60 41.00 25.90 29.40 29.90 35.20 43.50 TFusionOcc (T-SQ-12800) Res-50 C+L 46.70 30.76 30.67 21.27 33.12 38.96 22.09 27.54 30.47 22.87 21.65 29.86 42.63 27.18 29.80 30.91 36.81 46.26 TFusionOcc (T-SQ-IW-12800) Res-50 C+L 46.92 31.21 31.65 20.13 34.75 39.70 23.08 27.00 30.89 23.33 22.02 30.40 43.11 27.78 30.25 31.47 37.15 46.60 TFusionOcc (T-P-25600) Res-50 C+L 45.49 30.04 30.15 19.58 34.20 38.64 22.10 26.60 29.85 22.24 22.44 29.65 41.08 25.19 28.43 29.71 35.77 45.01 TFusionOcc (T-SQ-25600) Res-50 C+L 46.14 30.88 30.87 21.26 33.68 39.25 22.54 28.34 31.45 23.61 22.80 30.16 41.76 26.12 29.23 30.26 36.77 46.06 TFusionOcc (T-SQ-IW-25600) Res-50 C+L 47.01 31.47 31.96 20.78 34.57 39.73 24.02 27.82 31.19 23.82 23.05 30.67 43.17 27.18 30.13 31.18 37.52 46.71
 TABLE II: 3D semantic occupancy prediction results on SurroundOcc-nuScenes validation set.
-Our approach outperforms other existing methods with the same input modality.
-For readers’ reference, the bottom of the table presents results from three additional methods using different input modalities.
-All methods are trained with dense occupancy labels from SurroundOcc [30].
 Notion of modality: Camera (C), LiDAR (L), Radar (R).
@@
 TABLE III: 3D semantic occupancy prediction results on Occ3D-nuScenes benchmark.
-All methods are trained with dense occupancy labels from Occ3D [27].
-The camera visibility mask is not used during the training phase.
 Notion of modality: Camera (C), LiDAR (L), Radar (R).
 The best and second-best performance is indicated in bold red and bold blue, respectively.
-IV-F Challenging Scenarios Performance Analysis
-To further examine the performance and robustness of the proposed algorithm, we follow [18] to evaluate TFusionOcc with different T-primitive representations under challenging rainy and night-time scenarios and the experiment results are shown in Table IV and Table V, respectively.
-In the rainy scenario, raindrops on the camera lens cause focus blur and degrade the quality of semantic information provided by the camera branch.
-Furthermore, laser beam reflections during rainy conditions degraded the quality of the geometry information provided by the lidar branch.
-As a result, all algorithms in the benchmark exhibit varying degrees of performance degradation.
-However, our proposed method benefits from a multi-stage feature fusion design and the flexible T-Primitive representation, exhibits the least performance degradation in the benchmark, demonstrating high robustness against challenging rainy conditions.
-In the nighttime scenario, due to the nature of the camera sensor, which is sensitive to the ambient lightning conditions, the dim lighting condition during the nighttime leads to semantic information loss in the camera branch.
-Meanwhile, the depth-estimation accuracy from the camera branch drops.
-As a result, all algorithms in the benchmark experience severe performance degradation around 1.56% to 11.10% under the IoU metric and 8.89% to 12.50% under the mIoU metric.
-Among all performance degradation algorithms, our method still achieves SOTA performance, demonstrating superior feature fusion and information filtering capabilities.
-Method Backbone Modality IoU mIoU $\bullet$ barrier $\bullet$ bicycle $\bullet$ bus $\bullet$ car $\bullet$ const.
-veh.
-$\bullet$ motorcycle $\bullet$ pedestrian $\bullet$ traffic cone $\bullet$ trailer $\bullet$ truck $\bullet$ drive.
-surf.
-$\bullet$ other flat $\bullet$ sidewalk $\bullet$ terrain $\bullet$ manmade $\bullet$ vegetation InverseMatrixVT3D [17] R101-DCN C 29.72 (-0.31%) 18.99 (+0.11%) 18.55 14.29 22.28 30.02 10.19 15.20 10.03 9.71 13.28 20.98 37.18 23.47 27.74 17.46 10.36 23.13 GaussianFormer [9] R101-DCN C 27.37 (-2.46%) 16.96 (-2.14%) 18.16 9.58 21.09 26.83 8.04 10.13 7.80 5.84 12.66 18.24 35.53 18.51 27.79 19.23 11.04 20.85 Co-Occ [21] R101 C 28.90 (-1.10%) 19.70 (-0.60%) 22.10 17.60 26.30 30.80 10.90 9.90 8.20 9.70 11.40 19.30 39.00 22.20 32.60 23.00 11.50 21.30 GaussianFormer2-256 [7] R101-DCN C 31.14 (-0.60%) 20.36 (-0.46%) 19.84 13.52 26.89 31.65 10.82 15.16 9.04 8.41 13.72 21.84 40.51 24.57 32.21 20.65 12.64 24.33 SurroundOcc [30] R101-DCN C 30.57 (-0.92%) 19.85 (-0.45%) 21.40 12.75 25.49 31.31 11.39 12.65 8.94 9.48 14.51 21.52 35.34 25.32 29.89 18.37 14.44 24.78 Inverse++ [19] R101-DCN C 31.32 (-0.41%) 20.66 (-0.25%) 22.52 13.79 25.49 31.80 11.70 16.72 11.14 10.12 12.29 22.25 38.78 23.93 31.62 21.14 12.65 24.61 Co-Occ [21] R101 C+L 40.30 (-0.80%) 26.60 (-0.50%) 26.60 19.10 37.60 37.20 15.90 20.30 16.30 12.30 23.30 27.00 41.00 22.80 35.20 24.60 27.80 39.30 OccFusion [18] R101-DCN C+L 42.67 (-0.86%) 26.68 (-0.87%) 20.91 18.39 35.26 36.19 17.69 19.05 19.40 17.08 23.86 28.86 38.28 26.37 31.44 21.35 29.48 43.22 OccFusion [18] R101-DCN C+L+R 42.67 (-1.29%) 27.39 (-0.88%) 27.82 21.10 36.00 37.10 17.23 21.67 20.34 17.46 20.93 28.57 38.99 24.72 31.96 21.26 29.64 43.53 GaussianFusionOcc [23] R101-DCN C+L 44.28 (-0.88%) 29.19 (-1.02%) 28.10 19.84 36.28 38.90 18.11 21.13 26.14 17.95 25.79 29.92 41.72 27.35 34.99 22.85 33.84 44.10 GaussianFusionOcc [23] R101-DCN C+L+R 44.36 (-0.84%) 29.86 (-0.51%) 28.40 19.88 38.87 39.33 20.61 26.05 25.66 17.97 26.07 31.02 41.70 24.94 35.27 24.08 33.90 44.00 OccCylindrical [16] Res-50 C+L 44.08 (-0.86%) 28.07 (-0.60%) 24.50 22.66 33.47 36.79 16.96 23.63 23.79 23.49 22.70 28.14 41.67 22.32 33.55 21.44 30.58 43.45 DAOcc [33] Res-50 C+L 44.51 (-0.49%) 29.65 (-0.85%) 28.50 19.63 38.01 38.98 21.42 20.19 24.63 21.52 23.60 31.12 40.56 26.68 35.05 24.94 34.29 45.28 TFusionOcc (T-SQ-12800) Res-50 C+L 45.79 (-0.91%) 30.28 (-0.48%) 29.32 22.06 35.67 38.47 18.43 23.29 26.19 24.65 23.34 28.40 41.79 28.40 34.99 26.16 35.75 47.54 TFusionOcc (T-SQ-IW-12800) Res-50 C+L 45.81 (-1.11%) 30.74 (-0.47%) 30.54 21.85 37.99 39.11 20.43 21.56 26.44 24.82 23.39 28.74 41.99 28.68 35.70 26.86 36.08 47.63 TFusionOcc (T-P-25600) Res-50 C+L 44.77 (-0.72%) 29.22 (-0.82%) 27.86 21.11 36.24 38.33 17.79 18.33 25.25 23.53 25.50 28.59 40.27 24.95 33.54 24.82 34.90 46.57 TFusionOcc (T-SQ-25600) Res-50 C+L 45.40 (-0.74%) 30.38 (-0.50%) 29.18 22.52 36.87 38.61 19.41 22.19 26.98 24.84 24.78 28.70 41.21 27.39 34.74 25.57 35.69 47.44 TFusionOcc (T-SQ-IW-25600) Res-50 C+L 46.25 (-0.76%) 30.82 (-0.65%) 30.38 21.70 38.23 39.28 20.29 19.78 26.75 25.35 24.88 29.05 42.35 28.14 35.91 26.52 36.49 48.01
-TABLE IV: 3D semantic occupancy prediction results on SurroundOcc-nuScenes validation rainy scenario subset.
-All methods are trained with dense occupancy labels from SurroundOcc [30].
-Notion of modality: Camera (C), Lidar (L), Radar (R).
+In both benchmarks, TFusionOcc (T-SQ-IW) settings achieve the best and second-best performance, demonstrating the importance of the primitives’ shape variety and robustness.
+Specifically, our proposed method excels at detecting background objects, including drivable surfaces, other flat areas, sidewalks, terrain, man-made structures, and vegetation.
+In addition, our model performs significantly well on relatively small foreground objects, such as barriers, bicycles, cars, motorcycles, pedestrians, and traffic cones.
+Notably, our method, which only adopts a single 3D supervision signal, achieved par-performance in comparison with algorithms that leverage multi-task learning to receive dual supervision signals, such as DAOcc [35] in the SurroundOcc-nuScenes benchmark that incorporates an auxiliary 3D object detection head to leverage an extra 3D supervision signal during training, and SDGOcc [4] in the Occ3D-nuScenes benchmark, which incorporates an auxiliary 2D semantic segmentation head to utilize an extra 2D supervision signal during training.
+It is also worth noting that, as highlighted in the green and red rows in both benchmarks, for the same total number of primitive settings, the primitive with greater shape variety achieves better performance.
+In addition, increasing the number of primitives from 12,800 to 25,600 usually yields further gains, though the improvement is moderate compared with that from stronger primitive geometry.
+This suggests that primitive shape flexibility is a more decisive factor than simply increasing primitive count.
+To further examine the performance and robustness of the proposed algorithm, we follow [18] to evaluate TFusionOcc in settings under challenging rainy and night-time scenarios, and the experiment results are shown in Table IV.
+Method Backbone Modality Rainy IoU / mIoU Night IoU / mIoU Co-Occ [21] R101 C+L 40.30 / 26.60 35.60 / 14.60 OccFusion [18] R101-DCN C+L+R 42.67 / 27.39 41.01 / 16.61 GaussianFusionOcc [23] R101-DCN C+L 44.28 / 29.19 42.78 / 18.66 GaussianFusionOcc [23] R101-DCN C+L+R 44.36 / 29.86 42.51 / 18.45 OccCylindrical [16] Res-50 C+L 44.08 / 28.07 43.38 / 17.79 DAOcc [35] Res-50 C+L 44.51 / 29.65 42.85 / 18.53 TFusionOcc (T-SQ-12800) Res-50 C+L 45.79 / 30.28 44.74 / 18.81 TFusionOcc (T-SQ-IW-12800) Res-50 C+L 45.81 / 30.74 44.82 / 19.07 TFusionOcc (T-P-25600) Res-50 C+L 44.77 / 29.22 43.25 / 17.83 TFusionOcc (T-SQ-25600) Res-50 C+L 45.40 / 30.38 43.86 / 18.95 TFusionOcc (T-SQ-IW-25600) Res-50 C+L 46.25 / 30.82 44.97 / 19.67
+TABLE IV: 3D semantic occupancy prediction results on SurroundOcc-nuScenes validation rainy and nighttime scenario subset.
 The best and second-best performance is indicated in bold red and bold blue, respectively.
-Method Backbone Modality IoU mIoU $\bullet$ barrier $\bullet$ bicycle $\bullet$ bus $\bullet$ car $\bullet$ const.
-veh.
-$\bullet$ motorcycle $\bullet$ pedestrian $\bullet$ traffic cone $\bullet$ trailer $\bullet$ truck $\bullet$ drive.
-surf.
-$\bullet$ other flat $\bullet$ sidewalk $\bullet$ terrain $\bullet$ manmade $\bullet$ vegetation InverseMatrixVT3D [17] R101-DCN C 22.48 (-7.55%) 9.99 (-8.89%) 10.40 12.03 0.00 29.94 0.00 9.92 4.88 0.91 0.00 17.79 29.10 2.37 10.80 9.40 8.68 13.57 GaussianFormer [9] R101-DCN C 20.30 (-9.53%) 9.07 (-10.03%) 6.11 8.70 0.00 25.75 0.00 10.44 2.85 0.55 0.00 17.26 30.65 2.95 12.53 9.94 6.65 10.71 Co-Occ [21] R101 C 18.90 (-11.10%) 9.40 (-10.90%) 4.50 9.30 0.00 29.50 0.00 8.40 3.50 0.00 0.00 15.10 29.40 0.60 12.40 11.50 10.70 15.50 GaussianFormer2-256 [7] R101-DCN C 21.19 (-10.55%) 10.14 (-10.68%) 5.25 9.29 0.00 29.33 0.00 13.65 5.80 0.90 0.00 20.22 31.80 1.94 14.83 10.48 5.96 12.72 SurroundOcc [30] R101-DCN C 24.38 (-7.11%) 10.80 (-9.50%) 10.55 14.60 0.00 31.05 0.00 8.26 5.37 0.58 0.00 18.75 30.72 2.74 12.39 11.53 10.52 15.77 Inverse++ [19] R101-DCN C 23.70 (-8.03%) 10.93 (-9.98%) 8.87 10.19 0.00 32.62 0.00 11.77 7.46 0.72 0.00 22.20 32.95 2.15 13.01 9.79 8.61 14.48 Co-Occ [21] R101 C+L 35.60 (-5.50%) 14.60 (-12.50%) 8.40 16.40 0.00 37.20 0.00 13.80 10.90 0.40 0.00 24.30 36.40 2.20 14.60 17.00 19.90 31.70 OccFusion [18] R101-DCN C+L 40.87 (-2.66%) 15.87 (-11.68%) 13.28 17.53 0.00 36.42 0.00 16.16 11.37 1.42 0.00 25.71 32.64 0.63 16.06 20.87 24.52 37.27 OccFusion [18] R101-DCN C+L+R 41.01 (-2.95%) 16.61 (-11.66%) 15.70 16.26 0.00 38.09 0.00 22.18 13.24 0.08 0.00 25.92 33.15 1.57 16.08 21.09 24.51 37.83 GaussianFusionOcc [23] R101-DCN C+L 42.78 (-2.38%) 18.66 (-11.55%) 16.09 12.27 0.00 39.82 0.00 27.66 13.68 0.07 0.00 38.25 40.10 2.07 19.64 19.82 29.55 39.61 GaussianFusionOcc [23] R101-DCN C+L+R 42.51 (-2.69%) 18.45 (-11.92%) 12.15 11.47 0.00 39.77 0.00 29.58 15.27 0.04 0.00 37.08 37.13 2.58 19.94 20.63 29.42 40.10 OccCylindrical [16] Res-50 C+L 43.38 (-1.56%) 17.79 (-10.88%) 16.19 10.04 0.00 37.84 0.00 25.63 13.28 0.23 0.00 32.93 39.59 3.89 17.89 21.61 26.99 38.47 DAOcc [33] Res-50 C+L 42.85 (-2.15%) 18.53 (-11.97%) 17.18 15.07 0.00 39.05 0.00 25.77 14.70 0.07 0.00 32.51 33.95 4.02 19.78 22.58 30.91 40.87 TFusionOcc (T-SQ-12800) Res-50 C+L 44.74 (-1.96%) 18.81 (-11.95%) 16.01 17.51 0.00 39.84 0.00 23.82 12.76 1.13 0.00 32.79 37.67 3.48 19.70 21.42 31.41 43.40 TFusionOcc (T-SQ-IW-12800) Res-50 C+L 44.82 (-2.10%) 19.07 (-12.14%) 16.81 15.53 0.00 40.36 0.00 21.27 13.33 3.79 0.00 34.89 38.04 3.92 19.86 21.89 31.67 43.72 TFusionOcc (T-P-25600) Res-50 C+L 43.25 (-2.24%) 17.83 (-12.21%) 14.12 10.10 0.00 38.45 0.00 21.99 13.10 1.16 0.00 32.05 34.60 4.61 18.91 23.14 30.76 42.35 TFusionOcc (T-SQ-25600) Res-50 C+L 43.86 (-2.28%) 18.95 (-11.93%) 17.10 16.09 0.00 40.17 0.00 23.21 15.13 2.97 0.00 33.83 36.18 3.75 20.12 20.48 31.30 42.94 TFusionOcc (T-SQ-IW-25600) Res-50 C+L 44.97 (-2.04%) 19.67 (-11.80%) 17.98 19.96 0.00 40.69 0.00 23.30 14.76 3.85 0.00 34.49 38.35 3.43 19.70 22.42 31.95 43.79
-TABLE V: 3D semantic occupancy prediction results on SurroundOcc-nuScenes validation night scenario subset.
-All methods are trained with dense occupancy labels from SurroundOcc [30].
-Notion of modality: Camera (C), Lidar (L), Radar (R).
-The best and second-best performance is indicated in bold red and bold blue, respectively.
-IV-G Performance Analysis On nuScenes-C Dataset
-To further assess the robustness of the proposed algorithm in extreme conditions, such as snow, fog, and sensor malfunctions, we use the nuScenes-C dataset and evaluate TFusionOcc at different levels of corruption and severity.
-The results of the experiment for several camera corruption settings are shown in Tables VI, VII, and X, and the results for several lidar corruption settings are shown in Tables VIII, IX, and XI.
-Note that since the nuScenes-C dataset only provides single-lidar-sweep corruption data, and our model leverages 10 lidar sweeps.
-Hence, when evaluating our model across different lidar corruption settings, we use only a single corrupted lidar sweep from the nuScenes-C dataset, and the model performance under the clean setting differs from that in Table II due to a much sparser lidar point cloud.
-In the camera corruption scenario, our model shows strong robustness against camera crashes, frame loss, brightness, and fog, as measured by the mIoU metric.
+In both challenging scenarios, our proposed method under the T-SQ-IW setting achieved the best and second-best performance, demonstrating its robustness and the importance of the geometric shape representation capability of primitives.
+IV-E Performance Analysis on nuScenes-C Dataset
+To thoroughly assess the robustness of the proposed algorithm under extreme conditions, such as snow, fog, and sensor malfunctions, we use the nuScenes-C dataset and evaluate TFusionOcc across varying levels of corruption and severity.
+The results of the experiment for several camera corruption settings are shown in Table V.
+Method Clean IoU / mIoU Cam Crash IoU / mIoU Frame Lost IoU / mIoU Color Quant IoU / mIoU Motion Blur IoU / mIoU Bright IoU / mIoU Low Light IoU / mIoU Fog IoU / mIoU Snow IoU / mIoU OccCylindrical 44.94 / 28.67 41.76 / 23.38 40.84 / 21.87 43.64 / 25.38 41.45 / 22.57 44.78 / 27.88 43.48 / 22.17 44.59 / 27.38 38.92 / 18.29 DAOcc 45.00 / 30.50 42.75 / 22.35 42.16 / 20.78 43.88 / 27.46 42.75 / 26.25 43.72 / 26.78 41.25 / 21.15 43.92 / 26.87 40.72 / 19.42 TFusionOcc (T-P-25600) 45.49 / 30.04 42.59 / 22.34 41.97 / 20.78 43.14 / 24.42 40.45 / 20.04 44.81 / 29.03 41.24 / 18.62 44.14 / 27.43 40.58 / 15.62 TFusionOcc (T-SQ-25600) 46.14 / 30.88 42.59 / 23.50 41.93 / 22.04 43.03 / 25.65 38.36 / 20.25 45.41 / 29.92 41.99 / 20.98 44.90 / 28.85 40.09 / 16.33 TFusionOcc (T-SQ-IW-25600) 47.01 / 31.47 42.61 / 24.39 41.77 / 23.03 43.97 / 26.14 38.58 / 21.06 46.31 / 30.58 42.19 / 21.67 45.95 / 29.69 40.37 / 17.45
+TABLE V: 3D semantic occupancy prediction results on nuScenes-C Camera Corruption benchmark. The best and second-best performance is indicated in bold red and bold blue, respectively.
+In the camera corruption scenario, our model shows strong robustness against camera crashes, frame lost, brightness, and fog, as measured by the IoU and mIoU metric.
 Nevertheless, the motion blur and snow scenarios cause significant performance degradation in our model.
 With around 5.04% to 8.43% and 4.91% to 6.64% degradation under the IoU metric, 10.00% to 10.63% and 14.02% to 14.55% degradation under the mIoU metric for the motion blur and snow corruption case, respectively.
-In the lidar corruption scenario, our model is sensitive to lidar point cloud density and quality degradation, demonstrating its firm reliance on lidar point cloud quality.
-Method Metric Clean Cam Crash Frame Lost Color Quant Motion Blur Bright Low Light Fog Snow OccCylindrical IoU 44.94 41.76 (-3.18%) 40.84 (-4.10%) 43.64 (-1.30%) 41.45 (-3.49%) 44.78 (-0.16%) 43.48 (-1.46%) 44.59 (-0.35%) 38.92 (-6.02%) DAOcc IoU 45.00 42.75 (-2.25%) 42.16 (-2.84%) 43.88 (-1.12%) 42.75 (-2.25%) 43.72 (-1.28%) 41.25 (-3.75%) 43.92 (-1.08%) 40.72 (-4.28%) TFusionOcc (T-P-25600) IoU 45.49 42.59 (-2.90%) 41.97 (-3.52%) 43.14 (-2.35%) 40.45 (-5.04%) 44.81 (-0.68%) 41.24 (-4.25%) 44.14 (-1.35%) 40.58 (-4.91%) TFusionOcc (T-SQ-25600) IoU 46.14 42.59 (-3.55%) 41.93 (-4.21%) 43.03 (-3.11%) 38.36 (-7.78%) 45.41 (-0.73%) 41.99 (-4.15%) 44.90 (-1.24%) 40.09 (-6.05%) TFusionOcc (T-SQ-IW-25600) IoU 47.01 42.61 (-4.40%) 41.77 (-5.24%) 43.97 (-3.04%) 38.58 (-8.43%) 46.31 (-0.70%) 42.19 (-4.82%) 45.95 (-1.06%) 40.37 (-6.64%)
-TABLE VI: 3D semantic occupancy prediction IoU results on nuScenes-C Camera Corruption benchmark.
-All methods are evaluated with different corruption type input images provided by nuScenes-C [31] and ground-truth occupancy labels from SurroundOcc [30].
-Method Metric Clean Cam Crash Frame Lost Color Quant Motion Blur Bright Low Light Fog Snow SurroundOcc mIoU 20.30 11.60 (-8.70%) 10.00 (-10.30%) 14.03 (-6.27%) 12.41 (-7.89%) 19.18 (-1.12%) 12.15 (-8.15%) 18.42 (-1.88%) 7.39 (-12.91%) Inverse++ mIoU 20.91 11.76 (-9.15%) 10.22 (-10.69%) 14.88 (-6.03%) 12.84 (-8.07%) 19.71 (-1.20%) 12.05 (-8.86%) 18.86 (-2.05%) 7.62 (-13.29%) OccCylindrical mIoU 28.67 23.38 (-5.29%) 21.87 (-6.80%) 25.38 (-3.29%) 22.57 (-6.10%) 27.88 (-0.79%) 22.17 (-6.50%) 27.38 (-1.29%) 18.29 (-10.38%) DAOcc mIoU 30.50 22.35 (-8.15%) 20.78 (-9.72%) 27.46 (-3.04%) 26.25 (-4.25%) 26.78 (-3.72%) 21.15 (-9.35%) 26.87 (-3.63%) 19.42 (-11.08%) TFusionOcc (T-P-25600) mIoU 30.04 22.34 (-7.70%) 20.78 (-9.26%) 24.42 (-5.62%) 20.04 (-10.00%) 29.03 (-1.01%) 18.62 (-11.42%) 27.43 (-2.61%) 15.62 (-14.42%) TFusionOcc (T-SQ-25600) mIoU 30.88 23.50 (-7.38%) 22.04 (-8.84%) 25.65 (-5.23%) 20.25 (-10.63%) 29.92 (-0.96%) 20.98 (-9.90%) 28.85 (-2.03%) 16.33 (-14.55%) TFusionOcc (T-SQ-IW-25600) mIoU 31.47 24.39 (-7.08%) 23.03 (-8.44%) 26.14 (-5.33%) 21.06 (-10.41%) 30.58 (-0.89%) 21.67 (-9.80%) 29.69 (-1.78%) 17.45 (-14.02%)
-TABLE VII: 3D semantic occupancy prediction mIoU results on nuScenes-C Camera Corruption benchmark.
-All methods are evaluated with different corruption type input images provided by nuScenes-C [31] and ground-truth occupancy labels from SurroundOcc [30].
-Method Metric Clean Beam Missing Cross Sensor Cross Talk Incomplete Echo Motion Blur Wet Ground Fog Snow OccCylindrical IoU 35.89 30.73 (-5.16%) 25.01 (-10.88%) 33.29 (-2.60%) 35.37 (-0.52%) 32.44 (-3.45%) 35.04 (-0.85%) 25.79 (-10.10%) 34.00 (-1.89%) DAOcc IoU 42.39 38.17 (-4.22%) 33.79 (-8.60%) 33.68 (-8.71%) 42.01 (-0.38%) 38.05 (-4.34%) 41.67 (-0.72%) 31.47 (-10.92%) 37.01 (-5.38%) TFusionOcc (T-P-25600) IoU 40.29 32.75 (-7.54%) 25.65 (-14.64%) 36.25 (-4.04%) 39.67 (-0.62%) 34.79 (-5.50%) 38.17 (-2.12%) 26.44 (-13.85%) 36.88 (-3.41%) TFusionOcc (T-SQ-25600) IoU 40.53 33.00 (-7.53%) 25.89 (-14.64%) 36.60 (-3.93%) 39.90 (-0.63%) 34.67 (-5.86%) 38.74 (-1.79%) 26.54 (-13.99%) 36.88 (-3.65%) TFusionOcc (T-SQ-IW-25600) IoU 41.53 34.09 (-7.44%) 26.82 (-14.71%) 37.50 (-4.03%) 40.94 (-0.59%) 35.54 (-5.99%) 39.84 (-1.69%) 27.10 (-14.43%) 37.70 (-3.83%)
-TABLE VIII: 3D semantic occupancy prediction IoU results on nuScenes-C Lidar Corruption benchmark.
-All methods are evaluated with different corruption type input images provided by nuScenes-C [31] and ground-truth occupancy labels from SurroundOcc [30].
-Method Metric Clean Beam Missing Cross Sensor Cross Talk Incomplete Echo Motion Blur Wet Ground Fog Snow OccCylindrical mIoU 24.06 19.83 (-4.23%) 16.20 (-7.86%) 22.60 (-1.46%) 20.93 (-3.13%) 20.85 (-3.21%) 23.31 (-0.75%) 17.32 (-6.74%) 22.38 (-1.68%) DAOcc mIoU 27.73 22.96 (-4.77%) 18.41 (-9.32%) 22.86 (-4.87%) 24.45 (-3.28%) 23.66 (-4.07%) 27.25 (-0.48%) 19.73 (-8.00%) 24.53 (-3.20%) TFusionOcc (T-P-25600) mIoU 26.13 20.49 (-5.64%) 15.38 (-10.75%) 23.76 (-2.37%) 22.69 (-3.44%) 21.39 (-4.74%) 25.05 (-1.08%) 17.44 (-8.69%) 23.15 (-2.98%) TFusionOcc (T-SQ-25600) mIoU 27.03 21.07 (-5.96%) 15.61 (-11.42%) 24.73 (-2.30%) 23.30 (-3.73%) 22.15 (-4.88%) 26.06 (-0.97%) 18.40 (-8.63%) 24.37 (-2.66%) TFusionOcc (T-SQ-IW-25600) mIoU 27.50 21.81 (-5.69%) 16.49 (-11.01%) 24.94 (-2.56%) 24.11 (-3.39%) 22.62 (-4.88%) 26.57 (-0.93%) 18.63 (-8.87%) 24.37 (-3.13%)
-TABLE IX: 3D semantic occupancy prediction mIoU results on nuScenes-C Lidar Corruption benchmark.
-All methods are evaluated with different corruption type input lidar point cloud provided by nuScenes-C [31] and ground-truth occupancy labels from SurroundOcc [30].
-Corruption Severity IoU mIoU $\bullet$ barrier $\bullet$ bicycle $\bullet$ bus $\bullet$ car $\bullet$ const.
-veh.
-$\bullet$ motorcycle $\bullet$ pedestrian $\bullet$ traffic cone $\bullet$ trailer $\bullet$ truck $\bullet$ drive.
-surf.
-$\bullet$ other flat $\bullet$ sidewalk $\bullet$ terrain $\bullet$ manmade $\bullet$ vegetation CameraCrash Easy 43.16 26.44 23.34 16.62 28.80 35.79 19.34 21.30 28.47 19.01 18.46 24.14 36.89 19.19 25.07 27.42 35.32 43.94 CameraCrash Mid 42.27 23.37 17.66 11.56 26.79 33.91 14.58 16.88 26.48 16.44 13.69 19.75 35.87 16.54 22.43 24.70 34.07 42.56 CameraCrash Hard 42.39 23.35 20.30 11.23 26.54 34.40 15.27 13.93 25.80 17.87 11.15 21.39 35.24 14.59 22.44 25.18 34.52 43.77 FrameLost Easy 44.33 27.42 25.40 16.12 30.44 36.89 19.15 21.10 28.59 20.72 18.97 26.33 38.65 21.48 26.20 28.11 35.81 44.84 FrameLost Mid 41.33 22.19 15.55 10.88 24.67 33.26 13.95 14.55 25.62 15.66 11.70 19.94 33.63 13.21 21.59 24.17 33.98 42.73 FrameLost Hard 39.66 19.48 9.91 9.09 22.02 31.26 11.74 11.63 23.98 12.70 8.77 15.30 30.99 9.11 18.80 21.75 33.03 41.52 ColorQuant Easy 46.55 30.61 31.18 19.09 33.46 39.63 22.86 26.15 30.57 23.20 22.87 30.29 42.26 25.40 28.71 30.42 37.19 46.44 ColorQuant Mid 44.19 26.96 25.94 14.03 30.21 38.26 19.43 18.13 29.03 22.00 18.56 26.63 38.54 19.68 24.39 26.26 35.72 44.54 ColorQuant Hard 41.18 20.85 13.87 8.24 22.61 34.82 12.93 8.22 26.70 18.66 9.86 20.29 33.70 9.37 19.14 33.79 42.30 20.85 MotionBlur Easy 41.43 26.89 28.45 15.39 31.42 37.33 20.59 19.98 29.74 23.25 19.18 25.99 36.16 17.03 22.63 25.86 34.73 42.48 MotionBlur Mid 37.26 18.80 15.08 8.51 21.43 33.12 11.83 13.43 22.58 17.93 6.36 13.80 29.91 3.61 12.99 20.31 31.56 38.34 MotionBlur Hard 37.06 17.48 12.18 7.54 19.65 32.08 9.80 12.42 20.60 16.37 4.50 12.63 29.47 2.63 11.26 19.46 31.17 37.93 Brightness Easy 46.90 31.35 31.83 20.44 34.29 39.73 23.55 27.70 31.19 23.89 23.07 30.52 43.06 27.37 29.94 30.92 37.45 46.57 Brightness Mid 46.45 30.67 31.12 19.86 32.69 39.37 22.29 26.38 31.08 23.73 22.70 29.40 42.35 27.07 28.96 30.20 37.26 46.31 Brightness Hard 45.59 29.71 30.29 19.08 31.22 38.80 20.78 25.06 30.92 23.76 21.97 28.14 40.90 25.49 27.29 28.93 36.92 45.88 LowLight Easy 43.86 25.38 19.52 14.11 30.88 37.84 16.55 20.19 28.43 17.38 13.94 25.64 37.39 15.11 23.22 25.82 35.57 44.48 LowLight Mid 42.12 21.69 11.25 11.54 27.65 36.35 11.88 18.08 26.75 12.07 5.65 22.09 34.43 8.43 19.90 23.02 34.45 43.44 LowLight Hard 40.60 17.93 3.52 6.49 23.37 34.05 7.87 15.49 23.81 5.33 1.57 17.02 32.01 3.56 16.69 20.34 33.18 42.59 Fog Easy 46.37 30.57 30.82 19.55 34.30 39.40 21.91 25.99 30.65 23.35 22.12 30.06 42.13 26.77 28.91 29.86 37.10 46.27 Fog Mid 46.03 29.89 30.01 18.93 33.72 39.29 20.89 25.12 30.42 22.91 21.21 29.32 41.37 25.36 27.89 28.93 36.87 46.07 Fog Hard 45.44 28.61 29.00 17.43 32.02 38.90 18.92 24.15 30.04 22.41 19.61 27.90 40.02 22.29 25.69 27.37 36.47 45.58 Snow Easy 41.26 20.46 12.34 8.18 24.01 34.48 14.53 13.38 26.04 20.12 6.13 19.37 32.52 6.58 12.73 21.68 33.02 42.31 Snow Mid 39.61 15.81 7.13 4.17 16.34 30.79 7.63 11.86 22.00 16.87 4.93 9.97 27.63 2.30 7.69 15.13 30.04 38.46 Snow Hard 40.23 16.08 5.29 4.19 16.00 31.17 8.68 11.28 22.39 16.59 4.58 10.08 29.80 2.64 9.17 14.27 31.40 39.78
-TABLE X: The performance of TFusionOcc(T-SQ-IW-25600) on the nuScene-C [31] dataset under different camera corruption and severity settings.
-Corruption Severity IoU mIoU $\bullet$ barrier $\bullet$ bicycle $\bullet$ bus $\bullet$ car $\bullet$ const.
-veh.
-$\bullet$ motorcycle $\bullet$ pedestrian $\bullet$ traffic cone $\bullet$ trailer $\bullet$ truck $\bullet$ drive.
-surf.
-$\bullet$ other flat $\bullet$ sidewalk $\bullet$ terrain $\bullet$ manmade $\bullet$ vegetation BeamMissing Light 38.03 24.83 25.39 12.35 30.38 33.20 17.80 20.84 23.97 16.98 19.35 26.19 37.83 21.87 23.66 24.77 26.92 35.79 BeamMissing Moderate 34.28 21.94 22.45 9.84 28.26 29.26 15.39 17.79 20.83 12.97 17.61 23.96 35.44 19.69 20.99 22.06 23.07 31.41 BeamMissing Heavy 29.97 18.67 18.85 6.93 25.51 24.95 13.02 13.23 15.80 11.40 15.58 21.18 32.63 17.35 18.16 18.45 19.03 26.71 CrossSensor Light 31.71 20.08 19.55 7.95 27.85 27.18 14.63 13.75 17.86 11.99 17.34 23.06 34.50 19.05 19.57 19.78 20.08 27.19 CrossSensor Moderate 26.52 16.07 15.50 5.00 23.42 21.71 12.15 7.42 13.75 9.30 14.33 19.55 31.52 15.85 16.60 16.43 14.97 19.66 CrossSensor Heavy 22.23 13.32 11.50 3.77 21.31 16.56 6.96 9.04 10.43 7.01 13.48 15.76 28.99 14.64 14.00 13.66 11.28 14.64 CrossTalk Light 39.77 26.33 28.50 14.08 31.14 35.44 18.51 21.00 25.73 17.29 19.97 27.30 39.52 24.07 26.08 27.46 29.48 35.63 CrossTalk Moderate 37.47 24.93 27.82 12.62 30.10 34.11 16.63 19.01 24.07 15.91 19.14 26.14 38.85 23.78 25.07 26.82 28.03 30.82 CrossTalk Heavy 35.26 23.57 26.75 11.55 29.19 32.80 15.39 16.61 22.60 14.60 18.47 25.09 38.09 23.34 23.41 25.86 26.44 26.93 IncompleteEcho Light 41.18 25.74 28.97 12.08 28.39 30.45 15.70 19.21 27.68 19.35 17.63 24.15 39.80 23.88 26.37 27.75 30.51 39.94 IncompleteEcho Moderate 41.01 24.66 28.96 10.54 25.72 26.81 13.15 17.27 27.67 19.26 15.40 21.58 39.74 23.90 26.36 27.73 30.52 39.93 IncompleteEcho Heavy 40.62 21.93 28.92 8.66 17.25 16.89 7.41 13.52 27.62 19.08 8.97 14.64 39.57 23.92 26.34 27.71 30.53 39.91 MotionBlur Light 38.10 24.89 25.61 12.99 30.89 33.74 19.39 20.91 24.15 14.99 19.40 26.77 36.30 21.25 23.10 24.59 26.09 38.00 MotionBlur Moderate 35.44 22.56 22.40 11.01 29.50 31.09 18.24 18.10 20.87 11.53 18.13 25.43 33.88 18.98 20.68 22.26 22.74 36.10 MotionBlur Heavy 33.07 20.42 19.24 9.15 28.01 28.37 17.26 15.57 17.81 8.80 16.49 23.64 32.07 17.44 18.77 20.20 19.85 34.03 WetGround Light 40.53 26.99 28.86 15.18 31.66 36.33 19.80 23.18 27.60 19.65 20.73 28.03 37.07 21.62 23.82 27.84 30.50 39.99 WetGround Moderate 39.84 26.57 28.86 15.12 31.65 36.23 19.77 22.70 27.53 19.86 20.71 27.91 35.13 19.21 21.98 28.01 30.50 39.98 WetGround Heavy 39.14 26.14 28.80 14.79 31.57 36.14 19.74 22.69 27.47 19.87 20.68 27.85 33.08 17.10 20.01 27.99 30.49 39.98 Fog Light 36.86 24.38 27.36 14.09 28.96 33.78 15.06 20.04 23.85 18.70 17.81 23.49 38.44 22.32 24.71 23.95 25.84 31.69 Fog Moderate 28.39 20.05 22.93 12.52 26.29 29.36 12.23 17.41 19.34 16.90 11.32 16.80 34.83 20.18 21.50 19.90 16.92 22.37 Fog Heavy 16.04 11.46 13.35 6.10 15.92 19.10 7.93 8.75 9.78 10.17 3.34 7.81 26.48 14.66 12.86 11.56 6.17 9.35 Snow Light 39.38 24.78 28.29 14.52 29.78 34.99 14.71 19.91 23.43 17.00 18.80 25.34 39.33 23.29 25.23 26.10 27.60 28.23 Snow Moderate 37.52 24.42 28.34 14.26 29.77 34.75 14.91 19.96 22.36 15.96 18.51 24.98 38.93 23.35 25.34 26.44 27.15 25.73 Snow Heavy 36.20 23.91 28.11 13.47 29.74 34.20 15.34 18.98 21.07 13.80 18.17 25.16 38.42 23.08 25.28 26.47 25.68 25.65
-TABLE XI: The performance of TFusionOcc(T-SQ-IW-25600) on the nuScene-C [31] dataset under different lidar corruption and severity settings.
-IV-H Performance Analysis On Varying Distance
-IV-H1 Performance Variation Trend At Different Distances.
-Following [18], we take the ego vehicle as the origin and $R$ as the radius.
-By adjusting the length of $R$, we evaluate our proposed method against other SOTA algorithms at varying perception distance radius [20m, 25m, 30m, 35m, 40m, 45m, 50m], including challenging rainy and night scenarios.
-The trend of performance variation is shown in Figure 10.
-The trend of performance variation across the whole validation set, as shown in Figures 10(a) and 10(d), aligns with the performance benchmark in Table II.
-In the rainy scenario, the performance variation trend of our proposed method is similar to the results of the whole validation set experiment, except that the performance gap between different T-Primitive settings further narrows down to 0.11% to 2.90% under the mIoU metric and 0.44% to 1.74% under the IoU metric due to the semantic and geometric information loss from both sensors.
-In the nighttime scenario, we discover that our T-SQ-25600 setting outperforms the second-best T-SQ-IW-12800 setting within a 35-meter perception range under the mIoU metric, as shown in Figure 10(c).
-This phenomenon may stem from the fact that more primitives can better capture close-range details, leading to better robustness against challenging dim-lighting conditions.
-Figure 10: Performance variation trend for 3D semantic occupancy prediction task.
-(a) mIoU performance variation trend on the whole SurroundOcc-nuScenes validation set, (b) IoU performance variation on the whole SurroundOcc-nuScenes validation set, (c) mIoU performance variation trend on the SurroundOcc-nuScenes validation rainy scenario subset, (d) IoU performance variation on the SurroundOcc-nuScenes validation rainy scenario subset, (e) mIoU performance variation on the SurroundOcc-nuScenes validation night scenario subset, and (f) IoU performance variation on the SurroundOcc-nuScenes validation night scenario subset.
-Better viewed when zoomed in.
-IV-H2 Performance In Different Sectors Ranges
-[…]
+This phenomenon could stem from the multi-stage feature fusion design strategy.
+Since frame lost and cam crash, both corruption results in the complete loss of information from the camera side.
+Our feature fusion module at each stage can easily filter out unnecessary visual features and rely solely on LiDAR features for occupancy prediction.
+As for brightness, fog, and low-light corruption, these conditions primarily degrade camera-side information quality and can be compensated for on the LiDAR side through our multi-stage feature fusion design.
+However, motion blur corruption causes structured information loss via convolution, bringing in inaccurate visual features into the feature fusion stage, resulting in a low-quality fused feature.
+Similarly, snow corruption introduces additional salt-and-pepper noise into the multi-stage feature fusion pipeline, resulting in a low-quality fused feature.
+Nevertheless, we consistently observe that more expressive primitives achieve stronger robustness under most corruption settings, following the trend T-SQ-IW >> T-SQ >> T-P, which further highlights the importance of geometric shape flexibility.
+IV-F Performance Analysis on Varying Distance
 […]
@@
 Better viewed when zoomed in.
-IV-I Qualitative Study
-IV-I1 Qualitative Results On SurroundOcc-nuScenes Dataset
-To qualitatively study the difference in each setting of the proposed TFusionOcc, we demonstrate the visualization result of each setting in daytime, rainy, and night-time scenarios in Figure 12, with the main discrepancy area highlighted.
-In the daytime scenarios as shown at the top of Figure 12, in the orange square highlighted area, except for the T-P-25600 setting, all other settings successfully detect those two pedestrians.
-What is more, T-SQ-IW Primitive preserves better the geometric contour shape of background buildings, such as sharper building edges and more accurate building boundaries, than T-P primitive and T-SQ primitive, demonstrating the importance of the geometric shape flexibility of intermediate primitives.
-In rainy scenarios, as shown in the middle of Figure 12, in the dark red square highlighted area.
-The better geometric shape representation capability of the primitives could more accurately capture the geometric structure of the object, leading to better detection performance; T-SQ-IW-25600 achieved the best performance due to its best geometric shape representation capability, and T-P-25600 the worst, stemming from the ellipsoid geometric shape prior.
-Similarly, in the nighttime scenarios, as shown in the bottom of Figure 12, in the square highlighted area.
-Even with severe occlusions and a vehicle headlight glare pollution problem, we observed that greater geometric shape flexibility of primitives leads to better detection performance, which aligns with rainy and daytime phenomena.
-Figure 12: Qualitative results under different settings for daytime, rainy, and nighttime scenarios displayed in the upper, middle, and bottom sections, respectively. Better viewed when zoomed in.
-Furthermore, to demonstrate the effectiveness and robustness of our method, we compare TFusionOcc in three different T-Primitive settings, such as T-P-25600, T-SQ-25600 and T-SQ-IW-25600, against other SOTA algorithms in challenging daytime, rainy and nighttime scenarios, as shown in Figure 13, with the primary difference area highlighted in each scenario.
-In the daytime scenario, as shown at the top of Figure 13, in the orange square highlighted area, benefit from multi-task learning, which provides an additional supervision signal, DAOcc successfully detects distant pedestrians and achieves similar performance as TFusionOcc.
-However, in the dark-red-square-highlighted area, DAOcc fails to detect the distant vehicle, demonstrating the superior performance of our proposed method.
-In the rainy scenario, as shown in the middle of Figure 13, in the purple square highlighted area, our proposed method not only successfully detects all dynamic objects, but also preserves their general contours well.
-In the highlighted dark-blue-square area, our approach achieves par performance with other SOTA algorithms for detecting challenging pedestrians.
-In the nighttime scenario, as shown at the bottom of Figure 13, our proposed method successfully detects very distant dynamic objects in the highlighted dark-red-square area, demonstrating its superior long-range dynamic object perception capability.
+[…]
+IV-G Qualitative Study
+To demonstrate the effectiveness and robustness of our method, we compare TFusionOcc in three different T-Primitive settings, such as T-P-25600, T-SQ-25600 and T-SQ-IW-25600, against other SOTA algorithms in challenging daytime, rainy and nighttime scenarios, as shown in Fig.
+10, with the primary difference area highlighted in each scenario.
 Figure 10: Qualitative results compare against other SOTA algorithms for daytime, rainy, and nighttime scenarios displayed in the upper, middle, and bottom sections, respectively.
 Better viewed when zoomed in.
-IV-I2 Qualitative Results On nuScenes-C Dataset
-Figure 14: Qualitative results compare against other SOTA algorithms under fog and snow scenarios, displayed in the top and bottom sections, respectively. Better viewed when zoomed in.
-Since the nuScenes dataset does not contain fog and snow scenarios, which are common driving conditions in the real world, we adopted the nuScenes-C dataset to evaluate our model across three settings within the same scene under fog and snow conditions, with the severity of each condition set to hard, and compare the visualization results with other SOTA algorithms.
-The visualization results are shown in Figure 14, with the main area of the discrepancy highlighted in dark blue and dark red squares.
-Under foggy conditions, the surround-view images suffer significant visibility degradation due to strong atmospheric scattering, leading to reduced illumination, low contrast, and severely blurred object boundaries, especially in distant regions of the scene.
-Compared to baseline SOTA algorithms, our method, especially the T-SQIW-25600 setting, successfully detects all dynamic pedestrian objects in the scene (as shown by the dark blue squares) and also accurately captures the background objects’ contour as well (as shown by the dark red squares), demonstrating the robustness of the proposed algorithm under foggy conditions.
-Nevertheless, in the snow condition, the snowfall captured in the surround-view image appears to introduce salt-and-pepper noise to the image, and snowflakes occlude some objects’ contours, even the main body parts.
-In this case, our proposed method is less tolerant of salt-and-pepper noise than the baseline SOTA algorithms.
-IV-J Model Efficiency
-Method Modality Latency (ms) ($\downarrow$) Memory (GB) ($\downarrow$) Params SurroundOcc* [30] C 472 5.98 180.51M InverseMatrixVT3D* [17] C 447 4.41 67.18M OccFusion(C+R)* [18] C+R 588 5.56 92.71M OccFusion(C+L)* [18] C+L 591 5.56 92.71M OccFusion(C+L+R)* [18] C+L+R 601 5.78 114.97M OccCylindrical* [16] C+L 398 10.63 111.62M TFusionOcc (T-SQ-12800) C+L 278.46 3.68 95.62M TFusionOcc (T-SQ-IW-12800) C+L 281.47 3.88 95.96M TFusionOcc (T-P-25600) C+L 385.49 4.73 97.51M TFusionOcc (T-SQ-25600) C+L 388.05 4.13 97.59M TFusionOcc (T-SQ-IW-25600) C+L 391.06 4.04 98.24M
-TABLE XII: Efficiency comparison of multi-modal 3D semantic occupancy prediction on the nuScenes validation set.
-Algorithm results marked with * were taken from the paper that introduced the model and were measured on a different GPU.
-We evaluated the real-time performance of our model in various settings on an RTX 4090 GPU and presented the results in Table XII.
-The experiment results show the trade-off between the total number of primitives and the model inference speed.
-More primitives require more rendering operations, leading to higher latency.
-Surprisingly, all five settings have similar total model parameters.
-Based on the main performance benchmark shown in Table II, the T-SQ-IW-12800 setting delivers the best balance between detection and real-time performance.
-IV-K Ablation Study
-IV-K1 Ablation Study On Skeleton Merge Module
-Main-Skele Augment Overlap Remove Range Filter IoU ($\uparrow$) mIoU ($\uparrow$) 46.24 (Baseline) 30.79 (Baseline) 46.20 (-0.04%) 30.70 (-0.09%) 46.07 (-0.17%) 30.76 (-0.03%) 46.08 (-0.16%) 30.69 (-0.1%) 41.96 (-4.28%) 29.92 (-0.87%) 29.35 (-16.89%) 22.31 (-8.48%)
-TABLE XIII: Ablation study on the Skeleton Merge Module. Main-Skele: Lidar-side main skeleton branch. Augment: Camera-side augment skeleton branch. $\uparrow$:the higher, the better.
+In the daytime scenario, as shown at the top of Fig.
+10, in the orange square highlighted area, benefit from multi-task learning, which provides an additional supervision signal, DAOcc successfully detects distant pedestrians and achieves similar performance as TFusionOcc.
+However, in the dark-red-square-highlighted area, DAOcc fails to detect the distant vehicle, demonstrating the superior performance of our proposed method.
+In the rainy scenario, as shown in the middle of Fig.
+10, in the purple square highlighted area, our proposed method not only successfully detects all dynamic objects, but also preserves their general contours well.
+In the highlighted dark-blue-square area, our approach achieves par performance with other SOTA algorithms for detecting challenging pedestrians.
+In the nighttime scenario, as shown at the bottom of Fig.
+10, our proposed method successfully detects very distant dynamic objects in the highlighted dark-red-square area, demonstrating its superior long-range dynamic object perception capability.
+IV-H Model Efficiency
+Method Modality Latency (ms) ($\downarrow$) Memory (GB) ($\downarrow$) Params M-CONet [30] C+L 416.00 12.10 125.43M OccCylindrical [16] C+L 361.39 5.97 111.62M OccMamba-384 [10] C+L 448.30 12.1 111.08M OccMamba-128 [10] C+L 269.30 6.8 64.58M TFusionOcc (T-SQ-12800) C+L 278.46 3.68 95.62M TFusionOcc (T-SQ-IW-12800) C+L 281.47 3.88 95.96M TFusionOcc (T-P-25600) C+L 385.49 4.73 97.51M TFusionOcc (T-SQ-25600) C+L 388.05 4.13 97.59M TFusionOcc (T-SQ-IW-25600) C+L 391.06 4.04 98.24M
+TABLE VI: Model efficiency comparison of different methods.
+The experiments are performed on a single RTX4090 using six multi-camera images and 10 sweeps of LiDAR.
+For input image resolution, all methods adopt $1600\times 900$.
+We evaluated the real-time performance of our model across various settings, comparing it with several other SOTA approaches on an RTX 4090 GPU, and presented the results in Table VI.
+The experiment results show the trade-off among the total number of primitives, memory consumption, and model inference speed.
+More primitives require more rendering operations and memory, leading to higher latency but better performance.
+Based on the main performance benchmark shown in Table II, the T-SQ-IW-12800 setting delivers the best balance between prediction and real-time performance.
+IV-I Ablation Study
+IV-I1 Ablation Study on Different Distribution and Primitives
+Prob Distribution Query Number IoU ($\uparrow$) Gain ($\uparrow$) mIoU ($\uparrow$) Gain ($\uparrow$) T-P / G-P 25600 44.77 / 44.71 – / – 29.22 / 29.20 – / – T-SQ / G-SQ 25600 45.40 / 44.51 +0.63 / -0.20 30.38 / 30.25 +1.16 / +1.05 T-SQ-IW / G-SQ-IW 25600 46.23 / 45.58 +0.83 / +1.07 30.77 / 30.66 +0.39 / +0.41
+TABLE VII: Ablation study on different probability distribution and primitives.
+Gain denotes the vertical performance change within each series (T-series / G-series).
+T refers to the Student’s t-distribution kernel, and G refers to the Gaussian kernel.
+Table VII presents the ablation study results for different probability distributions and primitive settings.
+For each row in the table, for the same type of primitive and the same total number of primitives, with different distribution settings, the model obtained better performance, benefiting from the robustness property of the Student’s t-distribution.
+Notably, the geometric prediction capability benefits more than the semantic prediction capability.
+What’s more, by examining the table vertically, we can see that, regardless of the probability distribution kernel setting, the primitive with greater shape variety performs better, demonstrating the importance of the primitive’s geometric shape representation capability.
+IV-I2 Ablation Study on Basis Field of T-SQ-IW Primitive
+Basis field IoU ($\uparrow$) mIoU ($\uparrow$) $B_{1}\sim B_{24}$ 46.23 (Baseline) 30.77 (Baseline) $B_{10}\sim B_{24}$ 32.98 (-13.25%) 23.17 (-7.60%) $B_{1}\sim B_{18}$ 36.20 (-10.03%) 24.11 (-6.66%) $B_{1}\sim B_{9}$ 21.83 (-24.40%) 11.60 (-19.17%) No Inverse-Warp 20.24 (-25.99%) 13.87 (-16.90%)
+TABLE VIII: Ablation study on basis field in T-SQ-IW primitive.
+$B_{10}\sim B_{24}$: primitive loses basic constant, linear, shear and taper deformation capability.
+$B_{1}\sim B_{18}$: primitive loses quadratic, radial bulge and smooth corner deformation capability.
+$B_{1}\sim B_{9}$: primitive further loses shear, twist and bending deformation capability.
+Table VIII presents the results of the ablation study for each set of basis fields in the primitive T-SQ-IW.
+Since $B_{19}\sim B_{24}$ provides the highest level deformation, including quadratics, radial bulge, and smooth corner deformation, losing those basis fields causes 10.04% IoU and 6.66% mIoU degradation.
+In addition, because $B_{10}\sim B_{18}$ provides the second-highest level of deformation, including full shear, twist, and bending, losing $B_{10}\sim B_{24}$ results in performance drop.
+Surprisingly, the performance drop after losing all basis fields is similar to the loss $B_{10}\sim B_{24}$.
+This result is very likely because the remaining $B_{1}\sim B_{9}$ only provide basic constant, linear, shears, and taper deformation, which leads to limited improvement in the shape variety of the primitive.
+This result also demonstrates the importance of the high-level deformation basis field, which significantly improved the shape variety of the primitive, resulting in better overall 3D occupancy prediction performance.
+IV-I3 Ablation Study on Skeleton Merge Module
+Main-Skele Augment Overlap Remove Range Filter IoU ($\uparrow$) mIoU ($\uparrow$) 46.24 (Baseline) 30.79 (Baseline) 46.20 (-0.04%) 30.70 (-0.09%) 46.07 (-0.17%) 30.76 (-0.03%) 46.08 (-0.16%) 30.69 (-0.10%) 41.96 (-4.28%) 29.92 (-0.87%) 29.35 (-16.89%) 22.31 (-8.48%)
+TABLE IX: Ablation study on the Skeleton Merge Module. Main-Skele: LiDAR-side main skeleton branch. Augment: Camera-side augment skeleton branch.
 Range Filter $r$ IoU ($\uparrow$) mIoU ($\uparrow$) 5 m 46.24 (Baseline) 30.79 (Baseline) 4 m 46.21 (-0.03%) 30.75 (-0.04%) 3 m 46.09 (-0.15%) 30.76 (-0.03%) 2 m 45.72 (-0.52%) 30.69 (-0.10%) 1 m 44.64 (-1.60%) 30.32 (-0.47%) 0.5 m 43.60 (-2.64%) 30.04 (-0.75%)
-TABLE XIV: Ablation study on range filter hyperparameter $r$. $\uparrow$:the higher, the better.
-We conducted an ablation study for each component of the skeleton merge module and demonstrated the results of the experiment in Table XIII.
-The experiment results prove the significant impact of the main skeleton structure, which supports the fundamental 29.92% mIoU and 41.96% IoU performance, and the auxiliary impact of the augment structure, which further boosts 0.87% mIoU and 4.28% IoU performance.
-The main skeleton, based on lidar, provides detailed 3D geometric information of the surrounding scene, and the augmented skeleton, based on the camera, provides additional information to mitigate occlusion impact in the scene.
-Furthermore, we studied the impact of the range filter hyperparameter $r$ on the overall performance of the model and present the results in Table XIV.
-The results show that a smaller filter range forces the augmented skeleton to get too close to the main skeleton, preventing it from overcoming the lidar sensor’s occlusion limitation.
-IV-K2 Ablation Study On Depth-Fusion-Based 3D Deformable Attention Module
+TABLE X: Ablation study on range filter hyperparameter $r$.
+We conducted an ablation study for each component of the skeleton merge module and presented the results in Table IX.
+The experiment results prove the significant impact of the main skeleton structure, which supports the fundamental 29.92% mIoU and 41.96% IoU performance, and the auxiliary impact of the augmented skeleton, which further boosts 0.87% mIoU and 4.28% IoU performance.
+The main skeleton, based on LiDAR, provides precise 3D geometric information about the surrounding scene, especially the front faces of objects, and the augmented skeleton, based on the camera, provides additional information to mitigate the partial occlusion.
+Furthermore, we studied the impact of the range filter hyperparameter $r$ on the model’s overall performance and present the results in Table X.
+The results show that a smaller filter range forces the augmented skeleton to get too close to the main skeleton, preventing it from overcoming the LiDAR sensor’s occlusion limitation, resulting in performance degradation.
+IV-I4 Ablation Study on Fused Depth Map Guided 3D Deformable Attention Module
 Vis-Depth Fusion Depth-Fusion Vis-Depth LiDAR-Depth IoU ($\uparrow$) mIoU ($\uparrow$) 46.24 (Baseline) 30.79 (Baseline) 46.06 (-0.18%) 30.57 (-0.22%) 39.80 (-6.44%) 25.28 (-5.51%) 44.01 (-2.23%) 28.98 (-1.81%) 43.99 (-2.25%) 28.93 (-1.86%) 40.33 (-5.91%) 24.71 (-6.08%)
-TABLE XV: Ablation study on Depth-Fusion-Based 3D Deformable Attention Module.
-$\uparrow$:the higher, the better.
-Vis-Depth Fusion: add the fused depth feature back to the visual feature, resulting in a depth-aware visual feature.
-Depth-Fusion: fuse the visual depth feature with the lidar depth feature.
-Vis-Depth: visual depth feature.
-Lidar-Depth: lidar depth feature.
-$\uparrow$:the higher, the better.
-We conducted an ablation study with respect to each component of the depth-fusion-based 3D deformable attention module and presented the experiment results in Table XV.
-In practice, Vis-Depth provides dense but less accurate depth maps, while Lidar-Depth provides sparse but more accurate depth maps.
-The experiment shows that dense depth maps contribute significantly to the performance, even though they are less accurate.
-Relying solely on sparse but accurate depth maps from lidar results in a significant performance degradation of 5.5%-6.0%.
-Furthermore, adding fused depth maps back to the original visual features results in depth-aware visual features contributing the second-largest performance improvement, around 2.2%.
-The experiment results demonstrate the necessity of multi-stage multi-modality feature fusion to boost model performance.
-IV-K3 Ablation Study On MGCAFusion Module
-Skip-Connect WSFusion GCFusion IoU ($\uparrow$) mIoU ($\uparrow$) 46.24 (Baseline) 30.79 (Baseline) 38.60 (-7.64%) 18.18 (-12.61%) 39.39 (-6.85%) 18.57 (-12.22%) 24.20 (-22.04%) 19.91 (-10.88%)
-TABLE XVI: Ablation study on the MGCAFusion Module.
+TABLE XI: Ablation study on Depth-Fusion-Based 3D Deformable Attention Module.
+Vis-Depth Fusion: add the fused depth maps back to the visual features, resulting in depth-aware visual features.
+Depth-Fusion: fuse the visual depth maps with the LiDAR depth maps.
+Vis-Depth: visual depth maps.
+LiDAR-Depth: LiDAR depth maps.
+We conducted an ablation study for each component of the fused-depth-guided 3D deformable attention module and presented the results in Table XI.
+In practice, Vis-Depth provides dense but less accurate depth maps that are good at capturing the partial occlusion of object structure.
+In contrast, LiDAR-Depth provides sparse but more accurate depth maps, especially good at capturing the front faces of objects in the scene.
+The experiment shows that Vis-Depth maps contribute significantly to the performance, even though they are less accurate.
+Relying solely on sparse but accurate LiDAR-Depth maps results in a performance degradation of 6.44% IoU and 5.51% mIoU, respectively.
+Furthermore, adding fused depth maps back to the original visual features results in depth-aware visual features contributing the second-largest performance improvement.
+The experiment results demonstrate the importance of depth map density and the necessity of multi-source depth map fusion and vis-depth feature fusion to boost model performance.
+IV-I5 Ablation Study on GCWSFusion Module
+We conducted an ablation study for each major component of the GCWSFusion module and present the experiment results in Table XII.
+Gate-atten Weight-sum Skip-Connect WSFusion GCFusion IoU ($\uparrow$) mIoU ($\uparrow$) 46.24 (Baseline) 30.79 (Baseline) 45.52 (-0.72%) 30.68 (-0.11%) 41.34 (-4.90%) 27.12 (-3.67%) 38.60 (-7.64%) 18.18 (-12.61%) 39.39 (-6.85%) 18.57 (-12.22%) 24.20 (-22.04%) 19.91 (-10.88%)
+TABLE XII: Ablation study on the GCWSFusion Module.
 Skip-Connet: skip-connection structure at the last feature fusion stage.
@@
 GCFusion: gated concatenation fusion part.
-$\uparrow$:the higher, the better.
-We conducted an ablation study for each major component of the MGCAFusion module and present the experiment results in Table XVI.
-The experiment results show that the weighted summation and gated concatenation fusion components contribute almost equally to the model’s final performance, with the former slightly more important.
+Gate-atten: Gate attention mechanism in GCFusion module.
+Weight-sum: Weighted summation mechanism in the WSFusion module.
+The results of the experiment show that the weight-summation mechanism plays a relatively more important role than the gate-attention mechanism in the overall GCWSFusion module.
+However, the WSFusion and GCFusion components contribute almost equally to the model’s final performance, with WSFusion slightly more important.
 Missing each part will cause 12.22% to 12.61% mIoU and 6.85% to 7.64% IoU performance degradation.
-The skip-connection structure plays an important role during the feature fusion.
-V Conclusion and Future work
-This paper presents TFusionOcc, a novel Student’s t-distribution-based object-centric multi-stage multi-sensor fusion framework.
-Our framework in the sensor fusion part adopts a multi-stage feature-level fusion strategy, including an early-stage, a middle-stage, and a late-stage fusion, to better integrate complementary information across modalities and mitigate the drawbacks of each modality.
-We also adopt an object-centric modeling strategy by leveraging the Student’s t-distribution and the T-Mixture model as a probability kernel, combined with several typical primitives, such as the general T-Primitive and the superquadric, to achieve greater robustness against outliers.
-We further propose a primitive called the deformable superquadric primitive (superquadric with inverse-warp), which offers greater geometric shape representation flexibility, thereby boosting the model’s 3D geometric shape-capturing capability in the driving scene.
-The extensive experiments conducted on the nuScenes [2] and nuScenes-C [31] datasets, including challenging cases of rainy, night, fog, snow and sensor-malfunction, demonstrate that our proposed approach not only achieves SOTA performance but also exhibits superior robustness.
-In our future research endeavors, we will investigate how to utilize a much smaller set of primitives and fewer rendering operations to achieve similar performance.
+The skip-connection structure also plays an important role during feature fusion; without it, the mIoU and IoU degrade by 22.04% and 10.88%, respectively.
+V Conclusion
+This paper presents TFusionOcc, a T-primitive-based object-centric multi-sensor fusion framework for 3D semantic occupancy prediction.
+The proposed method is built upon a family of Student’s t-distribution-based T-primitives, among which the deformable T-Superquadric with inverse warping serves as the key geometry-enhancing primitive.
+To further improve robustness, we introduce a unified probabilistic formulation based on Student’s t-distribution and the T-mixture model for joint occupancy and semantic modeling, together with a tightly coupled multi-stage fusion architecture for effective camera–LiDAR integration.
+Extensive experiments on SurrounOcc-nuScenes, Occ3D-nuScenes, and nuScenes-C demonstrate that TFusionOcc achieves strong performance and robustness in the most challenging conditions.
+In future work, we plan to further improve long-range perception and robustness under structured visual degradations, such as motion blur and snow.